A Complete Bibliography of Publications in Formal Aspects of Computing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

12 June 2023 Version 1.16

Title word cross-reference

$119.99 [985]. $129.00 [985]. $19.99 [934]. 2 [809]. 4 [471, 1010]. 6 × 2 [985]. $75.00 [934]. ^{++} [595]. ^{3} [854]. _{d} [486]. χ [445]. EB^{3} [855]. L^{∗} [1040]. Z [842]. μ [493, 907, 308]. ν [518]. Ω [325]. P [99]. ϕ [644]. π [723, 576]. ψ [644]. × [934].

* [1045].

-based [1040]. -calculus [723, 576, 907]. -complete [99]. -DOF [1010]. -phase [809]. -slot [471].

//doi.org/10.1007/978 [966]. //doi.org/10.1007/978-3-662-56641-1 [966]. //www.concrete [952].

//www.concrete-semantics.org/ [952]. //www.decision [965]. //www.decision-procedures.org/ [965].

’05 [513].

1 [966]. 1394 [308, 417].

2 [520]. 2-MDFAs [520]. 2007 [566]. 2009 [604]. 2010 [637]. 2016 [959].

3 [978]. 3rd [604].

517 [553]. 5th [604].

978-0-19-874782-6 [934]. 978-0-19-874783-3 [934]. 978-3-319-10542-0 [952]. 978-3-642-54117-9 [985].

978-3-662-50496-3 [965]. 978-3-662-52444-2 [985]. 978-3-662-56639-8 [966]. 978-3-662-56641-1 [966, 966].

Abadi [210]. abortive [517]. abridged [241]. Abstract [346, 397, 313, 427, 896, 835, 239, 252, 912, 671, 464, 197]. Abstraction [301, 351, 821, 350, 962, 491, 665, 889, 621]. Abstraction-Based [301]. Abstractions [771, 977, 63, 770]. AC [575]. accelerated [969]. Accelerating [845]. access [829, 620, 636, 602]. Accessing [162]. ACL2 [997]. across [778]. Action [433, 183, 510, 396, 244, 148, 261, 470, 452, 591]. action-based [261, 452]. actions [618, 110]. Active [860, 830]. Activity [1055, 348]. Actors [444]. ad [231, 757, 932]. Ada [112, 672]. adaptation [878, 873]. Adaptive [298, 946, 694]. adaptivity [874]. adequate [150]. ADL [348]. Adopted [411]. Advanced [431, 992]. Advances [387]. again [22]. against [962, 582]. agent [651, 889, 110, 683]. agent-based [683]. agents [900, 225]. aid [224, 1011]. Alexander [966]. Algebra [342, 620, 395, 370, 624, 52, 132, 227, 549, 675, 761, 484, 577, 1046, 142, 974, 707, 820, 617, 708, 636, 271]. Algebraic [366, 304, 512, 344, 861, 480, 254, 628, 252, 638, 477, 890, 63, 833, 195, 91, 98, 841, 264, 623, 611, 814]. algebras [151, 237, 23, 908, 44]. Algorithm [1016, 313, 1020, 477, 592, 262, 721, 720, 103, 25, 520, 863]. Algorithms [172, 542, 760, 913, 665, 272, 581, 929, 740, 710, 14, 801]. allow [949]. Alloy [556, 818]. Almost [492, 287]. Alphabet [282, 281]. alternating [579]. Alternative [947, 603]. Alternatives [316]. alway [265]. Ambient [844]. ambients [575]. amendment [991, 1006]. Amílcar [898]. Amir [644]. AMS [456]. analogue [456]. Analogy [400, 192]. Analogy-Based [400]. Analyse [418]. Analyses [323]. Analysing [850, 1054, 317, 806, 41]. Analysis [316, 337, 697, 169, 521, 1002, 845, 982, 571, 901, 899, 535, 971, 140, 545, 912, 575, 517, 201, 977, 1020, 812, 490, 470, 525, 831, 528, 989, 467, 698, 631, 468, 896, 1027, 802, 964, 885, 502, 14, 685, 629, 667]. analyzing [839]. Ancient [401]. André [1005]. Android [955]. Angelic [515]. Angle [1030]. annotated [911]. anonymity [819, 768]. Apple [8]. applicability [114]. Application [340, 778, 975, 466, 822, 919, 940]. Applications [92, 764, 431, 1001, 984, 478, 1061, 930, 847, 844, 1004]. applicative [743]. Applied [510, 423]. Approach [305, 352, 360, 397, 385, 594, 955, 273, 642, 679, 651, 552, 638, 16, 210, 694, 930, 503, 660, 848, 1014, 747, 942, 820, 578, 117, 839, 223, 1037, 1038, 647, 963]. Approaches [1012, 370, 827, 822, 992]. Approximation [443, 821, 520]. approximations [982]. ArcAngel [429]. Architectural [393, 226, 726, 990]. Architecture [321, 340, 859, 527, 701]. architectures [869, 971, 653]. Argos [469]. arithmetic [543]. array [154]. ASAP [492]. ASM [874, 778, 438]. ASM-based [874]. ASMs [651]. Aspects [553, 511, 65]. Assembling [995]. Assertion [696, 534]. Assertion-based [696, 534]. assertional [474, 25]. assertions [174]. Assessment [1026]. assignment [154]. assignments [130]. assistant [597]. Assisted [927]. associations [843]. associative [57]. Assume [794, 957]. Assume-guarantee [794, 957]. Assumption [911, 182, 225]. assumption/commitment [225]. assumptions [94]. Assurance [1050, 694]. Assuring [677]. ASTDs [798]. Asynchronous [189, 336, 286, 884, 941, 467, 139]. asynchronously [225]. ATM [354, 307]. Atomic [618, 474, 989]. Atomicity [746, 78]. atomized [278]. attacker [571]. attacks [962]. attentional [611]. attribute [41]. Audio [298]. augmented [1031]. August [927]. authorization [877]. authors [65]. AUTOFOCUS [359]. Automata [403, 368, 329, 845, 1018, 484, 780, 853, 739, 917, 1024, 581, 646, 244, 501, 806, 834, 596, 1044]. automata-based [834]. Automated [663, 957, 447, 682, 976, 14, 439, 918, 774]. Automatic [74, 763, 310, 362, 339, 634, 947, 968]. Automating [972, 629]. Automatizing [170]. automotive [584]. autonomous [1054, 1008, 900, 1031]. Auxiliary [335]. avionics [850]. avoiding [577]. aware [978, 1046]. awareness [844]. axiom [53]. Axiomatic [81, 813, 1037]. Axiomatisation [454]. Axiomatization [281]. Axioms [27, 282].

B [655, 776, 352, 558, 714, 862, 540, 828, 656, 691, 886, 650, 979, 1014, 682, 972, 495, 775, 811, 1011, 653]. Backus [566]. backwards [267]. Bakery [863]. Balancing [827]. Barrier [454, 1034]. Based [378, 430, 169, 301, 400, 286, 1002, 982, 874, 594, 904, 696, 62, 513, 535, 1026, 679, 131, 753, 484, 544, 1016, 937, 829, 784, 912, 797, 652, 873, 787, 562, 630, 261, 938, 930, 592, 974, 1063, 903, 981, 849, 605, 272, 189, 1014, 452, 569, 822, 683, 506, 727, 779, 619, 534, 929, 213, 783, 887, 715, 157, 755, 684, 97, 956, 1012, 196, 1056, 122, 733, 940, 834, 1040, 205, 852, 825, 629]. basic [594, 129, 205]. basis [105, 974]. Battery [978]. Battery-aware [978]. be [391]. before [939]. Behavior [697, 749]. behavioral [849, 466, 875]. Behaviour [317, 845, 58, 626, 613, 26]. Behavioural [339, 63, 775, 936]. behaviours [119]. beliefs [900]. Berlin [985]. Bernhard [992]. between [273, 505, 805, 171]. Beyond [601, 17]. bigraphs [785, 738]. BigrTiMo [1025]. Binding [397, 738, 671]. biphase [139, 521]. Bird [216]. Birkhoff [922]. bisimilarity [99]. Bisimulation [998, 904, 107, 853, 250, 550]. Bisimulations [573, 867, 725]. bit [471]. black [673, 863]. black-box [673]. Blaming [640]. blending [525]. blocking [760, 234]. blocks [38]. Bloem [986]. bodies [224]. Book [296, 90, 179, 985, 230, 966, 1004, 206, 15, 126, 29, 952, 72, 146, 95, 934, 243, 1005, 986, 135, 236, 992, 212, 6, 965, 194, 54, 20, 47, 120, 200, 187]. Boolean [910, 969]. Börger [966]. bounded [809, 668, 926]. bounded-time [809]. bounds [645]. Bowen [934]. box [673]. Boyer [97, 113]. branches [939]. Branching [22]. Broad [390]. Broad-Spectrum [390]. Broadcast [465]. BSP [670]. Büchi [1024, 596]. budgeting [787]. buffer [89]. buffers [648]. bug [1061]. bugs [939]. build [537]. Building [861, 678, 39, 717]. Burstall [388, 389, 387]. Bus [418]. business [609, 635]. button [678]. bytecode [568].

C [927, 658, 831, 889]. CA [785]. Cache [333]. CADP [882, 854, 855]. CafeOBJ [902]. Calculate [101]. calculational [276]. Calculator [322]. calculi [573, 922]. Calculus [440, 294, 413, 460, 312, 431, 1035, 315, 578, 318, 523, 254, 176, 231, 797, 723, 576, 178, 266, 13, 754, 713, 184, 508, 461, 241, 757, 21, 455, 277, 118, 907, 86, 129, 1025, 844]. call [215, 1015]. callbacks [519]. calls [519]. Cameo [603]. Camera [1016]. Can [427]. CaPiTo [674]. Capture [577, 611]. Capture-avoiding [577]. carrying [533]. Cars [1030]. Case [380, 299, 356, 316, 415, 307, 845, 624, 1064, 527, 978, 946, 70, 258, 783, 944, 802, 516, 666, 190]. case-study [190]. Cases [1050, 946]. Cash [361, 356, 354, 357, 358, 355]. Cash-Point [361, 356, 354, 357, 358]. CASL [393]. Categorical [842, 124]. categorically [101]. Category [124]. Category-theoretic [124]. causally [503]. CCP [469]. CCS [131, 229, 226, 14]. CCS-based [131]. Celebrating [387]. Cell [1010]. centred [797]. centric [873, 876]. certifiably [939]. certificates [1034]. certification [555]. certified [624]. Certifying [472]. Chain [31]. Chalice [712]. challenge [537]. challenges [532, 678]. Changing [311]. channel [716, 983, 1047, 810]. Channels [377]. characterisation [98]. Characterisations [723]. characterization [738]. charts [779]. checked [104, 477]. Checker [1054, 652, 862, 79, 339, 514]. checkers [904]. Checking [300, 306, 1051, 421, 729, 613, 679, 937, 1061, 840, 459, 540, 460, 464, 1021, 560, 739, 487, 178, 668, 677, 645, 889, 695, 942, 697, 578, 906, 866, 876, 847, 737, 485, 743, 926, 591, 824, 504, 986]. checking-based [679]. checks [556]. chemical [754]. chemistry [198]. chip [567]. choice [93, 714, 278, 733]. CICS [162]. circuit [104]. circuits [74, 750, 751, 980, 472, 271, 279]. circular [957, 836]. Circus [435, 586, 672]. Clarification [332]. Clarke [986]. Class [343, 74, 565, 784, 903, 153]. classes [197]. classification [667]. classroom [597]. clause [909]. client [879, 640]. client/server [879]. Clock [232, 542]. clocked [751, 279]. closed [488]. Closure [426]. closures [31, 240]. Cloud [874, 1014]. CML [847]. Co [301, 224, 907]. co-invariant [224]. co-logic [907]. Co-operating [301]. Code [342, 962, 533, 624, 1029, 677, 889, 712, 631, 704]. Code-carrying [533]. Cognitive [317, 363, 340, 612]. Coherency [333]. Coinduction [998]. colimits [37]. Collaborative [1008, 985]. Collection [387]. collector [981, 152]. coloured [883, 565]. combination [542, 458, 278]. combinational [750]. combinator [105]. combinators [661]. Combining [352, 383, 142, 802]. comedy [175].

Command [23, 744]. commitment [225]. Communicating [130, 402, 64, 259, 581, 467, 569, 279, 1027, 495, 948, 225, 806, 509]. Communication [377, 286, 1002, 760, 41, 884, 873, 476, 189, 139]. communication-centric [873]. Communications [406, 140, 878, 567]. compact [1020]. companion [966]. Comparative [423, 527, 70]. Comparing [459]. Comparison [386, 901, 941, 958, 598, 195, 818]. Compensation [788]. compilation [508, 506, 896, 279]. compiled [704]. Compiler [531, 168, 507, 228, 46]. Compilers [346, 638]. Complete [903, 473, 99, 1039, 125, 483, 632, 851]. Completeness [458, 897, 138]. completions [94]. complex [1029]. complexity [863, 923]. Compliant [1060]. Component [342, 383, 866, 874, 984, 787, 930, 848, 849, 929, 887]. component-based [930, 849, 929, 887]. Component-wise [866]. components [613, 475]. composability [859]. Composing [345]. composite [1014]. Composition [412, 444, 344, 148, 748, 868, 452, 217]. Compositional [433, 93, 830, 1055, 248, 386, 683, 687, 205, 920, 232, 576, 479, 834]. compositionally [469]. compositions [695]. Comprehensive [1059]. Computable [250]. Computation [950, 888, 791, 969, 267, 137, 733]. Computations [512, 511]. Computer [927, 340, 387, 535, 604, 552, 477, 735, 2]. computer-based [535]. Computerized [299]. Computing [867, 314, 349, 65, 553]. concepts [33, 784, 605, 264]. Concerned [807]. Concise [1004]. Concolic [703]. Concrete [952]. Concurrency [434, 929, 252, 603, 924, 890, 974, 827, 11, 919, 600, 601, 944, 157]. Concurrency-preserving [929]. Concurrent [498, 166, 436, 170, 347, 171, 883, 982, 536, 588, 776, 941, 182, 780, 1039, 562, 16, 479, 77, 261, 32, 863, 923, 665, 981, 195, 839, 97, 196, 685, 719, 260]. Conditional [217]. Conditions [402]. confidentiality [800, 766]. configuration [112, 635]. configurations [835]. Configured [379]. confluence [894]. Conformance [368, 373, 51, 871, 677]. Congress [349]. Conjunctive [315, 896]. connection [273]. conquer [216]. consequences [553]. conservative [117]. considerations [253]. Consistency [979, 1059, 1060, 540, 858, 235]. Consistency-preserving [979]. Consistent [630, 503]. constants [40]. constrained [462]. Constraint [895, 894, 715, 622, 740]. Constraint-based [715]. Constraints [443, 910, 239, 759, 633]. construct [653]. Constructing [754, 504]. Construction [535, 861, 565, 105, 868, 660, 474, 942]. constructions [254, 277, 86]. constructive [9]. container [954]. Content [448]. Contention [421, 423]. Context [63, 668, 698, 844]. context-dependent [698]. Continuation [330]. Continuous [778, 312, 694]. continuous/discrete [778]. Contract [852, 794]. Contract-based [852]. Contracts [600, 833, 601]. contractually [268]. Contribution [387]. control [778, 672, 793, 620, 660, 849, 103, 602, 502, 636]. control-flow [502]. controllable [582]. controller [1018, 1008]. controllers [1009, 915, 486, 988, 1034]. Conventions [411]. Convergence [171]. conversations [877, 925]. conversion [564]. converters [228]. convex [488]. Cooperating [375]. Copeland [934]. Coq [568]. Coquand [86]. coroutines [896]. Correct [343, 868, 154, 506, 46, 801]. Correct-by-construction [868]. correcting [983]. Correction [1006]. Correctness [333, 366, 512, 863, 511, 58, 239, 642, 584, 1039, 259, 1058, 103, 715, 758, 635]. Cost [917]. costs [984]. Counterexample [1042].

Counterexample-guided [1042].

Coupling [1055]. courses [595]. coverage [793]. criteria [650, 502]. criterion [78]. critical [1031, 730]. Critics [324]. CRL [493, 308]. Cryptographic [767, 571, 641, 543]. CSMA [785]. CSMA/CA [785]. CSP [551, 352, 383, 428, 540, 606, 1003, 768, 249, 737, 1047, 482, 729, 351, 495, 350]. CSP# [799]. csp2B [352]. CtCoq [321]. CTL* [684]. CTML [950]. CTS [785, 810]. custom [961]. Cut [939, 697, 897]. cyber [1005]. cyber-physical [1005]. cyclic [845, 717, 257].

Daniel [965]. Data [367, 406, 304, 427, 165, 215, 570, 365, 350, 496, 793, 651, 1039, 961, 640, 125, 736, 876, 257, 837, 1053, 770, 690, 843, 514, 190, 154, 943]. data-based [1063]. data-centric [876]. data-enriched [837]. data-flow [651]. database [69, 71, 808]. Dataflow [167, 442, 695]. datatypes [319, 453]. David [36]. DC [502]. Deadline [421]. deadlines [739]. deadlock [131, 983]. deadlocks [498]. Dec [566]. decidability [667]. Deciding [99, 853, 550]. decision [684, 1040, 659, 965]. Decisions [510]. declarative [900]. Decoding [367]. decomposition [916, 622]. deduction [896]. Deductive [579, 542, 745]. Deep [385, 1033, 804, 1038]. Defining [41]. Definition [293, 129, 256, 755]. definitions [235]. Delimited [376]. delta [991, 1006]. Demand [167]. demonstrably [46]. Demonstrating [363]. Denotational [116, 1047, 1012, 814, 658, 103]. Dense [345, 761, 460]. dense-time [460]. dependable [536, 901]. dependent [58, 698]. Dependently [399]. deployment [930]. Derivation [776, 392, 946, 592, 801, 814]. derivations [643, 562]. derivative [513]. derived [25]. Deriving [626, 330, 904, 761, 224]. describe [605]. described [559]. Description [348, 488, 899, 45]. Design [305, 321, 362, 316, 323, 313, 400, 19, 874, 778, 985, 536, 104, 486, 638, 916, 1046, 271, 605, 528, 788, 990, 988, 631, 947, 253, 733, 114, 808]. design-based [733]. Designing [924, 269]. designs [614]. desires [900]. Detection [422, 955, 893, 1061, 812, 631]. determined [221]. determinism [50, 840]. Deterministic [373, 374, 719, 582, 473, 848, 452]. Deva [129]. Development [416, 380, 406, 430, 2, 349, 655, 558, 787, 1021, 9, 328, 847, 887, 264, 222, 86, 1056, 811, 1011]. developments [223]. devices [615]. diagnose [960]. Diagrams [165, 339, 672, 607, 546, 690, 735, 519]. diet [918]. difference [740]. differential [1037]. digital [794]. Dijkstra [410, 709]. Directed [346, 574, 786, 82]. Dirk [1035]. Disciplined [453]. DisCo [361]. Discovering [478, 983]. Discovery [774, 671]. Discrete [227, 778, 178, 731, 114]. Dispenser [355]. distinguishing [946, 805]. Distributed [380, 413, 446, 375, 313, 170, 108, 172, 835, 680, 941, 486, 1021, 477, 503, 476, 722, 162, 272, 258, 813, 1056, 710, 504]. Distributing [403]. Distributions [370]. diversity [884]. divide [216]. divide-and-conquer [216]. divided [108]. Do [3]. Do-it-yourself [3]. DOF [1010]. doi.org [966]. Domain [305, 273, 822, 631]. domains [899]. Don’t [420]. double [488]. Draheim [1035]. Drawing [1049]. Driven [167, 869, 868, 988, 1011, 814]. Driving [1030]. DSL [858]. dual [907]. duality [176, 201]. Duration [460, 266, 312, 431, 578, 462, 178, 461]. Duration-constrained [462]. during [635]. Dynamic [933, 877, 901, 1063, 650, 664, 822, 876, 632, 485, 634, 824, 745]. Dynamical [408, 1032]. dynamically [149]. dynamics [755, 943].

E6 [555]. EB [854]. eBook [966]. echelon [888]. Editing [326]. Editor [400]. Editorial [1048, 320, 424, 432, 572, 953, 593, 494, 513, 585, 616, 686, 765, 772, 649, 741, 219, 441, 463, 832, 856, 1057, 872, 681, 1007, 610, 1028, 7, 61, 164, 188, 274, 381, 405, 414, 489, 524, 530, 935, 56, 67, 499, 826, 905, 1052, 625, 693, 967, 857, 297, 973, 43, 303, 1, 49, 96, 128, 141, 371, 409, 449, 457, 547, 554, 781, 795, 994, 676, 539, 481, 914, 657, 1036, 864, 865, 880, 599, 891, 987, 263, 699, 838, 173, 1017, 513]. Edmund [986]. eds [985, 986]. Edsger [410]. effects [1022, 632]. Efficient [982, 571, 1053, 565, 490, 453, 932, 1015]. Egon [966]. egress [698]. Eiffel [551, 603, 455]. eISBN [966]. elastic [930]. electronic [131, 560, 557, 556, 555]. electrophysiology [611]. elementary [642, 185, 98, 790]. Eliminating [53]. Elucidating [665]. embedded [985, 787, 889, 926]. embedding [718]. Embeddings [396]. embeds [469]. Emergence [689]. emergency [698]. empirical [901]. encapsulation [534]. Encoding [367, 948]. Encodings [314, 770]. energy [1018]. enforcement [1013]. Engineering [445, 1062, 529, 937, 868, 563]. enhancement [810]. Enhancements [998]. Enhancing [1051, 1033]. enriched [837]. ensuring [871]. enterprise [984]. environment [902]. environments [1009]. equalities [910]. Equality [110]. Equational [908, 28, 395]. equations [273, 639, 1037]. Equivalence [256, 904, 150, 894, 870, 107, 94, 903, 836, 250]. equivalence-completions [94]. equivalences [59]. equivalent [570]. ERC [455]. Errata [553]. Erratum [636]. Error [156, 802, 810]. error-prone [810]. errors [242, 175]. Essence [782]. Essential [264]. Esterel [362]. Estimating [984]. Estimation [1016, 975, 946, 114]. Ethernet [1002]. EUR [952, 965]. Europe [450]. evaluate [1001]. Evaluation [167, 298, 505, 893, 10, 929]. Event [873, 11, 196, 8, 714, 828, 974, 1063, 814, 655, 776, 558, 714, 828, 656, 691, 886, 979, 1014, 972, 775, 811, 1011, 653]. Event-B [655, 776, 558, 714, 828, 656, 691, 886, 979, 1014, 972, 775, 811, 1011, 653]. Event-based [873, 974]. event-driven [814]. event/data [1063]. event/data-based [1063]. events [24, 30, 35, 42, 48, 55, 60, 66, 73, 106, 76, 80, 84, 111, 115, 121, 127, 136, 498]. everyday [767]. everyone [766]. Eves [561]. evolving [220, 149]. EWD1300 [411]. Example [282, 355, 981, 272]. exclusion [762, 923, 14]. executable [112]. Execution [446, 443, 845, 1001, 1029, 1041]. Execution-Time [443]. exercise [808]. existence [1037]. existential [1045]. Exits [184]. Expectations [531]. experiences [594]. experiment [57]. Experiments [691, 542]. explained [1041]. Explanation [760, 193]. Explicit [232, 116, 942]. Exploiting [1031]. explored [471]. Explorer [590]. Exploring [497, 725, 437]. explosives [161, 123]. expressions [221, 110, 462]. Expressive [288, 875, 482]. expressiveness [827]. Extended [313, 1023, 1045, 252, 860, 916, 71, 119, 1040, 799]. Extending [177, 153, 808]. extension [245, 239, 110]. Extensional [970, 442]. External [714]. extra [777]. extremal [247].

FAC [513]. Face [311]. facile [198]. Factorisation [59]. FACTum [990]. failure [1002, 746, 205, 175]. failure-based [205]. failures [509, 960, 570]. fair [490, 269]. fairness [94, 11, 158]. False [422]. falsification [621]. families [155]. Family [1062, 937]. family-based [937]. fast [25]. Fault [918, 750, 751, 104, 809, 931, 87]. fault-tolerance [931, 87]. fault-tolerant [104]. faulty [716]. FDR [590]. feature [937]. feature-oriented [937]. Features [431]. February [927]. feedback [751].

FermaT [801]. Fifty [999]. file [688, 704]. filesystem [537]. Filter [718, 1038]. Filter-embedding [718]. filters [1043]. Finding [977, 843, 556]. fine [792]. fine-grained [792]. Finite [403, 512, 860, 723, 946, 248, 59]. FireWire [418, 421, 419, 423]. First [451, 353, 349, 32, 687]. First-Order [353]. Fitzgerald [985]. fixed [50]. flash [703]. flaws [955]. Flexible [602, 606]. Flow [207, 165, 975, 878, 651, 103, 502]. flows [748]. Floyd [709]. fold [701]. fold-in [701]. fold-out [701]. folds [319, 453]. Forever [420]. form [888]. Formal [1002, 380, 769, 298, 309, 494, 299, 1004, 971, 300, 289, 311, 1020, 812, 1050, 386, 525, 882, 789, 281, 165, 293, 65, 291, 1010, 511, 964, 423, 349, 808, 943, 624, 874, 955, 778, 800, 1001, 505, 45, 256, 607, 815, 552, 486, 598, 871, 261, 694, 995, 503, 660, 199, 827, 270, 161, 595, 989, 563, 139, 123, 268, 770, 755, 803, 948, 885, 223, 253, 940, 1059, 193, 1025, 799, 936, 450, 553]. Formalisation [888, 445, 804]. formalised [33, 730]. Formalising [883, 318]. Formalism [382, 201, 360, 216, 605]. formalisms [901]. Formalization [1016, 340, 295, 140, 622, 567, 502]. Formalizing [254, 688, 568, 858, 819]. Formally [945, 869, 756, 1041]. formation [774, 755]. Formula [1016, 163]. Formulae [1019]. formulas [908]. Forthcoming [24, 30, 35, 42, 48, 55, 60, 66, 73, 106, 76, 80, 84, 111, 115, 121, 127, 136]. Foundation [512, 535, 45, 797, 890, 86, 940, 1059]. Foundations [531, 886, 992, 1065, 842, 266, 1005]. Four [280, 1049]. fragments [575]. Frama [831]. Frama-C [831]. frames [664, 634]. Framework [920, 400, 220, 859, 980, 252, 1029, 607, 912, 695, 608, 816, 467, 988, 803, 641, 623]. free [93, 592, 762, 235]. FreeRTOS [815]. FRS [927]. FSP [646]. full [870]. fully [954, 197]. Fun [423]. Function [344, 251]. Functional [58, 16, 82, 694, 138, 257, 567, 137]. functionality [800]. functions [478]. Fundamentals [1004]. fusion [718]. futures [830]. fuzzy [821].

gain [673]. GALS [882]. game [641]. game-theoretic [641]. games [917, 1013]. garbage [981, 152]. Gaussian [885]. general [859, 576]. Generalisations [395]. Generalised [376, 319, 544, 375, 890, 453]. generalized [464]. Generating [624, 374, 650, 841]. Generation [346, 851, 750, 751, 93, 604, 846, 574, 793, 783, 947]. Generator [1026, 44, 224]. generators [961, 677]. Generic [412, 632, 1044]. Genius [410]. Gentle [644]. geographical [1053]. geometric [622]. Gerard [1004]. Gerwin [952]. Giant [644]. Gift [1011]. global [680]. goal [626, 786]. goal-directed [786]. GomX [978]. GomX-3 [978]. Gordon [927]. Gorm [985]. GPU [969]. GPU-accelerated [969]. GR [1045, 1019]. grained [792]. Grammar [375]. Grammars [1060, 41]. Graph [341, 1060, 589, 1064, 918, 633]. Graphical [322]. Graphs [207, 396, 696, 108]. grid [1002]. GRL [882]. ground [535]. groups [714, 843]. Guarantee [289, 648, 794, 957, 890, 981, 260]. guarantees [1045]. guards [142]. Guest [572, 494, 499, 539, 481, 513]. Guide [1004, 65, 934]. guided [1042, 612].

halting [39]. Handbook [986]. Handel [658]. Handel-C [658]. handlers [1022]. Handling [894]. Hardback [952, 934, 985, 965]. hardware [228, 497, 506, 543, 1000]. hardware/software [497]. Haskell [717]. heap [712]. heap-manipulating [712]. Heidelberg [985]. Heinz [817]. Helmut [986]. helps [960]. Hennessy [13]. Henzinger [986]. Herman [721, 720]. Heterogeneous [291]. heuristic [755]. hierarchical [688, 799]. hierarchically [108, 843]. hierarchy [784]. High [418]. High-Performance [418]. Higher [763, 1023, 33, 478, 215, 137]. Higher-order [1023]. Highly [536]. Historical [998]. history [258]. Hoare [999, 742, 186, 142, 177, 335]. hoc [231, 757, 932]. HOL [888, 952, 819, 318, 996, 1010, 885, 326]. home [93]. homological [624]. Hopfield [40]. Horn [909, 907]. http [952, 965]. https [966]. Human [340, 552, 802, 611]. Humble [662]. Humorous [662]. Huth [992]. Hybrid [975, 748, 1063, 431, 331, 920, 922]. hybridised [823]. HYPE [748]. hypertraces [100].

Identification [418]. Identify [416, 415, 422, 419, 308, 417]. IEEE [416, 418, 785, 421, 415, 422, 419, 308, 423, 417, 810]. if [420]. II [751, 588, 328, 466]. III [780]. ILP [626]. Impact [253, 735]. imperative [275, 267]. implementation [82, 582, 259, 983, 929]. implementations [769, 492, 63, 483, 779, 945]. implementing [41]. Implicit [289]. impredicative [112]. Improved [740]. improvement [956]. Improving [916]. Inadequate [398]. incompleteness [892]. Inconsistent [292]. Incremental [416, 313, 926, 574, 558, 792, 866, 152]. independent [870]. induction [133, 63, 44, 976]. Inductive [766, 155, 992, 1042, 738, 747, 277]. industrial [871, 847]. Industry [1004]. Inequations [427]. Inference [169, 255, 122]. Inferring [1032]. Infinite [171, 183]. Informatics [992]. Information [673, 62, 975, 878, 951, 822]. infrastructure [1049]. infusion [802]. inherent [735]. inheritance [732, 51, 37]. initial [273, 908]. Injection [1010]. input [851]. input/output [851]. inspired [604]. Institution [394]. institutions [1063, 823]. instruction [653]. Integer [512, 1060]. Integrated [494, 382, 467]. Integrating [32, 563, 811, 800]. Integration [309, 456, 231, 1050, 291, 655, 630, 848]. intelligence [1031]. Intensional [442]. intensive [1061]. Inter [648]. Inter-process [648]. interacting [613, 545]. Interaction [340, 552, 615]. interactions [879]. Interactive [363, 540, 324, 990, 1023, 611]. interconnection [40]. interest [727]. Interface [564, 322, 316, 323, 283, 325, 778, 614, 248, 204]. Interfaces [369]. interference [771, 568]. Interleaving [341, 549, 11, 919]. Interlocking [299, 300, 1021, 1053]. Internal [588, 285, 714]. Internet [298]. interoperability [949]. interoperating [475]. interpretation [912]. interpretations [185]. interpreters [10]. Interpreting [207]. interruptions [702]. intersection [88]. Interval [565, 511]. intransitive [933, 726]. Introducing [777]. Introduction [415, 392, 384, 209, 88]. Invariance [407, 716]. Invariant [594, 619, 690, 224, 574, 972]. Invariant-based [619]. Invariants [240, 329, 583, 731, 774, 534, 841]. inversion [267]. inverted [251]. Investigating [309, 981]. investigation [131]. Invocation [289]. IPv6 [337]. Isabelle [888, 1050, 952, 819, 17, 996]. Isabelle/HOL [888, 952, 819, 996]. Isabelle/SACM [1050]. ISBN [985, 966, 952, 934, 965]. isolated [618]. Issue [513, 530]. Issues [19]. iterative [988]. ITSEC [555].

J [927]. Jack [934]. Jacobian [964]. Jan [147, 180]. January [637, 934]. Jape [322]. Java [730, 893, 563, 634, 1015]. JavaSPI [945]. John [985, 566, 700]. Jonathan [934]. June [604].

K-Putput [991, 1006]. Kaisa [773].

Kalman [1043]. KAOS [778]. kernel [807]. kinematic [964]. KIV [557]. Kleene [709]. Klein [952]. Knaster [290]. knowledge [571]. Kobayashi [286]. Kroening [965].

L [147, 180, 325]. labeled [846]. Lambda [105, 1035]. Lamport [210]. Landin [604]. lane [915]. Language [445, 434, 828, 429, 348, 390, 519, 1062, 19, 83, 496, 45, 881, 870, 924, 941, 69, 116, 950, 568, 5, 4, 257, 808, 814]. language-independent [870]. Languages [305, 288, 252, 638, 639, 661, 122, 1065, 629]. large [682, 969]. Larsen [985]. Late [510]. lattice [92]. law [672, 57]. Laws [825]. layer [663]. Layered [710, 83, 476]. layers [584, 665]. lazy [647]. LCF [996]. Learning [1034, 1029, 860, 889, 1040, 596]. Least [315]. Legislation [1011]. Legislation-driven [1011]. length [946, 740]. lengths [645]. lenses [991, 1006]. lessons [479]. Level [555, 137, 814, 1062]. libraries [761]. library [954]. lifted [977]. Light [1010]. Limits [270, 981, 737]. line [790]. Linear [396, 443, 329, 1060, 286, 742, 761, 565, 22, 544, 79, 759, 27, 886, 189, 461, 820, 708, 841, 596, 659]. Linearizability [1000, 654, 713, 928]. Lines [1062]. linked [544]. Links [385]. Lip [310]. Lip-Synchronisation [310]. Lisp [997]. List [109, 392, 258]. literate [223]. liveness [158, 1037]. load [612]. Local [338, 407, 982, 680, 162]. local-analysis-based [982]. localities [144, 143]. lock [592]. lock-free [592]. Löf’s [57]. Logic [763, 345, 292, 430, 332, 335, 353, 392, 417, 232, 286, 999, 895, 742, 33, 949, 458, 654, 648, 186, 620, 636, 705, 897, 79, 909, 142, 159, 32, 1063, 327, 518, 27, 886, 13, 177, 189, 461, 569, 906, 632, 138, 715, 523, 53, 238, 907, 158, 137, 976, 18, 203, 596, 745]. Logical [531, 266, 220, 459, 328, 267, 1005]. logics [22, 922, 526, 538, 548, 117, 821]. long [163]. looking [939]. Loop [422, 224]. loops [621, 841]. Lord [644]. lost [989]. LOTOS [317, 646, 293, 119, 205]. Lovely [325]. low [978]. Lower [645]. LTL [942, 866]. Lustre [469].

M [986]. Machine [104, 1029, 377, 362, 373, 895, 582, 402, 473, 848, 39]. Machines [376, 346, 375, 372, 374, 883, 500, 716, 860, 946, 82, 1046, 59, 495, 39]. Macintosh^{T M} [8]. mail [131]. maintenance [108]. malware [893]. Management [1059, 1060, 485]. manager [8]. maneuvers [915]. Manifest [899]. manipulating [712]. Manna [993, 238]. MapReduce [718]. Marcel [985]. March [637, 566]. Marek [796]. Mark [934, 139, 521]. Markov [917, 1040]. Martin [57]. Martin-Löf’s [57]. Matching [531, 738, 921, 25]. mathematical [86, 992]. mathematics [643]. MATLAB [852]. MATLAB-style [852]. matrices [799]. matrix [888, 852]. Maude [902]. Maximal [892, 867]. maximality [846]. Maximally [347]. maximum [108]. may [724]. may/must [724]. MC [502]. MC/ [502]. McCarthy [700]. mCRL2 [695]. MDESL [1012]. MDFAs [520]. mealy [851]. Measure [1019]. measurement [199, 950]. mechanical [480, 750, 751, 262, 275]. Mechanically [416, 139, 152]. Mechanised [692]. Mechanising [561]. Mechanism [298, 1003, 471]. Mechanized [958, 287]. Mechanizing [479]. Meertens [216]. Meet [441, 709]. MEGA [325]. membranes [573]. Memoirs [387]. Memoriam [773, 817]. memory [730, 756, 974, 222, 1000]. Message [404, 207, 237]. Message-Passing [404, 237]. messaging [789]. Meta [10, 630, 129]. meta-calculus [129]. Meta-interpreters [10]. metadata [887]. metamodelling [804]. metastability [471]. MetateM [209]. Method [438, 291, 488, 800, 1039, 210, 211, 563, 4, 556, 260, 1015]. methodologies [461]. Methodology [380, 92, 787, 269, 609]. Methods [309, 494, 1004, 311, 450, 423, 349, 500, 256, 995, 270, 595, 1049, 841, 2, 114]. Metric [171, 273]. metrics [943]. Mexitl [317]. Michael [992, 36, 927]. Michel [662]. Microprocessors [366]. Milestones [997]. Milner [637, 13]. Mind [401]. mini [537]. Minimal [322, 526, 538, 548]. minimisation [248]. minimization [1061, 520, 1044]. Miranda [18, 203]. mission [825]. mission-based [825]. ML [124, 71, 391, 400]. mm [985, 934]. Mobile [378, 679, 575, 91, 1025, 337]. mobility [829, 1047]. MOBY [357]. MOBY/ [357]. modal [538, 548]. Model [869, 1054, 613, 1026, 300, 840, 306, 1051, 460, 938, 739, 487, 178, 668, 889, 421, 355, 339, 578, 876, 847, 743, 956, 591, 824, 1059, 286, 679, 551, 603, 730, 815, 937, 924, 1061, 652, 459, 991, 1006, 862, 234, 671, 261, 79, 464, 1021, 560, 526, 868, 497, 606, 903, 677, 189, 695, 986, 581, 816, 697, 1058, 866, 139, 737, 485, 21, 556, 947, 926, 514, 670, 733, 940, 193, 936, 635]. Model-based [1026, 938, 956, 903]. Model-checker [339]. Model-checking [613, 460, 739, 178, 695, 737]. Model-driven [869]. model-finding [556]. model-oriented [21]. model-to-model [936]. Modeling [434, 476, 1003, 559, 1027, 810, 932, 778, 980, 607, 1055, 803, 1031, 966]. Modelling [298, 785, 881, 357, 302, 37, 359, 920, 573, 901, 456, 615, 748, 882, 612, 611, 519]. Models [360, 366, 398, 1062, 845, 626, 1064, 627, 535, 776, 492, 630, 1008, 27, 754, 979, 682, 698, 783, 770, 666, 1000, 821, 806, 205]. modes [174]. Modular [1026, 1022, 68, 124]. modularisation [77]. Modularising [71]. Modulated [334]. modules [145]. Modulo [608, 894]. MOF [628]. Monadicity [395]. monads [661]. Mondex [558, 561, 560, 557, 559, 556, 555]. monitoring [929, 925, 516]. monotone [951]. Moore [97, 113]. Morphisms [394, 102]. movie [113]. MPSoC [943]. MSCs [465]. MSVL [813]. Multi [360, 1059, 1062, 745, 915, 131, 984, 651, 544, 703, 929]. multi-agent [651]. multi-component [984]. Multi-formalism [360]. multi-lane [915]. Multi-level [1062]. Multi-Model [1059]. multi-pattern-based [544]. multi-process [131]. multi-sector [703]. multi-threaded [745, 929]. multiary [991, 1006]. multiparty [878, 877, 925]. Multiple [991, 37, 1006]. Multitraces [100]. mural [192]. MUS [380]. must [724]. Mutation [440, 587]. mutual [975, 762, 923, 976, 14].

names [876]. Natural [346, 881, 615]. natural-language [881]. NDB [69]. necessarily [488]. need [711]. neighbourhood [458]. nested [319, 82, 453]. nested-stack [82]. Net [403, 408, 565, 576, 970, 272, 98]. nets [883, 26, 93, 846, 752, 32, 876, 40, 667]. Network [583, 91, 1034]. Networks [1055, 348, 329, 1033, 1002, 1064, 913, 812, 754, 646, 757, 969, 824, 225, 1038, 932]. Neural [1033, 1038, 1034]. neurobiological [806]. NewThink [218]. Nipkow [952]. Nitpick [337]. nominal [577]. Non [373, 374, 626, 50, 760, 694, 490, 582, 771, 568, 11, 452, 255, 841]. non-blocking [760]. non-controllable [582]. non-determinism [50]. Non-Deterministic [373, 582, 452]. non-functional [694]. non-interference [771, 568]. non-interleaving [11]. non-linear [841]. non-repudiation [490]. non-structural [255]. non-Zeno [626]. Nondeterminacy [284]. nondeterminism [183, 515, 735]. nondeterministic [946, 903, 28]. noninterference [933, 687, 729, 726]. Nonlinear [1032, 744]. normal [251]. Normalising [57]. Notational [411]. Notations [291]. Note [389, 286]. notions [726]. novel [839]. number [764]. NUT [358]. obfuscation [962, 892]. obfuscations [893]. Obituary [425, 36, 898]. Object [343, 153, 391, 19, 83, 252, 784, 638, 116, 532, 214, 534, 455, 589, 383, 428, 385, 197, 384]. Object-extending [153]. Object-Oriented [391, 19, 83, 784, 638, 116, 532, 455, 589]. Object-Z [383, 428, 385, 197, 384]. Objects [302, 830, 234, 767, 736]. obligations [38, 656]. observable [98]. observational [840, 59]. OCL [559]. Ode [388]. Ofer [965]. OLAP [820]. Oliver [992]. one [105]. one-combinator [105]. Online [451]. onto [154]. Ontologies [1051]. opacity [958]. operating [301]. operation [703]. Operational [909, 784]. Operations [285, 588, 777, 908, 191]. operators [868, 156]. optical [980, 885]. Optimal [1018, 497, 740]. Optimality [511]. Optimising [862]. Optimized [168]. Optimizing [913]. orbit [978]. orchestration [791]. orchestrations [758]. Order [763, 353, 171, 33, 273, 245, 26, 231, 478, 306, 862, 32, 908, 215, 44, 100, 666, 1023]. order-sorted [245, 908]. O’Regan [1004]. Oriented [391, 19, 83, 252, 937, 784, 797, 638, 116, 595, 532, 249, 21, 455, 589]. orthogonal [550]. Orwellian [218]. OSEK [1003]. OSEK/VDX [1003]. other [867]. Output [376, 851]. outputs [588]. Overtaking [336]. Overview [382]. Oxford [934]. OZ [350].

Packetised [298]. pAMN [516].

Paperback [966, 934]. Papers [387, 513]. paradigm [517, 225]. Parallel [413, 404, 19, 83, 606, 269, 89, 278, 154]. Parallelization [216]. Parameterization [1023]. Parameterized [951, 329, 619]. Parametric [170, 528, 231, 229]. Parametrized [281]. Paramorphisms [85]. Park [36]. Parlog [10]. part [588, 780, 725, 327, 328]. Partial [26, 893, 306, 117, 666, 273, 505, 150, 780, 862, 10, 921, 896, 100, 758]. Partial-order [306]. Partiality [284]. partition [1044]. partitioned [670]. partitioning [497]. partitions [109]. pass [506]. Passing [404, 318, 64, 130, 237]. patching [972]. path [904]. pattern [746, 544, 908, 25]. patterns [798, 605, 990, 887, 834]. Peirce [151]. PEPA [525]. Performance [298, 418, 470, 525, 467, 253, 114]. Periodic [336]. permission [955]. permissions [829]. Permutation [314]. persistent [1039]. Personal [389, 995]. Perspective [998, 669, 831]. pervasive [789, 844]. Peter [985, 604]. Petri [883, 403, 26, 565, 846, 576, 970, 752, 272, 876]. Phase [418, 809, 422]. physical [663, 1005]. physical-layer [663]. Physigrams [615]. pi [754]. pi-calculus [754]. piecewise [1038]. PIF [262]. Place [311]. planning [478]. platform [703]. Platzer [1005]. Plausibility [363]. PLC [357]. plugins [545]. Pnueli [644, 238]. Point [361, 356, 354, 357, 358, 50]. pointers [640]. policy [679, 602]. poly [675]. poly-threading [675]. polyhedra [488]. polymorphic [245]. polymorphism [231]. polynomial [581]. POP [390]. Portrait [410]. Pose [1016]. position [1020]. positioning [1054]. possibilistic [771]. potency [892]. Power [288, 1023, 485, 875, 191, 482]. pp [985, 966, 952, 934, 965]. Practical [352, 1012, 578, 818]. Practice [441, 595, 853]. Practice-oriented [595]. practitioner [210]. practitioners [966]. praise [707]. pre [1038]. pre-filter [1038]. precise [869, 1015]. precondition [176]. preconditions [75]. Predicate [338, 546, 191, 265, 32, 621]. predicates [680]. predictive [685]. Preface [728, 706]. prefix [148]. prefixes [148]. prehistory [995]. PREMO [302]. preorders [723]. presence [640, 242]. preservation [705]. preserved [110]. Preserving [635, 565, 979, 929, 834]. preset [805]. Presheaves [379]. Press [934]. Presupposition [175]. prevention [989]. Principle [172, 63]. principles [861, 687]. Priority [247, 160]. ProB [862, 682]. Probabilistic [1051, 724, 312, 421, 720, 1042, 937, 723, 812, 853, 938, 470, 725, 722, 771, 528, 1035, 711, 617, 119, 969, 485, 733]. probabilities [724, 160]. Probability [370, 669, 249, 247]. Problem [354, 956, 39]. Procedure [508, 480, 684, 659]. Procedures [338, 365, 369, 38, 134, 202, 965]. procedures.org/ [965]. Process [142, 12, 611, 370, 149, 52, 132, 148, 227, 254, 648, 131, 529, 484, 237, 563, 623, 1025, 635]. Processes [304, 281, 404, 160, 1023, 627, 143, 133, 64, 130, 259, 467, 569, 609, 28, 279, 81, 250, 523, 948, 1040, 509]. processors [154]. producing [1049, 543]. Product [1062, 790]. product-line [790]. productivity [909]. ProFeat [937]. Professor [927, 817, 36]. profiles [875]. Program [342, 102, 545, 9, 430, 836, 284, 280, 861, 505, 642, 870, 912, 691, 974, 23, 715, 264, 597]. Programming [413, 390, 1060, 895, 594, 496, 536, 515, 732, 941, 797, 124, 16, 718, 82, 116, 906, 661, 708, 907, 670, 1065, 825]. Programs [166, 442, 431, 347, 392, 331, 400, 174, 745, 776, 544, 911, 807, 562, 786, 897, 242, 275, 269, 621, 532, 34, 1022, 87, 208, 617, 215, 277, 267, 97, 907, 634, 685, 852, 659, 260, 968]. Progress [182, 678, 265, 562, 202, 211, 75]. progress-based [562]. projection [34, 919]. Prolog [174]. PROMELA [840, 464, 889]. Promotion [315]. prone [810]. Proof [38, 305, 330, 322, 1050, 237, 134, 324, 822, 906, 543, 192, 113, 58, 276, 643, 239, 793, 870, 478, 792, 560, 656, 64, 63, 262, 474, 269, 972, 461, 103, 960, 86, 976]. Proof-based [822]. ProofNavigator [597]. ProofPower [734]. ProofPower-Z [734]. Proofs [959, 284, 400, 624, 104, 958, 961, 275, 972, 1058, 715]. Proofs-as-Programs [400]. Propagation [1033, 911]. Properties [166, 421, 565, 812, 31, 822, 1013, 129]. Property [574, 840, 705, 265, 503, 677, 682, 834]. Property-directed [574]. Protection [1003]. Protective [283]. Protocol [310, 180, 308, 307, 417, 769, 493, 564, 228, 140, 674, 557, 139, 770, 521, 181, 416, 418, 421, 415, 419, 423]. protocols [618, 571, 663, 41, 490, 767, 683, 619, 749, 641, 945]. Prototypes [531]. Prototyping [168, 803]. provable [819]. Provably [506, 801]. Provably-correct [506]. prove [713]. Proved [416]. Prover [323, 753, 997, 97, 113]. prover-based [753]. Proving [744, 654, 897, 326, 894, 764, 17, 597]. PSL [447]. PSL/Sugar [447]. Ptolemy [466]. Publication [451]. pump [802]. Pure [291, 997]. purity [632]. purpose [656]. purse [560, 556, 555]. purses [557]. PuRSUE [1009]. push [678]. push-button [678]. pushdown [824]. Putput [991, 1006]. PVS [493, 764, 948, 521].

Quantified [835]. Quantitative [1043, 943]. Quantities [408]. quantum [980, 968]. Quasi [373]. Quest [359].

Railway [299, 300, 1055, 660, 1049]. RAISE [1021, 69, 560, 4, 743]. Ramifications [471]. randomized [710]. Raschke [966]. rational [727]. RC [502]. RC/ [502]. Reachability [575]. reaction [754]. reactions [468]. Reactive [331, 58, 702, 233, 750, 751, 881, 786, 234]. Read [989, 703]. Ready [871]. Real [52, 132, 368, 306, 348, 238, 384, 439, 529, 133, 459, 787, 786, 261, 764, 583, 177, 546, 208, 711, 837, 623, 232]. Real-Time [368, 348, 384, 439, 306, 238, 529, 133, 459, 787, 786, 261, 583, 177, 546, 208, 837, 623, 232].

Reasoning [149, 927, 174, 289, 364, 786, 340, 447, 170, 281, 633, 758, 172, 815, 459, 830, 479, 957, 159, 974, 836, 28, 619, 632, 25, 137, 710, 811]. Reconciling [711]. reconfiguration [901]. reconfigurations [849]. reconsidered [117]. Records [399]. recovery [809, 701]. Recursion [176, 133, 369, 23, 13]. Recursive [338, 846, 134, 202, 257, 519]. Redesign [445]. reduced [846]. Reduction [306, 763, 862, 918, 581, 647]. redundancy [971]. refactoring [979]. Refine [513]. Refinement [440, 33, 343, 367, 433, 334, 496, 245, 438, 614, 406, 294, 435, 51, 434, 436, 428, 413, 510, 798, 483, 823, 315, 617, 779, 249, 429, 437, 365, 439, 233, 480, 588, 777, 663, 962, 484, 807, 705, 780, 517, 828, 640, 125, 886, 13, 12, 713, 89, 184, 979, 712, 508, 452, 569, 711, 683, 737, 215, 455, 690, 556, 570, 482, 689, 775, 238, 196, 928, 190, 1056, 225, 1044, 692, 589, 726]. Refinement-based [779, 484, 1056]. Refinement-oriented [249]. refinement-testing [737]. refinements [618, 535, 526, 538, 548, 514, 591]. Refining [345, 285, 900, 369]. refusals [569]. refusals-based [569]. refutation [740]. region [949]. regular [462]. reification [665, 21]. relabelling [229]. Relating [353, 735, 928]. relation [805]. Relational [588, 436, 780, 895, 808]. relations [904, 931, 780, 981, 259]. Relationship [171, 505]. relaxation [454]. relaxed [939]. reliability [1002, 971, 159]. rely [648, 890, 981, 260, 289]. rely-guarantee [648, 890, 260]. Rely/ [289]. rely/guarantee [981]. Remote [365, 162]. Reo [695]. replicated [259]. reporting [1020]. repository [507]. representation [571]. Representational [364]. repudiation [490]. requirements [850, 881, 694, 161, 123]. Researcher [662]. resolution [909, 740, 1015]. RESOLVE [678]. resources [162, 523]. Response [161].

Responsiveness [475, 541]. rest [157]. restorable [755]. restricted [724]. retaining [724]. Retargeting [228]. Retrenchment [334, 655, 746, 750, 751, 496]. reuse [792]. Reversible [879, 733]. Review [985, 966, 1004, 952, 934, 1005, 986, 992, 965, 500, 72, 95, 243, 135]. Reviews [296, 90, 179, 230, 206, 15, 126, 29, 146, 236, 212, 6, 194, 54, 20, 47, 120, 200, 187]. Revising [327, 328]. Revisited [180, 290, 933, 81, 203, 181]. revivals [541]. rewrite [185, 240]. Rewriting [417, 575, 918, 17]. Reynolds [782]. Richer [246]. Right [287]. Rigorous [887, 609]. rippling [972]. RISC [597]. risk [1046]. risk-aware [1046]. RiskStructures [1046]. Ritchie [36]. river [755]. Robin [934, 637]. Robotic [1010, 1009]. robust [1018]. Robustness [1033]. Rod [388, 389, 387]. Roderick [986]. Rodin [655]. Rodrigues [1016]. ROLA [989]. role [877, 605]. Root [421, 423]. RT [382]. RT-Z [382]. rTiMo [963]. RTS [785, 810]. RTS/CTS [785, 810]. rule [125, 31, 156]. Rules [330, 894, 134, 269, 834]. run [873]. run-time [873]. Running [267]. Runtime [1013, 925]. Rüthing [992].

SACM [1050]. Safe [770, 1034]. Safety [1026, 202, 158, 1030, 161, 123, 1031, 730]. safety-critical [1031, 730]. salience [612]. SAN [525]. sanity [850]. SAT [982, 1049]. satisfiability [942, 608]. Scalable [698, 592]. scale [682]. scenarios [559]. schedulability [208]. schedulers [724, 722]. scheduling [978, 776]. Schema [430, 1060]. Schema-Based [430]. schematics [1049]. scheme [727, 810]. SCI [333]. Science [387]. scientist [604]. SCOOP [551, 602]. screw [964]. SCTL [380]. SCTL-MUS [380]. SDL [45]. SDLV [1030]. second [231]. section [959]. sector [703]. Secure [406, 878, 727]. Security [490, 731, 769, 955, 679, 840, 491, 557, 771, 683, 103, 619, 770, 945]. Segall [262]. selection [650]. Self [878, 1030, 720, 727]. Self-adaptation [878]. Self-Driving [1030]. self-interest [727]. self-stabilisation [720]. Semantic [732, 385, 252, 924, 843, 205]. Semantics [330, 343, 346, 446, 442, 447, 952, 341, 165, 221, 1035, 287, 280, 1012, 83, 62, 509, 628, 112, 658, 256, 829, 784, 941, 492, 909, 464, 116, 327, 606, 278, 103, 661, 1047, 586, 775, 948, 100, 666, 197, 834, 1025, 799, 814]. semantics-based [62]. semantics.org/ [952]. semaphores [762]. Semiotic [323]. semiring [718]. sensor [1064, 812]. separation [949, 648, 204, 506, 976]. sequence [735]. sequencer [271]. sequences [744, 805, 645, 504]. Sequential [509, 431, 89, 172, 148, 532, 214]. Sequentiality [167]. Sere [773]. Serial [418]. Sernadas [898]. server [879]. Service [356, 357, 358, 797, 671, 491, 695, 833, 803]. service-centred [797]. service-oriented [797]. services [674, 1014, 758]. Set [697, 9, 1038, 653]. Sets [374]. SH [301]. SH-Verification [301]. Shallow [396]. shaper [947]. Shared [413, 717, 760, 830, 974, 81, 222, 948, 260]. shared-memory [974]. Shared-Variable [413, 760]. Sharing [169, 785]. Sharp [334]. Should [391]. showing [211]. SI [271]. Signalling [307, 1056]. signature [527]. signing [794]. Simple [759, 778, 201, 568]. Simpson [471]. Simulation [334, 931, 501, 12, 195, 214]. simulations [534]. Simulink [607]. single [125]. singleton [509, 570]. Sintzoff [662]. Sketch [333]. slice [696]. slices [939]. Slicing [242, 581, 696, 505]. Sliding [180, 493, 181]. Sliding-Window [180, 181]. slot [471]. small [71]. smart [1002]. smells [631]. SMV [419]. Snepscheut [147, 180]. Social [323]. softcover [985]. Software [380, 386, 220, 869, 507, 966, 1029, 498, 598, 694, 497, 703, 831, 563, 21, 701, 926]. solution [497, 755]. Solutions [639, 273]. solving [982, 956]. Some [198, 479]. sorted [245, 908, 44]. sorting [913]. sound [1039, 939, 929, 945, 692]. Soundness [667]. Space [295, 132, 565, 98]. spaces [747]. Spanning [313]. Spark [1001]. spatial [736]. Special [959, 513, 530]. Specific [305]. Specification [239, 316, 333, 560, 378, 204, 288, 168, 532, 569, 468, 225, 417, 58, 1009, 92, 45, 16, 77, 69, 32, 518, 71, 161, 609, 631, 123, 268, 21, 157, 556, 641, 264, 623, 253]. Specifications [393, 363, 368, 345, 383, 386, 283, 379, 350, 1045, 92, 842, 526, 538, 548, 248, 199, 13, 848, 650, 581, 5, 68, 218, 213, 743, 854, 855]. Specified [428, 69]. Specifying [361, 8, 285, 162, 756]. Spectrum [390, 924]. SPIN [418, 840, 339, 486, 818]. Sprevak [934]. Springer [985, 966, 952, 965]. SRML [627]. stabilisation [720]. Stability [167]. stable [541]. stack [82, 592, 122]. stacks [674]. Standard [292, 332]. Standardization [311]. standards [871]. star [1038]. Starvation [762]. Starvation-free [762]. state [883, 565, 860, 946, 248, 921, 969, 783, 157, 98, 59, 196, 520, 519, 799]. state-based [157, 196]. Statechart [339]. Statechartable [752]. Stateflow [779]. Statemate [446]. statements [278]. states [93, 498]. static [144]. Statistical [1054, 1061, 975]. stays [157]. steady [969]. steady-state [969]. steered [1029]. Steering [1030]. Steffen [992]. step [507, 356]. Stepwise [1021, 712, 776]. Stochastic [1055, 920, 785, 917, 467, 711, 811]. stop [522]. stopping [778, 522]. storage [703]. storing [161, 123]. StPowla [627]. strand [747]. strategic [549]. strategies [719]. Strategy [435, 829, 392]. Stream [376, 375, 378, 373, 374, 402, 797, 582, 473, 848].

Stream-Based [378, 797]. Streamlining [562]. Strichman [965]. strict [257]. string [893, 759]. strong [59]. Structural [428, 837, 849, 255]. structure [26]. Structured [276, 643, 377, 222, 842, 5, 843, 223]. structures [149, 717, 544, 1039, 828, 979]. structuring [68]. studies [944]. Studio [362]. Study [380, 299, 356, 316, 415, 423, 307, 624, 1064, 527, 766, 70, 802, 516, 190]. Stuttering [426]. style [643, 922, 65, 852]. stylized [869]. subscripted [156]. Subset [339]. substation [1002]. Substitutability [343, 214]. substitution [577, 156, 53]. subtyping [231, 214, 255]. Sugar [447]. suitable [977]. suites [851]. summation [477]. super [39]. super-Turing [39]. Superposition [233, 452]. support [480, 540, 828, 596, 692]. survey [769]. Switched [1032]. switching [668]. symbiosis [944]. Symbolic [370, 685, 659, 1033, 1041, 1029]. symmetry [647]. Synchronisable [302]. Synchronisation [310, 454, 670]. synchronization [542, 991, 1006]. synchronizing [805]. synchronous [974, 497, 468]. Syntax [397, 82]. Synthesis [403, 1018, 869, 1009, 564, 1042, 794, 1008, 213, 543]. Synthesizing [915, 809, 849]. System [433, 321, 363, 300, 874, 955, 778, 1054, 846, 131, 558, 456, 870, 792, 77, 1021, 64, 688, 71, 789, 258, 528, 466, 942, 813, 822, 704, 98, 97, 222, 86, 591, 1056, 1011, 814]. system-level [814]. systematic [105]. Systems [380, 445, 406, 368, 377, 428, 408, 375, 386, 378, 170, 402, 1032, 301, 336, 171, 1059, 920, 349, 58, 835, 220, 702, 982, 742, 233, 750, 751, 850, 901, 535, 584, 679, 529, 183, 881, 1042, 680, 951, 1061, 182, 873, 459, 787, 16, 479, 261, 1008, 938, 248, 583, 470, 660, 32, 237, 1063, 476, 722, 731, 903, 882, 546, 849, 889, 528, 185, 452, 195, 813, 622, 214, 876, 768, 929, 837, 887, 851, 1053, 839, 755, 1010, 701, 240, 1027, 735, 623, 885, 196, 579, 611, 740, 2, 1025, 1031, 232, 985, 1005]. systolic [74, 34].

tableau [684]. tableau-based [684]. Tables [344, 251]. Tabular [437]. Tactic [429, 241]. Tactics [326, 692]. tale [1064]. Tank [516]. Tarski [290]. task [666]. tasks [815]. tccp [912]. TCOZ [385]. TCSP [385]. teaching [594, 643, 598, 944]. technique [936]. Techniques [306, 350, 763, 958, 525, 713, 195, 506]. teleo [786]. teleo-reactive [786]. Teller [362]. template [808]. Temporal [345, 705, 77, 232, 458, 654, 79, 526, 538, 548, 32, 886, 461, 569, 919, 213, 238, 158, 821, 596]. Temporal-logic [705]. Term [17, 575]. Termination [185, 744, 716, 606, 659]. Terms [105, 576]. Test [793, 374, 783, 651, 946, 722, 851]. Test-data [793]. Testing [440, 702, 500, 107, 304, 373, 582, 402, 755, 351, 587, 753, 723, 871, 724, 938, 487, 503, 903, 473, 848, 858, 703, 737, 960, 501, 834, 504, 502, 673]. Tests [368, 961, 650, 959]. Textbook [287]. TFA [1015]. Their [512, 431, 618, 535, 553, 444, 396, 622]. Theorem [323, 753, 764, 997, 17, 235]. theorems [495]. theoretic [124, 641]. Theoretical [1012]. Theories [275, 395, 515, 732, 77, 533, 734, 608]. Theory [401, 1004, 441, 304, 399, 3, 50, 62, 143, 853, 57, 9, 664, 833, 774, 218, 929, 240, 790, 964, 472]. Thomas [986]. Thread [549, 675]. Threaded [749, 745, 929]. threading [675]. Three [721, 398]. TIC [118]. Time [368, 787, 295, 312, 293, 443, 348, 384, 439, 845, 58, 52, 227, 1001, 809, 522, 529, 829, 306, 133, 873, 459, 786, 261, 79, 460, 583, 159, 917, 177, 546, 208, 711, 837, 238, 623, 160, 232]. Time-budgeting [787]. Timed [368, 925, 469, 845, 1018, 565, 881, 484, 492, 739, 461, 244, 118, 1013, 729]. timetable [956]. Timing [1003, 802]. TK [9]. to-alway [265]. Tobias [952]. together [487]. tokens [721]. tolerance [931, 87]. tolerant [104]. Tool [301, 596, 540, 828, 1049, 847]. Toolbox [8]. tools [861, 542, 598, 506, 4]. topological [1058]. TOrPEDO [1058]. Tournaments [923]. Trace [592, 150, 928]. Trace-based [592]. traces [183, 780, 503]. traffic [915]. train [778]. traits [792]. tramway [1054]. transactional [756]. transactions [557, 989]. transducers [244]. Transformation [386, 341, 87, 23, 816, 213, 940, 834, 629]. transformation-based [213, 629]. transformational [25]. Transformations [251, 858, 153, 837, 715, 589, 936]. transformer [265]. Transformers [338, 191]. Transforming [910]. transition [846, 32, 731, 528, 942, 851, 799]. Translating [646]. translation [652, 576, 154]. translation-based [652]. Treatment [289, 480, 81]. Tree [416, 418, 313, 415, 422, 419, 308, 417, 950, 1044]. Trees [697, 750, 751, 918]. Tribute [662]. Triple [1026, 1060]. trusted [2]. trustworthy [757]. Turing [934, 39]. Turski [796]. Two [259, 1064, 309, 760, 958, 519]. Type [376, 169, 255, 399, 122, 3, 239, 57, 736, 466, 235, 472]. Typed [844, 399, 245]. Types [304, 88, 427, 174, 496, 245, 246]. typing [491].

UI [325]. UK [934]. UML

[883, 456, 360, 559, 339, 563]. Un- [490]. unbounded [50, 183]. uncertainty [1018]. unchanged [157]. under-determined [221]. underspecification [735]. Unification [280, 190, 733]. Unified [434, 1050, 519, 643, 642, 848]. Unifying [949, 304, 461, 734, 515, 732]. unique [273]. Units [341, 77, 199]. UNITY [201, 1024, 156, 53]. universe [110]. University [934, 956]. unprivileged [807]. Untanglings [839]. Update [450]. updates [989]. Uppaal [1054, 310, 739, 521]. upper [584]. Use [406, 377, 271, 139, 783, 666, 559]. UseCase [655]. UseCase-wise [655]. User [322, 316, 323, 325, 613, 614, 531, 807]. Using [1001, 310, 356, 418, 815, 357, 289, 447, 386, 199, 722, 713, 282, 355, 339, 485, 268, 701, 653, 626, 883, 982, 1018, 613, 536, 317, 228, 41, 300, 913, 840, 862, 1021, 560, 248, 691, 886, 1003, 882, 889, 621, 595, 697, 876, 768, 887, 1010, 1013, 329, 514, 758, 225, 806, 854, 855, 14, 801, 1011, 799]. UTP [587, 1047, 586, 1012, 948, 963]. UTP-CSP [1047].

Valediction [580]. Validate [355]. Validating [447, 491, 307, 535, 607, 559]. Validation [360, 359, 307, 350, 1053, 956, 623]. Value [318, 64, 130, 215]. Value-Passing [318, 64, 130]. Variability [940, 1061, 977]. Variability-based [940]. variability-intensive [1061]. Variable [413, 397, 760, 474, 260]. Variables [338, 335, 439, 81, 471, 948]. variants [519]. Variations [395, 790]. Variety [439]. VDM [186, 70, 595, 5, 68, 661]. VDMTools [355]. VDX [1003]. Veith [986]. Verhoef [985]. verifiable [537]. Verification [493, 542, 299, 310, 356, 680, 300, 364, 302, 557, 1003, 168, 621, 404, 339, 208, 813, 301, 1010, 612, 437, 329, 423, 1038, 854, 855, 417, 920, 1030, 1033, 74, 763, 982, 145, 861, 769, 480, 1001, 745, 980, 536, 679, 663, 529, 544, 498, 951, 545, 552, 791, 16, 1043, 598, 871, 930, 477, 487, 691, 660, 32, 923, 882, 546, 272, 789, 720, 1014, 816, 532, 939, 1022, 682, 822, 990, 609, 988, 779, 91, 944, 802, 97, 641, 597, 579, 634, 852, 968, 932, 943, 232, 936, 301]. verification-driven [988]. Verification-guided [612]. Verifications [282, 331, 254]. Verified [313, 279, 507, 954, 152]. verifier [678]. Verifying [527, 531, 792, 1039, 357, 258, 768, 704, 277, 419, 514, 507, 915, 756, 195, 139, 260]. Verlag [952, 965]. version [916, 241, 1040]. Versions [280]. versus [22, 771, 569]. VHDL [456]. via [750, 751, 672, 665, 677, 471, 1060, 1033]. view [144, 995]. views [630]. Virtual [1055, 222, 1015]. visual [629]. visualizing [754]. VLSI [114]. Volume [448, 992]. voting [727, 768]. vs [917, 452]. VSTTE [530]. VVSL [5, 68].

Wait [420]. WAM [245, 239]. Want [420]. Warner [566]. Warshall [709]. Way [377]. weak [867, 853, 1000]. Weakest [75, 176]. Weakness [1019]. web [491, 833, 758, 843]. weighted [1044]. Where [441]. White [863]. who [604]. wide [924]. wide-spectrum [924]. Wilson [934]. Window [180, 493, 181]. Winskel [287]. wireless [1064, 812, 932]. wise [866, 655]. within [733]. witnessing [1058]. Władysław [796]. workbench [941]. workflow [667]. working [487]. World [349]. worst [845]. wrappers [116]. write [163]. www.concrete-semantics.org [952]. www.decision-procedures.org [965]. Wybe [410].

x [952, 376, 500, 377, 375, 373, 582, 372, 374, 402, 473, 848, 39]. X-Machine [377, 373, 582, 402, 473, 848]. X-Machines [376, 375, 372, 374, 500, 39]. xSAP [1026]. xv [934]. xxi [966, 965]. xxii [985].

years [999]. Yonezawa [286]. yourself [3].

Z [294, 285, 383, 428, 652, 705, 561, 70, 292, 327, 328, 518, 162, 332, 385, 353, 734, 197, 384, 235, 246, 382]. Z/Eves [561]. Z2SAL [652]. Zemanek [817]. Zeno [626]. Zohar [993]. zonotope [1038]. ZRC [294].

References

Jones:1989:E

[1] C. B. Jones. Editorial. Formal Aspects of Computing, 1(1):1–3, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01887196.

Thomas:1989:DMT

[2] Martyn Thomas. Development methods for trusted computer systems. Formal Aspects of Computing, 1(1):5–18, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01887197.

Backhouse:1989:DIY

[3] Roland Backhouse, Paul Chisholm, Grant Malcolm, and Erik Saaman. Doit-yourself type theory. Formal Aspects of Computing, 1(1):19–84, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01887198.

Nielsen:1989:RLM

[4] Mogens Nielsen, Klaus Havelund, Kim Ritter Wagner, and Chris George. The RAISE language, method and tools. Formal Aspects of Computing, 1(1):85–114, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887199.

Middelburg:1989:VLS

[5] C. A. Middelburg. VVSL: A language for structured VDM specifications. Formal Aspects of Computing, 1(1):115–135, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887200.

Monahan:1989:BR

[6] Brian Monahan. Book reviews. Formal Aspects of Computing, 1(1):137–144, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01887201.

Cooke:1989:E

[7] D. J. Cooke. Editorial. Formal Aspects of Computing, 1(1):145–146, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01887202;   http://link.springer.
com/article/10.1007/BF01887202.

Burton:1989:SAM

[8] C. T. Burton, S. J. Cook, S. Gikas, J. R. Rowson, and S. T. Sommerville. Specifying the Apple Macintosh^{T M} Toolbox event manager. Formal Aspects of Computing, 1(1):147–171, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01887203.

Henson:1989:PDC

[9] Martin C. Henson. Program development in the constructive set theory TK. Formal Aspects of Computing, 1(1):173–192, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887204.

Huntbach:1989:MIP

[10] Matthew Huntbach. Meta-interpreters and partial evaluation in Parlog. Formal Aspects of Computing, 1(1): 193–211, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887205.

Kwiatkowska:1989:EFN

[11] Marta Z. Kwiatkowska. Event fairness and non-interleaving concurrency. Formal Aspects of Computing, 1(1): 213–228, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887206.

Jifeng:1989:PSR

[12] He Jifeng. Process simulation and refinement. Formal Aspects of Computing, 1(1):229–241, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/BF01887207.

Holmstrom:1989:RCS

[13] Sören Holmström. A refinement calculus for specifications in Hennessy– Milner logic with recursion. Formal Aspects of Computing, 1(1):242–272, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01887208.

Walker:1989:AAM

[14] D. J. Walker. Automated analysis of mutual exclusion algorithms using CCS. Formal Aspects of Computing, 1(1):273–292, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887209.

Cooke:1989:BR

[15] John Cooke, Crustian Calude, Will Harwood, and Dan Simpson. Book reviews. Formal Aspects of Computing, 1(1):293–301, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887210.

Dybjer:1989:FPA

[16] Peter Dybjer and Herbert P. Sander. A functional programming approach to the specification and verification of concurrent systems. Formal Aspects of Computing, 1(1):303–319, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01887211.

Nipkow:1989:TRB

[17] Tobias Nipkow. Term rewriting and beyond — theorem proving in Isabelle. Formal Aspects of Computing, 1(1):320–338, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887212.

Thompson:1989:LM

[18] Simon Thompson. A logic for Miranda. Formal Aspects of Computing, 1(1):339–365, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887213.

America:1989:IDP

[19] Pierre America. Issues in the design of a parallel object-oriented language. Formal Aspects of Computing, 1(1):366–411, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887214.

Parke:1989:BR

[20] Tom Parke and R. McLean. Book reviews. Formal Aspects of Computing, 1(1):412–415, March 1989. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01887215.

Oliveira:1990:RCM

[21] J. N. Oliveira. A reification calculus for model-oriented software specification. Formal Aspects of Computing, 2(1):1–23, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01888215.

Carmo:1990:BVL

[22] José Carmo and Amílcar Sernadas. Branching versus linear logics yet again. Formal Aspects of Computing, 2(1):24–59, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01888216.

Hesselink:1990:CAR

[23] Wim H. Hesselink. Command algebras, recursion and program transformation. Formal Aspects of Computing, 2(1):60–104, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01888217.

Anonymous:1990:FEa

[24] Anonymous. Forthcoming events 1990. Formal Aspects of Computing, 2(1):105–108, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01888218.

Partsch:1990:FPM

[25] H. A. Partsch and F. A. Stomp. A fast pattern matching algorithm derived by transformational and assertional reasoning. Formal Aspects of Computing, 2(1):109–122, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01888219.

Best:1990:POB

[26] Eike Best and Jörg Desel. Partial order behaviour and structure of Petri nets. Formal Aspects of Computing, 2(1):123–138, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01888220.

Hesselink:1990:AML

[27] Wim H. Hesselink. Axioms and models of linear logic. Formal Aspects of Computing, 2(1):139–166, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01888221.

Misra:1990:ERA

[28] Jayadev Misra. Equational reasoning about nondeterministic processes. Formal Aspects of Computing, 2(1):167–195, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01888222.

Denvir:1990:BR

[29] Tim Denvir, Rosamund Rawlings, Tom Parke, S. J. Goldsack, and Anthony Hall. Book reviews. Formal Aspects of Computing, 2(1):196–202, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01888223.

Anonymous:1990:FEb

[30] Anonymous. Forthcoming events 1990. Formal Aspects of Computing, 2(1):203–205, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01888224.

Hermann:1990:CPR

[31] Miki Hermann. Chain properties of rule closures. Formal Aspects of Computing, 2(1):207–225, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01888225.

He:1990:IPT

[32] Xudong He and John A. N. Lee. Integrating predicate transition nets with first order temporal logic in the specification and verification of concurrent systems. Formal Aspects of Computing, 2(1):226–246, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01888226.

Back:1990:RCF

[33] R. J. R. Back and J. von Wright. Refinement concepts formalised in higher order logic. Formal Aspects of Computing, 2(1):247–272, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01888227.

Lengauer:1990:PSP

[34] C. Lengauer and J. W. Sanders. The projection of systolic programs. Formal Aspects of Computing, 2(1):273–293, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01888228.

Anonymous:1990:FEc

[35] Anonymous. Forthcoming events 1990. Formal Aspects of Computing, 2(1):294–298, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01888229.

Paterson:1990:OPD

[36] M. S. Paterson. Obituary: Professor David Michael Ritchie Park. Formal Aspects of Computing, 2(1):299–300, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01888230;   http://link.springer.
com/article/10.1007/BF01888230.

Lin:1990:MMI

[37] Huimin Lin and Man-Chi Pong. Modelling multiple inheritance with colimits. Formal Aspects of Computing, 2(1):301–311, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01888231.

Ah-kee:1990:POB

[38] Alain Ah-kee. Proof obligations for blocks and procedures. Formal Aspects of Computing, 2(1):312–330, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01888232.

Stannett:1990:XMH

[39] Mike Stannett. X-machines and the halting problem: Building a super-Turing machine. Formal Aspects of Computing, 2(1):331–341, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01888233.

Soparkar:1990:ICH

[40] Nandit Soparkar and Abraham Silberschatz. On the interconnection constants of Hopfield nets. Formal Aspects of Computing, 2(1):342–358, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01888234.

Chapman:1990:DAI

[41] N. P. Chapman. Defining, analysing and implementing communication protocols using attribute grammars. Formal Aspects of Computing, 2(1): 359–392, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01888235.

Anonymous:1990:FEd

[42] Anonymous. Forthcoming events 1990. Formal Aspects of Computing, 2(1):393–396, March 1990. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01888236.

Gries:1991:E

[43] David Gries. Editorial. Formal Aspects of Computing, 3(1):1, January/ March 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01211432;   http://link.springer.
                                        com/article/10.1007/BF01211432.

Owe:1991:GIO

[44] Olaf Owe and Ole-Johan Dahl. Generator induction in order sorted algebras. Formal Aspects of Computing, 3(1): 2–20, January/March 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211433.

Broy:1991:TFF

[45] Manfred Broy. Towards a formal foundation of the specification and description language SDL. Formal Aspects of Computing, 3(1):21–57, January/ March 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211434.

Stepney:1991:DCC

[46] Susan Stepney, Dave Whitley, David Cooper, and Colin Grant. A demonstrably correct compiler. Formal Aspects of Computing, 3(1):58– 101, January/March 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211435.

Pitt:1991:BR

[47] David Pitt and Dan Simpson. Book reviews. Formal Aspects of Computing, 3(1):102–105, January/March 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211436.

Anonymous:1991:FEa

[48] Anonymous. Forthcoming events. Formal Aspects of Computing, 3(1):106– 108, January/March 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211437.

Jones:1991:E

[49] C. B. Jones. Editorial. Formal Aspects of Computing, 3(2):109, June 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01898398;   http://link.springer.
com/article/10.1007/BF01898398.

Barrett:1991:FPT

[50] Geoff Barrett. The fixed point theory of unbounded non-determinism. Formal Aspects of Computing, 3(2):110–128, June 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01898399.

Cusack:1991:RCI

[51] Elspeth Cusack. Refinement, conformance and inheritance. Formal Aspects of Computing, 3(2):129–141, June 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01898400.

Baeten:1991:RTP

[52] J. C. M. Baeten and J. A. Bergstra. Real time process algebra. Formal Aspects of Computing, 3(2):142–188, June 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01898401.

Sanders:1991:ESA

[53] Beverly A. Sanders. Eliminating the substitution axiom from UNITY logic. Formal Aspects of Computing, 3(2):189–205, June 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01898402.

Murphy:1991:BR

[54] David Murphy and J. W. Sanders. Book reviews. Formal Aspects of Computing, 3(2):206–212, June 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/BF01898403.

Anonymous:1991:FEb

[55] Anonymous. Forthcoming events. Formal Aspects of Computing, 3 (2):213–216, June 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01898404.

Denvir:1991:E

[56] Tim Denvir. Editorial. Formal Aspects of Computing, 3(3):217, July/ September 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF01245631;   http://link.
springer.com/article/10.1007/BF01245631.

Hedberg:1991:NAL

[57] Michael Hedberg. Normalising the associative law: An experiment with Martin-Löf’s type theory. Formal Aspects of Computing, 3(3):218– 252, July/September 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01245632.

Alagar:1991:FSP

[58] Vangalur S. Alagar and Geetha Ramanathan. Functional specification and proof of correctness for time dependent behaviour of reactive systems. Formal Aspects of Computing, 3(3):253–283, July/September 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01245633.

Qin:1991:FFS

[59] Huajun Qin and Philip Lewis. Factorisation of finite state machines under strong and observational equivalences. Formal Aspects of Computing, 3(3):284–307, July/September 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01245634.

Anonymous:1991:FEc

[60] Anonymous. Forthcoming events. Formal Aspects of Computing, 3(3):308– 312, July/September 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01245635.

Cooke:1991:E

[61] John Cooke. Editorial. Formal Aspects of Computing, 3(4):313–314, October/ December 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF01642505;   http://link.
springer.com/article/10.1007/BF01642505.

Baumann:1991:TSB

[62] Peter Baumann. Towards a semanticsbased information theory. Formal Aspects of Computing, 3(4):315–325, October/December 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01642506.

Hennicker:1991:CIP

[63] Rolf Hennicker. Context induction: A proof principle for behavioural abstractions and algebraic implementations. Formal Aspects of Computing, 3(4):326–345, October/December 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01642507.

Hennessy:1991:PSC

[64] M. Hennessy. A proof system for communicating processes with valuepassing. Formal Aspects of Computing, 3(4):346–366, October/December 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01642508.

Notarmarco:1991:FAC

[65] Christiane Notarmarco and Rod Mulvey. Formal aspects of computing: L^{A}TEX style guide for authors. Formal Aspects of Computing, 3(4):367– 376, October/December 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01642509.

Anonymous:1991:FEd

[66] Anonymous. Forthcoming events. Formal Aspects of Computing, 3(4):377– 378, October/December 1991. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF01642510;   http:
                                        //link.springer.com/article/10.
                                        1007/BF01642510.

Denvir:1992:E

[67] Tim Denvir. Editorial. Formal Aspects of Computing, 4(1):1–12, January 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01214954.

Middelburg:1992:MSV

[68] C. A. Middelburg. Modular structuring of VDM specifications in VVSL. Formal Aspects of Computing, 4 (1):13–47, January 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01214955.

George:1992:NDS

[69] Chris George. The NDB database specified in the RAISE specification language. Formal Aspects of Computing, 4(1):48–75, January 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01214956.

Hayes:1992:VZC

[70] Ian Hayes. VDM and Z: A comparative case study. Formal Aspects of Computing, 4(1):76–99, January 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01214957.

Kazmierczak:1992:MSS

[71] Edmund Kazmierczak. Modularising the specification of a small database system in extended ML. Formal Aspects of Computing, 4(1):100–142, January 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01214958.

Henglein:1992:BR

[72] Fritz Henglein. Book review. Formal Aspects of Computing, 4(1):143– 144, January 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF01214959;   http://link.
springer.com/article/10.1007/BF01214959.

Anonymous:1992:FEa

[73] Anonymous. Forthcoming events. Formal Aspects of Computing, 4(1): 145–147, January 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214960.

Abdulla:1992:AVC

[74] Parosh Abdulla. Automatic verification of a class of systolic circuits. Formal Aspects of Computing, 4(2):149–194, March 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01212335.

Lukkien:1992:WPP

[75] Johan J. Lukkien and Jan L. A. van de Snepscheut. Weakest preconditions for progress. Formal Aspects of Computing, 4(2):195–236, March 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01212336.

Anonymous:1992:FEc

[76] Anonymous. Forthcoming events. Formal Aspects of Computing, 4(2):237– 238, March 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF01212337;   http://link.
springer.com/article/10.1007/BF01212337.

Fiadeiro:1992:TTM

[77] J. Fiadeiro and T. Maibaum. Temporal theories as modularisation units for concurrent system specification. Formal Aspects of Computing, 4 (3):239–272, May 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01212304.

Anderson:1992:CA

[78] James H. Anderson and Mohamed G. Gouda. A criterion for atomicity. Formal Aspects of Computing, 4(3):273–298, May 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01212305.

Fisher:1992:MCL

[79] Michael Fisher. A model checker for linear time temporal logic. Formal Aspects of Computing, 4(3):299–319, May 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01212306.

Anonymous:1992:FEd

[80] Anonymous. Forthcoming events. Formal Aspects of Computing, 4 (3):320–322, May 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01212307.

Owe:1992:ATP

[81] Olaf Owe. Axiomatic treatment of processes with shared variables revisited. Formal Aspects of Computing, 4(4):323–340, July 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211310.

Fassbender:1992:ISD

[82] Heinz Fassbender and Heiko Vogler. An implementation of syntax directed functional programming on nested-stack machines. Formal Aspects of Computing, 4(4):341–375, July 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211311.

America:1992:LSP

[83] Pierre America and Jan Rutten. A layered semantics for a parallel objectoriented language. Formal Aspects of Computing, 4(4):376–408, July 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211312.

Anonymous:1992:FEe

[84] Anonymous. Forthcoming events. Formal Aspects of Computing, 4 (4):409–411, July 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211313.

Meertens:1992:P

[85] Lambert Meertens. Paramorphisms. Formal Aspects of Computing, 4(5): 413–424, September 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211391.

Seldin:1992:CCC

[86] Jonathan P. Seldin. Coquand’s calculus of constructions: A mathematical foundation for a proof development system. Formal Aspects of Computing, 4(5):425–441, September 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211392.

Liu:1992:TPF

[87] Zhiming Liu and Mathai Joseph. Transformation of programs for faulttolerance. Formal Aspects of Computing, 4(5):442–469, September 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF01211393.

Hindley:1992:TII

[88] J. Roger Hindley. Types with intersection: An introduction. Formal Aspects of Computing, 4(5):470–486, September 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211394.

Kay:1992:SPB

[89] Andrew Kay and Peter Lupton. Sequential to parallel buffer refinement. Formal Aspects of Computing, 4(5): 487–492, September 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211395.

Atkinson:1992:BR

[90] M. D. Atkinson, David Lester, and Lawrence C. Paulson. Book reviews. Formal Aspects of Computing, 4(5): 493–496, September 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211396.

Orava:1992:AVM

[91] Fredrik Orava and Joachim Parrow. An algebraic verification of a mobile network. Formal Aspects of Computing, 4(6):497–543, November 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211473.

Boudriga:1992:LSA

[92] Noureddine Boudriga, Fathi Elloumi, and Ali Mili. On the lattice of specifications: Applications to a specification methodology. Formal Aspects of Computing, 4(6):544–571, November 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211474.

Best:1992:CGH

[93] Eike Best, Ludmila Cherkasova, and Jörg Desel. Compositional generation of home states in free choice nets. Formal Aspects of Computing, 4(6): 572–581, November 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211475.

Francez:1992:ECF

[94] Nissim Francez, Ralph-J. J. Back, and Reino Kurki-Suonio. On equivalencecompletions of fairness assumptions. Formal Aspects of Computing, 4(6): 582–591, November 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211476.

Hussmann:1992:BR

[95] Heinrich Hussmann. Book review. Formal Aspects of Computing, 4(6): 592–593, November 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF01211477;   http:
                                        //link.springer.com/article/10.
                                        1007/BF01211477.

Jones:1992:E

[96] Cliff Jones and John Cooke. Editorial. Formal Aspects of Computing, 4(1S):595–596, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF03180563;   http://link.springer.
com/article/10.1007/BF03180563.

Russinoff:1992:VSC

[97] David M. Russinoff. A verification system for concurrent programs based on the Boyer–Moore prover. Formal Aspects of Computing, 4(1S):597–611, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF03180564.

Pomello:1992:ACE

[98] L. Pomello and C. Simone. An algebraic characterisation of elementary net system (observable) state space. Formal Aspects of Computing, 4(1S):612–637, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03180565.

Balcazar:1992:DBC

[99] José Balcázar, Joaquim Gabarró, and Miklós Sántha. Deciding bisimilarity is P -complete. Formal Aspects of Computing, 4(1S):638–648, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF03180566.

Shields:1992:MHP

[100] M. W. Shields. Multitraces, hypertraces and partial order semantics. Formal Aspects of Computing, 4(1S):649–672, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF03180567.

Fokkinga:1992:CC

[101] Maarten M. Fokkinga. Calculate categorically! Formal Aspects of Computing, 4(1S):673–692, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF03180568.

Burton:1992:PM

[102] C. T. P. Burton. Program morphisms. Formal Aspects of Computing, 4(1S):693–726, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF03180569.

Mizuno:1992:SFC

[103] Masaaki Mizuno and David Schmidt. A security flow control algorithm and its denotational semantics correctness proof. Formal Aspects of Computing, 4(1S):727–754, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF03180570.

Bevier:1992:MCP

[104] William R. Bevier and William D. Young. Machine checked proofs of the design of a fault-tolerant circuit. Formal Aspects of Computing, 4(1S):755–775, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03180571.

Fokker:1992:SCO

[105] Jeroen Fokker. The systematic construction of a one-combinator basis for Lambda–terms. Formal Aspects of Computing, 4(1S):776–780, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF03180572.

Anonymous:1992:FEb

[106] Anonymous. Forthcoming events. Formal Aspects of Computing, 4 (1S):781–782, November/December 1992. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF03180573;   http://link.springer.
com/article/10.1007/BF03180573.

Cleaveland:1993:TEB

[107] Rance Cleaveland and Matthew Hennessy. Testing equivalence as a bisimulation equivalence. Formal Aspects of Computing, 5(1):1–20, January 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211314.

Lentfert:1993:DMM

[108] P. J. A. Lentfert and S. D. Swierstra. Distributed maximum maintenance on hierarchically divided graphs. Formal Aspects of Computing, 5(1):21–60, January 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211315.

Bird:1993:LP

[109] R. S. Bird and O. de Moor. List partitions. Formal Aspects of Computing, 5(1):61–78, January 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211316.

Massart:1993:EAE

[110] T. Massart and R. Devillers. Equality of agent expressions is preserved under an extension of the universe of actions. Formal Aspects of Computing, 5(1):79–88, January 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211317.

Anonymous:1993:FEa

[111] Anonymous. Forthcoming events. Formal Aspects of Computing, 5(1):89– 90, January 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF01211318;   http://link.
                                        springer.com/article/10.1007/BF01211318.

Bucci:1993:EIS

[112] A. Bucci, P. Inverardi, and S. Martini. An “executable” impredicative semantics for the Ada configuration. Formal Aspects of Computing, 5(2):91–120, March 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211301.

Weber-Wulff:1993:PMP

[113] Debora Weber-Wulff. Proof movie — a proof with the Boyer–Moore prover. Formal Aspects of Computing, 5(2):121–151, March 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211302.

Williams:1993:ADP

[114] Alan Williams. The applicability of discrete performance estimation methods to VLSI design. Formal Aspects of Computing, 5(2):152–176, March 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211303.

Anonymous:1993:FEb

[115] Anonymous. Forthcoming events. Formal Aspects of Computing, 5(2): 177–180, March 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211304.

Hense:1993:DSO

[116] Andreas V. Hense. Denotational semantics of an object-oriented programming language with explicit wrappers. Formal Aspects of Computing, 5(3):181–207, May 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211554.

Owe:1993:PLR

[117] Olaf Owe. Partial logics reconsidered: A conservative approach. Formal Aspects of Computing, 5(3):208–223, May 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211555.

Quemada:1993:TTC

[118] Juan Quemada, David de Frutos, and Arturo Azcorra. TIC: A timed calculus. Formal Aspects of Computing, 5(3):224–252, May 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211556.

Miguel:1993:LEP

[119] Carlos Miguel, Angel Fernández, and Leon Vidaller. LOTOS extended with probabilistic behaviours. Formal Aspects of Computing, 5(3):253–281, May 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211557.

Roy:1993:BR

[120] Dyckhoff Roy and E. Moggi. Book reviews. Formal Aspects of Computing, 5(3):282–284, May 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211558.

Anonymous:1993:FEc

[121] Anonymous. Forthcoming events. Formal Aspects of Computing, 5 (3):285–288, May 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211559.

Stoddart:1993:TIS

[122] Bill Stoddart and Peter J. Knaggs. Type inference in stack based languages. Formal Aspects of Computing, 5(4):289–298, July 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01212404.

Mukherjee:1993:FSS

[123] Paul Mukherjee and Victoria Stavridou. The formal specification of safety requirements for storing explosives. Formal Aspects of Computing, 5(4):299–336, July 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.

1007/BF01212405. See response [161].

Dennis-Jones:1993:CMC

[124] Esther Dennis-Jones and David E. Rydeheard. Categorical ML — category-theoretic modular programming. Formal Aspects of Computing, 5(4):337–366, July 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01212406.

Gardiner:1993:SCR

[125] P. H. B. Gardiner and Carroll Morgan. A single complete rule for data refinement. Formal Aspects of Computing, 5(4):367–382, July 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01212407.

Davie:1993:BR

[126] Tony Davie and Simon Brock. Book reviews. Formal Aspects of Computing, 5(4):383–384, July 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01212408;   http://link.springer.
                                        com/article/10.1007/BF01212408.

Anonymous:1993:FEd

[127] Anonymous. Forthcoming events. Formal Aspects of Computing, 5 (4):385–387, July 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01212409.

Jones:1993:E

[128] Cliff Jones and John Cooke. Editorial. Formal Aspects of Computing, 5 (5):389–390, September 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF01212484;   http:
                                        //link.springer.com/article/10.
                                        1007/BF01212484.

Weber:1993:DBP

[129] Matthias Weber. Definition and basic properties of the Deva meta-calculus. Formal Aspects of Computing, 5(5): 391–431, September 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01212485.

Hennessy:1993:CPV

[130] M. Hennessy and A. Ingólfsdóttir. Communicating processes with valuepassing and assignments. Formal Aspects of Computing, 5(5): 432–466, September 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01212486.

Brebner:1993:CBI

[131] Gordon Brebner. A CCS-based investigation of deadlock in a multi-process electronic mail system. Formal Aspects of Computing, 5(5):467–479, September 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01212487.

Baeten:1993:RSP

[132] J. C. M. Baeten and J. A. Bergstra. Real space process algebra. Formal Aspects of Computing, 5(6): 481–529, November 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211247.

Davies:1993:RIR

[133] Jim Davies and Steve Schneider. Recursion induction for real-time processes. Formal Aspects of Computing, 5 (6):530–553, November 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211248.

Hesselink:1993:PRR

[134] Wim H. Hesselink. Proof rules for recursive procedures. Formal Aspects of Computing, 5(6):554–570, November 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211249.

Lins:1993:BR

[135] Rafael D. Lins. Book review. Formal Aspects of Computing, 5(6):571–572, November 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF01211250;   http://link.
                                        springer.com/article/10.1007/BF01211250.

Anonymous:1993:FEe

[136] Anonymous. Forthcoming events. Formal Aspects of Computing, 5(6):573, November 1993. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF01211251;   http://link.
                                        springer.com/article/10.1007/BF01211251.

Staples:1994:FLH

[137] John Staples, Peter J. Robinson, and Daniel Hazel. A functional logic for higher level reasoning about computation. Formal Aspects of Computing, 6(1):1–38, January 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211079.

Nickolas:1994:CFL

[138] Peter Nickolas. The completeness of functional logic. Formal Aspects of Computing, 6(1):39–59, January 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211080.

Moore:1994:FMA

[139] J. Strother Moore. A formal model of asynchronous communication and its use in mechanically verifying a biphase mark protocol. Formal Aspects of Computing, 6(1):60–91, January 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211081.

Bruns:1994:FAC

[140] Glenn Bruns and Stuart Anderson. The formalization and analysis of a communications protocol. Formal Aspects of Computing, 6(1):92–112, January 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211082.

Jones:1994:E

[141] Cliff Jones. Editorial. Formal Aspects of Computing, 6(2):113–114, March 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01221096;   http://link.springer.
com/article/10.1007/BF01221096.

Groote:1994:PAG

[142] Jan Friso Groote and Alban Ponse. Process algebra with guards: Combining Hoare logic with process algebra. Formal Aspects of Computing, 6(2):115–164, March 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01221097.

Boudol:1994:TPL

[143] G. Boudol, I. Castellani, M. Hennessy, and A. Kiehn. A theory of processes with localities. Formal Aspects of Computing, 6(2):165–200, March 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01221098.

Aceto:1994:SVL

[144] Luca Aceto. A static view of localities. Formal Aspects of Computing, 6(2):201–222, March 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01221099.

Antoniou:1994:VM

[145] Grigoris Antoniou. The verification of modules. Formal Aspects of Computing, 6(2):223–244, March 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01221100.

Horning:1994:BR

[146] James J. Horning. Book reviews. Formal Aspects of Computing, 6(2):245– 246, March 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF01221101;   http://link.
                                        springer.com/article/10.1007/BF01221101.

Dijkstra:1994:JVS

[147] Edsger W. Dijkstra. Jan L. A. van de Snepscheut. Formal Aspects of Computing, 6(3):247–249, May 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01215406.

Baeten:1994:SCA

[148] J. C. M. Baeten and J. A. Bergstra. On sequential composition, action prefixes and process prefix. Formal Aspects of Computing, 6(3):250–268, May 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01215407.

America:1994:RAD

[149] Pierre America and Frank de Boer. Reasoning about dynamically evolving process structures. Formal Aspects of Computing, 6(3):269–316, May 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01215408.

Bloom:1994:WPT

[150] Bard Bloom. When is partial trace equivalence adequate? Formal Aspects of Computing, 6(3):317–338, May 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01215409.

Brink:1994:PA

[151] Chris Brink, Katarina Britz, and Renate A. Schmidt. Peirce algebras. Formal Aspects of Computing, 6(3):339–358, May 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01215410.

Russinoff:1994:MVI

[152] David M. Russinoff. A mechanically verified incremental garbage collector. Formal Aspects of Computing, 6(4):359–390, July 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211305.

Lieberherr:1994:OEC

[153] Karl J. Lieberherr, Walter L. Hürsch, and Cun Xiao. Object-extending class transformations. Formal Aspects of Computing, 6(4):391–416, July 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211306.

Wray:1994:CTD

[154] J. P. Wray and A. Stewart. Correct translation of data parallel assignment onto array processors. Formal Aspects of Computing, 6(4):417–439, July 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211307.

Dybjer:1994:IF

[155] Peter Dybjer. Inductive families. Formal Aspects of Computing, 6 (4):440–465, July 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211308.

Prasetya:1994:EUS

[156] I. S. W. B. Prasetya. Error in the UNITY substitution rule for subscripted operators. Formal Aspects of Computing, 6(4):466–470, July 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211309.

Pitt:1994:RSU

[157] David Pitt and Paddy Byers. The rest stays unchanged (concurrency and state-based specification). Formal Aspects of Computing, 6(5): 471–494, September 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211864.

Sistla:1994:SLF

[158] A. Prasad Sistla. Safety, liveness and fairness in temporal logic. Formal Aspects of Computing, 6(5): 495–511, September 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211865.

Hansson:1994:LRA

[159] Hans Hansson and Bengt Jonsson. A logic for reasoning about time and reliability. Formal Aspects of Computing, 6(5):512–535, September 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211866.

Tofts:1994:PPP

[160] Chris Tofts. Processes with probabilities, priority and time. Formal Aspects of Computing, 6(5):536–564, September 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211867.

Larsen:1994:RFS

[161] Peter Gorm Larsen. Response to “The formal specification of safety requirements for storing explosives”. Formal Aspects of Computing, 6(5): 565–568, September 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211868. See [123].

Houston:1994:SDC

[162] Iain S. C. Houston and Mark B. Josephs. Specifying distributed CICS in Z: Accessing local and remote resources. Formal Aspects of Computing, 6(5):569–579, September 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/BF01211869.

Lamport:1994:HWL

[163] Leslie Lamport. How to write a long formula. Formal Aspects of Computing, 6(5):580–584, September 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211870.

Cooke:1994:E

[164] John Cooke. Editorial. Formal Aspects of Computing, 6(6):585, December 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF03259386;   http://link.springer.
com/article/10.1007/BF03259386.

Larsen:1994:FSD

[165] Peter Gorm Larsen, Nico Plat, and Hans Toetenel. A formal semantics of data flow diagrams. Formal Aspects of Computing, 6(6):586–606, December 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03259387.

Chandy:1994:PCP

[166] K. Mani Chandy. Properties of concurrent programs. Formal Aspects of Computing, 6(6):607–619, December 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03259388.

Avron:1994:SSD

[167] Arnon Avron and Nada Sasson. Stability, sequentiality and demand driven evaluation in dataflow. Formal Aspects of Computing, 6(6):620–642, December 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03259389.

Jifeng:1994:SVP

[168] He Jifeng and Jonathan Bowen. Specification, verification and prototyping of an optimized compiler. Formal Aspects of Computing, 6(6):643–658, December 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03259390.

Minamide:1994:SAB

[169] Yasuhiko Minamide. Sharing analysis based on type inference. Formal Aspects of Computing, 6(6):659–675, December 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03259391.

Inverardi:1994:APR

[170] Paola Inverardi, Corrado Priami, and Daniel Yankelevich. Automatizing parametric reasoning on distributed concurrent systems. Formal Aspects of Computing, 6(6):676–695, December 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03259392.

Stannett:1994:ICS

[171] Mike Stannett. Infinite concurrent systems — I. The relationship between metric and order convergence. Formal Aspects of Computing, 6(6): 696–715, December 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF03259393.

Stomp:1994:PSR

[172] F. A. Stomp and W.-R de Roever. A principle for sequential reasoning about distributed algorithms. Formal Aspects of Computing, 6(6):716–737, December 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03259394.

deBoer:1994:E

[173] F. S. de Boer, E.-R. Olderog, A. Ponse, and F.-J. de Vries. Editorial. Formal Aspects of Computing, 6(1S):741– 742, 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01213600;   http://link.springer.
com/article/10.1007/BF01213600.

Apt:1994:RAP

[174] Krzysztof R. Apt and Elena Marchiori. Reasoning about Prolog programs: From modes through types to assertions. Formal Aspects of Computing, 6(1S):743–765, 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01213601.

vanEijck:1994:PFC

[175] Jan van Eijck. Presupposition failure — a comedy of errors. Formal Aspects of Computing, 6(1S):766– 787, 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01213602.

Bonsangue:1994:WPC

[176] Marcello M. Bonsangue and Joost N. Kok. The weakest precondition calculus: Recursion and duality. Formal Aspects of Computing, 6(1S):788– 800, 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01213603.

Hooman:1994:EHL

[177] Jozef Hooman. Extending Hoare logic to real-time. Formal Aspects of Computing, 6(1S):801–825, 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01213604.

Hansen:1994:MCD

[178] Michael R. Hansen. Model-checking discrete duration calculus. Formal Aspects of Computing, 6(1S):826–845, 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01213605.

Atkinson:1994:BR

[179] M. D. Atkinson and Julian Bradfield. Book reviews. Formal Aspects of Computing, 6(1S):846–848, 1994. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01213606.

Hofstee:1995:JVS

[180] H. Peter Hofstee. On Jan L. A. van de Snepscheut’s “The Sliding-Window Protocol Revisited”. Formal Aspects of Computing, 7(1):1–2, January 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01214619;   http://link.springer.
com/article/10.1007/BF01214619.

See [181].

vandeSnepscheut:1995:SWP

[181] Jan L. A. van de Snepscheut. The sliding-window protocol revisited. Formal Aspects of Computing, 7(1): 3–17, January 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.

1007/BF01214620. See comment [180].

Costa:1995:PAC

[182] J. F. Costa and A. Sernadas. Progress assumption in concurrent systems. Formal Aspects of Computing, 7 (1):18–36, January 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214621.

Butler:1995:ASU

[183] Michael Butler and Carroll Morgan. Action systems, unbounded nondeterminism, and infinite traces. Formal Aspects of Computing, 7(1):37–53, January 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01214622.

King:1995:ERC

[184] Steve King and Carroll Morgan. Exits in the refinement calculus. Formal Aspects of Computing, 7(1):54–76, January 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01214623.

Lescanne:1995:TRS

[185] Pierre Lescanne. Termination of rewrite systems by elementary interpretations. Formal Aspects of Computing, 7(1):77–90, January 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214624.

Coenen:1995:HLV

[186] J. Coenen. Hoare’s logic and VDM. Formal Aspects of Computing, 7(1): 91–105, January 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214625.

vonWright:1995:BR

[187] Joakim von Wright, Peter Jeavons, Julian Bradfield, Muffy Thomas, and David Lester. Book reviews. Formal Aspects of Computing, 7(1): 106–110, January 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01214626.

Cooke:1995:E

[188] John Cooke. Editorial. Formal Aspects of Computing, 7(2):111–112, March 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01211601;   http://link.springer.
                                        com/article/10.1007/BF01211601.

Kobayashi:1995:ACM

[189] Naoki Kobayashi and Akinori Yonezawa. Asynchronous communication model based on linear logic. Formal Aspects of Computing, 7(2):113–149, March 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF01211602. See

note [286].

Spivey:1995:UCS

[190] J. M. Spivey. Unification: A casestudy in data refinement. Formal Aspects of Computing, 7(2):150–168, March 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211603.

Rewitzky:1995:PTP

[191] I. Rewitzky and C. Brink. Predicate transformers as power operations. Formal Aspects of Computing, 7(2):169–182, March 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211604.

Vadera:1995:PAM

[192] Sunil Vadera. Proof by analogy in mural. Formal Aspects of Computing, 7(2):183–206, March 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211605.

Todd:1995:FME

[193] Bryan S. Todd, FRCS and Richard Stamper. A formal model of explanation. Formal Aspects of Computing, 7(2):207–225, March 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211606.

Mukherjee:1995:BR

[194] Paul Mukherjee, John Stell, and Carron Kirkwood. Book reviews. Formal Aspects of Computing, 7(2):226–229, March 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211607.

Lynch:1995:CST

[195] Nancy Lynch and Roberto Segala. A comparison of simulation techniques and algebraic techniques for verifying concurrent systems. Formal Aspects of Computing, 7(3):231–265, May 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211073.

Sinclair:1995:ERS

[196] Jane Sinclair and Jim Woodcock. Event refinement in state-based concurrent systems. Formal Aspects of Computing, 7(3):266–288, May 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211074.

Smith:1995:FAS

[197] Graeme Smith. A fully abstract semantics of classes for Object-Z. Formal Aspects of Computing, 7(3):289–313, May 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211075.

Leth:1995:SFC

[198] Lone Leth and Bent Thomsen. Some facile chemistry. Formal Aspects of Computing, 7(3):314–328, May 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211076.

Hayes:1995:UUM

[199] Ian J. Hayes and Brendan P. Mahony. Using units of measurement in formal specifications. Formal Aspects of Computing, 7(3):329–347, May 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211077.

Walsh:1995:BR

[200] Toby Walsh, K. J. Turner, and Jan Joris Vereijken. Book reviews. Formal Aspects of Computing, 7 (3):348–351, May 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211078.

Dijkstra:1995:DSF

[201] Rutger M. Dijkstra. DUALITY: A simple formalism for the analysis of UNITY. Formal Aspects of Computing, 7(4):353–388, July 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211214.

Hesselink:1995:SPR

[202] Wim H. Hesselink. Safety and progress of recursive procedures. Formal Aspects of Computing, 7(4):389–411, July 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211215.

Thompson:1995:LMR

[203] Simon Thompson. A logic for Miranda, revisited. Formal Aspects of Computing, 7(4):412–429, July 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211216.

Hayes:1995:SIS

[204] I. J. Hayes and J. W. Sanders. Specification by interface separation. Formal Aspects of Computing, 7(4):430–439, July 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211217.

Valmari:1995:CFB

[205] Antti Valmari and Martti Tienari. Compositional failure-based semantic models for Basic LOTOS. Formal Aspects of Computing, 7(4):440–468, July 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211218.

Brock:1995:BR

[206] Simon Brock and P. Gibson. Book reviews. Formal Aspects of Computing, 7(4):469–472, July 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211219.

Ladkin:1995:IMF

[207] Peter B. Ladkin and Stefan Leue. Interpreting message flow graphs. Formal Aspects of Computing, 7(5): 473–509, September 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211629.

Liu:1995:VSR

[208] Zhiming Liu, Mathai Joseph, and Tomasz Janowski. Verification of schedulability for real-time programs. Formal Aspects of Computing, 7(5): 510–532, September 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211630.

Barringer:1995:MI

[209] H. Barringer, M. Fisher, D. Gabbay, G. Gough, and R. Owens. MetateM: An introduction. Formal Aspects of Computing, 7(5):533–549, September 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211631.

Engelhardt:1995:TPA

[210] Kai Engelhardt and Willem-Paul de Roever. Towards a practitioners’ approach to Abadi and Lamport’s method. Formal Aspects of Computing, 7(5):550–575, September 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211632.

Leino:1995:MSP

[211] K. Rustan M. Leino. A method for showing progress. Formal Aspects of Computing, 7(5):576–580, September 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211633.

Mitchell:1995:BR

[212] W. P. R. Mitchell, Alan Hamilton, A. J. McIsaac, and Iain Stewart. Book reviews. Formal Aspects of Computing, 7(5):581–585, September 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211634.

Noel:1995:TBS

[213] Philippe Noël. A transformation-based synthesis of temporal specifications.

Formal Aspects of Computing, 7(6): 587–619, November 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01210997.

Maung:1995:SSS

[214] Ian Maung. On simulation, subtyping and substitutability in sequential object systems. Formal Aspects of Computing, 7(6):620–651, November 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01210998.

Naumann:1995:DRC

[215] David A. Naumann. Data refinement, call by value and higher order programs. Formal Aspects of Computing, 7(6):652–662, November 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01210999.

Gorlatch:1995:PDC

[216] Sergei Gorlatch and Christian Lengauer. Parallelization of divide-and-conquer in the Bird–Meertens formalism. Formal Aspects of Computing, 7(6): 663–682, November 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211000.

Manohar:1995:CC

[217] Rajit Manohar, K. Rustan, and M. Leino. Conditional composition. Formal Aspects of Computing, 7(6): 683–703, November 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211001.

Mukherjee:1995:TOS

[218] Paul Mukherjee and Victoria Stavridou. A theory of Orwellian specifications with NewThink. Formal Aspects of Computing, 7(6):704–727, November 1995. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211002.

Broy:1996:E

[219] Manfred Broy and Cliff Jones. Editorial. Formal Aspects of Computing, 8(1):1–2, January 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01211048;   http:
//link.springer.com/article/10.
1007/BF01211048.

Alencar:1996:LFE

[220] Paulo S. C. Alencar and Carlos J. P. de Lucena. A logical framework for evolving software systems. Formal Aspects of Computing, 8(1):3–46, January 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211049.

Larsen:1996:SUD

[221] Peter Gorm Larsen and Bo Stig Hansen. Semantics of underdetermined expressions. Formal Aspects of Computing, 8(1):47–66, January 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211050.

Seidel:1996:SDV

[222] Karen Seidel and Paul Gardiner. Structured development of a virtual shared memory system. Formal Aspects of Computing, 8(1):67–85, January 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211051.

Simons:1996:ALS

[223] Martin Simons and Matthias Weber. An approach to literate and structured formal developments. Formal Aspects of Computing, 8(1):86–107, January 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211052.

Billington:1996:CIG

[224] David Billington and R. Geoff Dromey. The co-invariant generator: An aid in deriving loop bodies. Formal Aspects of Computing, 8(1):108–126, January 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211053.

Stolen:1996:SRN

[225] Ketil Stølen, Frank Dederichs, and Rainer Weber. Specification and refinement of networks of asynchronously communicating agents using the assumption/commitment paradigm. Formal Aspects of Computing, 8(2): 127–161, March 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214554.

Krishnan:1996:AC

[226] Padmanabhan Krishnan. Architectural CCS. Formal Aspects of Computing, 8(2):162–187, March 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214555.

Baeten:1996:DTP

[227] J. C. M. Baeten and J. A. Bergstra. Discrete time process algebra. Formal Aspects of Computing, 8(2):188–208, March 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01214556.

Brown:1996:RHC

[228] Geoffrey Brown, Wayne Luk, and John O’Leary. Retargeting a hardware compiler using protocol converters. Formal Aspects of Computing, 8(2):209–237, March 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01214557.

Hussak:1996:CPR

[229] Walter Hussak. On CCS with parametric relabelling. Formal Aspects of Computing, 8(2):238–244, March 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01214558.

Bowen:1996:BR

[230] Jonathan Bowen and Dieter Gollmann. Book reviews. Formal Aspects of Computing, 8(2):245–246, March 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01214559;   http://link.springer.
com/article/10.1007/BF01214559.

Castagna:1996:IPA

[231] Giuseppe Castagna. Integration of parametric and “ad hoc” second order polymorphism in a calculus with subtyping. Formal Aspects of Computing, 8(3):247–293, May 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214916.

Zhou:1996:CVR

[232] Ping Zhou, Jozef Hooman, and Ruurd Kuiper. Compositional verification of real-time systems with explicit clock temporal logic. Formal Aspects of Computing, 8(3):294–323, May 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01214917.

Back:1996:SRR

[233] R. J. R. Back and K. Sere. Superposition refinement of reactive systems. Formal Aspects of Computing, 8(3):324–346, May 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214918.

Duke:1996:BMR

[234] Roger Duke, Cecily Bailes, and Graeme Smith. A blocking model for reactive objects. Formal Aspects of Computing, 8(3):347–368, May 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214919.

Spivey:1996:CTF

[235] Michael Spivey. The consistency theorem for free type definitions in Z. Formal Aspects of Computing, 8(3):369–375, May 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01214920.

Matthews:1996:BR

[236] Brian Matthews, Andrew Blyth, and J. S. Fitzgerald. Book reviews. Formal Aspects of Computing, 8(3):376–378, May 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01214921.

Hennessy:1996:PSM

[237] M. Hennessy and H. Lin. Proof systems for message-passing process algebras. Formal Aspects of Computing, 8(4):379–407, July 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01213531.

Scholefield:1996:RTR

[238] David Scholefield. Real-time refinement in Manna and Pnueli’s temporal logic. Formal Aspects of Computing, 8(4):408–427, July 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01213532.

Beierle:1996:SCP

[239] Christoph Beierle and Egon Börger. Specification and correctness proof of a WAM extension with abstract type constraints. Formal Aspects of Computing, 8(4):428–462, July 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01213533.

Ronn:1996:ICT

[240] Stefan Rönn. Invariants and closures in the theory of rewrite systems. Formal Aspects of Computing, 8(4):463–478, July 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01213534.

Martin:1996:TCA

[241] A. P. Martin, P. H. B. Gardiner, and J. C. P. Woodcock. A tactic calculus — abridged version. Formal Aspects of Computing, 8(4):479–489, July 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01213535.

Harman:1996:SPP

[242] Mark Harman, Dan Simpson, and Sebastian Danicic. Slicing programs in the presence of errors. Formal Aspects of Computing, 8(4):490–497, July 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01213536.

Kirkwood:1996:BR

[243] Carron Kirkwood. Book review. Formal Aspects of Computing, 8(4):498, July 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01213537;   http://link.springer.
com/article/10.1007/BF01213537.

Lynch:1996:ATT

[244] Nancy Lynch and Frits Vaandrager. Action transducers and timed automata. Formal Aspects of Computing, 8(5):499–538, September 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/BF01211907.

Beierle:1996:RTW

[245] Christoph Beierle and Egon Börger. Refinement of a typed WAM extension by polymorphic order-sorted types. Formal Aspects of Computing, 8(5): 539–564, September 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211908.

Spivey:1996:RTZ

[246] Michael Spivey. Richer types for Z. Formal Aspects of Computing, 8(5): 565–584, September 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211909.

Smolka:1996:PEP

[247] Scott A. Smolka and Bernhard Steffen. Priority as extremal probability.

Formal Aspects of Computing, 8(5): 585–606, September 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211910.

Graf:1996:CMF

[248] Susanne Graf, Bernhard Steffen, and Gerald Lüttgen. Compositional minimisation of finite state systems using interface specifications. Formal Aspects of Computing, 8(5):607–616, September 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211911.

Morgan:1996:ROP

[249] Carroll Morgan, Annabelle McIver, Karen Seidel, and J. W. Sanders. Refinement-oriented probability for CSP. Formal Aspects of Computing, 8 (6):617–647, November 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01213492.

Ponse:1996:CPB

[250] Alban Ponse. Computable processes and bisimulation equivalence. Formal Aspects of Computing, 8(6): 648–678, November 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01213493.

Zucker:1996:TNI

[251] J. I. Zucker. Transformations of normal and inverted function tables. Formal Aspects of Computing, 8(6):

679–705, November 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01213494.

Breu:1996:ASF

[252] Ruth Breu and Elena Zucca. An algebraic semantic framework for object oriented languages with concurrency (extended abstract). Formal Aspects of Computing, 8(6):706–715, November 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01213495.

Sitaraman:1996:IPC

[253] Murali Sitaraman. Impact of performance considerations on formal specification design. Formal Aspects of Computing, 8(6):716–736, November 1996. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01213496.

Bezem:1997:FPA

[254] Marc Bezem, Roland Bol, and Jan Friso Groote. Formalizing process algebraic verifications in the calculus of constructions. Formal Aspects of Computing, 9(1):1–48, January 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01212523.

Palsberg:1997:TIN

[255] Jens Palsberg, Mitchell Wand, and Patrick O’Keefe. Type inference with non-structural subtyping. Formal Aspects of Computing, 9(1):49–67, January 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01212524.

Caswell:1997:EFS

[256] M. J. A. Caswell. Equivalence of formal semantics definition methods. Formal Aspects of Computing, 9(1):68–77, January 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01212525.

Nishimura:1997:SFL

[257] Susumu Nishimura. A strict functional language with cyclic recursive data. Formal Aspects of Computing, 9(1):78–97, January 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01212526.

Krogdahl:1997:VDL

[258] Stein Krogdahl and Olav Lysne. Verifying a distributed list system: A case history. Formal Aspects of Computing, 9(1):98–118, January 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01212527.

Koutny:1997:TIR

[259] Maciej Koutny, Luigi V. Mancini, and Giuseppe Pappalardo. Two implementation relations and the correctness of communicating replicated processes. Formal Aspects of Computing, 9(2):119–148, March 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211616.

Xu:1997:RGM

[260] Qiwen Xu, Willem-Paul de Roever, and Jifeng He. The rely-guarantee method for verifying shared variable concurrent programs. Formal Aspects of Computing, 9(2):149–174, March 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211617.

Fidge:1997:ABF

[261] C. J. Fidge and A. J. Wellings. An action-based formal model for concurrent real-time systems. Formal Aspects of Computing, 9(2):175–207, March 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211618.

Hesselink:1997:MPS

[262] Wim H. Hesselink. A mechanical proof of Segall’s PIF algorithm. Formal Aspects of Computing, 9(2):208–226, March 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211619.

Wing:1997:E

[263] Jeannette Wing. Editorial. Formal Aspects of Computing, 9(3):227–228, May 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01211083;   http://link.springer.
com/article/10.1007/BF01211083.

Sannella:1997:ECA

[264] Donald Sannella and Andrzej Tarlecki. Essential concepts of algebraic specification and program development. Formal Aspects of Computing, 9(3):229–269, May 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211084.

Dijkstra:1997:PTP

[265] Rutger M. Dijkstra and Beverly A. Sanders. A predicate transformer for the progress property ‘to-always’. Formal Aspects of Computing, 9 (3):270–282, May 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211085.

Hansen:1997:DCL

[266] Michael R. Hansen and Zhou Chaochen. Duration calculus: Logical foundations. Formal Aspects of Computing, 9(3):283–330, May 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211086.

Ross:1997:RPB

[267] Brian J. Ross. Running programs backwards: The logical inversion of imperative computation. Formal Aspects of Computing, 9(3):331–348, May 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211087.

OHalloran:1997:UFS

[268] C. O’Halloran, R. Arthan, and D. King. Using a formal specification contractually. Formal Aspects of Computing, 9(4):349–358, July 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211295.

Jutla:1997:MDP

[269] Charanjit S. Jutla and Josyula R. Rao. A methodology for designing proof rules for fair parallel programs. Formal Aspects of Computing, 9(4):359–378, July 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211296.

Kneuper:1997:LFM

[270] Ralf Kneuper. Limits of formal methods. Formal Aspects of Computing, 9(4):379–394, July 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01211297.

Josephs:1997:USA

[271] Mark B. Josephs and Andrew M. Bailey. The use of SI–Algebra in the design of sequencer circuits. Formal Aspects of Computing, 9(4):395–408, July 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211298.

Kindler:1997:PNB

[272] Ekkart Kindler, Wolfgang Reisig, Hagen Völzer, and Rolf Walter. Petri net based verification of distributed algorithms: An example. Formal Aspects of Computing, 9(4):409–424, July 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211299.

Baier:1997:CBI

[273] Christel Baier and Mila Majster-Cederbaum. The connection between initial and unique solutions of domain equations in the partial order and metric approach. Formal Aspects of Computing, 9(4):425–445, July 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211300.

Cooke:1997:E

[274] John Cooke. Editorial. Formal Aspects of Computing, 9(5–6):447, September 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01211454;   http://link.springer.
                                        com/article/10.1007/BF01211454.

Hesselink:1997:TMP

[275] Wim H. Hesselink. Theories for mechanical proofs of imperative programs. Formal Aspects of Computing, 9(5–6): 448–468, September 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01211455.

Back:1997:SCP

[276] Ralph Back, Jim Grundy, and Joakim von Wright. Structured calculational proof. Formal Aspects of Computing, 9(5–6):469–483, September 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01211456.

Parent-Vigouroux:1997:VPC

[277] Catherine Parent-Vigouroux. Verifying programs in the calculus of inductive constructions. Formal Aspects of Computing, 9(5–6):484–517, September 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211457.

Knijnenburg:1997:SCA

[278] Peter M. W. Knijnenburg and Joost N. Kok. The semantics of the combination of atomized statements and parallel choice. Formal Aspects of Computing, 9(5–6):518–536, September 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211458.

OLeary:1997:VCC

[279] John O’Leary, Geoffrey Brown, and Wayne Luk. Verified compilation of communicating processes into clocked circuits. Formal Aspects of Computing, 9(5–6):537–559, September 1997. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01211459.

Rewitzky:1998:UFV

[280] Ingrid Rewitzky and Chris Brink. Unification of four versions of program semantics. Formal Aspects of Computing, 10(1):1–29, September 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003923.

Korver:1998:FAA

[281] Henri Korver and Alex Sellink. A formal axiomatization for alphabet reasoning with parametrized processes. Formal Aspects of Computing, 10(1): 30–42, September 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003924.

Korver:1998:EVU

[282] Henri Korver and Alex Sellink. Example verifications using alphabet axioms. Formal Aspects of Computing, 10(1):43–58, September 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003925.

Leavens:1998:PIS

[283] Gary T. Leavens and Jeannette M. Wing. Protective interface specifications. Formal Aspects of Computing, 10(1):59–75, September 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003926.

Morris:1998:PNP

[284] Joseph M. Morris and Alexander Bunkenburg. Partiality and nondeterminacy in program proofs. Formal Aspects of Computing, 10(1): 76–96, September 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003927.

Derrick:1998:SRI

[285] John Derrick, Eerke Boiten, Howard Bowman, and Maarten Steen. Specifying and refining internal operations in Z. Formal Aspects of Computing, 10 (2):125–159, November 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050007.

deGroote:1998:NKY

[286] Philippe de Groote and Guy Perrier. A note on Kobayashi’s and Yonezawa’s “Asynchronous Communication Model Based on Linear Logic”. Formal Aspects of Computing, 10(2): 160–170, November 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050008. See [189].

Nipkow:1998:WAR

[287] Tobias Nipkow. Winskel is (almost) right: Towards a mechanized semantics textbook. Formal Aspects of Computing, 10(2):171–186, November 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050009.

Hayes:1998:EPS

[288] Ian J. Hayes. Expressive power of specification languages. Formal Aspects of Computing, 10(2):187–192, November 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050010.

Dingel:1998:TFT

[289] J. Dingel, D. Garlan, S. Jha, and D. Notkin. Towards a formal treatment of implicit invocation using Rely/ guarantee reasoning. Formal Aspects of Computing, 10(3):193–213, March 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650050011.

Jonker:1998:KTR

[290] Jan Eppo Jonker. Knaster–Tarski revisited. Formal Aspects of Computing, 10(3):214–232, March 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001650050012.

Paige:1998:HNP

[291] Richard F. Paige. Heterogeneous notations for pure formal method integration. Formal Aspects of Computing, 10(3):233–242, March 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001650050013.

Henson:1998:SLZ

[292] Martin C. Henson. The standard logic of Z is inconsistent. Formal Aspects of Computing, 10(3):243–247, March 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650050014.

Leonard:1998:FDT

[293] Luc Léonard and Guy Leduc. A formal definition of time in LOTOS. Formal Aspects of Computing, 10 (3):248–266, March 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050015.

Cavalcanti:1998:ZRC

[294] Ana Cavalcanti and Jim Woodcock. ZRC — a refinement calculus for Z. Formal Aspects of Computing, 10 (3):267–289, March 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050016.

Hehner:1998:FTS

[295] Eric C. R. Hehner. Formalization of time and space. Formal Aspects of Computing, 10(3):290–306, March 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050017.

Anonymous:1998:BR

[296] Anonymous. Book reviews. Formal Aspects of Computing, 10(3): 307–310, March 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050018.

Gnesi:1998:E

[297] Stefania Gnesi and Diego Latella. Editorial. Formal Aspects of Computing, 10(4):311–312, April 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s001650050019;
http://link.springer.com/article/
10.1007/s001650050019.

Bernardo:1998:FPM

[298] Marco Bernardo, Roberto Gorrieri, and Marco Roccetti. Formal performance modelling and evaluation of an adaptive mechanism for packetised audio over the Internet. Formal Aspects of Computing, 10(4):313–337, April 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650050020.

Boralv:1998:CSF

[299] Arne Borälv. Case study: Formal verification of a computerized railway interlocking. Formal Aspects of Computing, 10(4):338–360, April 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650050021.

Cimatti:1998:FVR

[300] A. Cimatti, F. Giunchiglia, G. Mongardi, D. Romano, F. Torielli, and P. Traverso. Formal verification of a railway interlocking system using model checking. Formal Aspects of Computing, 10(4):361–380, April 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650050022.

Ochsenschlager:1998:SVT

[301] P. Ochsenschläger, J. Repp, R. Rieke, and U. Nitsche. The SH-Verification tool — abstraction-based verification of co-operating systems. Formal Aspects of Computing, 10(4):381–404, April 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050023.

Faconti:1998:MVP

[302] G. P. Faconti and M. Massink. Modelling and verification of PREMO synchronisable objects. Formal Aspects of Computing, 10(4):405–434, April 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050024.

Groote:1998:E

[303] Jan Friso Groote, Bas Luttik, and Jos van Wamel. Editorial. Formal Aspects of Computing, 10(5–6): 435, May 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s001650050025;   http://
link.springer.com/article/10.1007/
s001650050025.

Gaudel:1998:TAD

[304] Marie-Claude Gaudel and Perry R. James. Testing algebraic data types and processes: A unifying theory. Formal Aspects of Computing, 10 (5–6):436–451, May 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050026.

Anderson:1998:DPA

[305] Stuart Anderson and Konstantinos Tourlas. Design for proof: An approach to the design of domain–specific languages. Formal Aspects of Computing, 10(5–6):452–468, May 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050027.

Dams:1998:POR

[306] Dennis Dams, Rob Gerth, Bart Knaack, and Ruurd Kuiper. Partialorder reduction techniques for realtime model checking. Formal Aspects of Computing, 10(5–6):469–482, May 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050028.

Vassiliou-Gioles:1998:CSP

[307] Theofanis Vassiliou-Gioles and Ina Schieferdecker. Case study in protocol validation: Validating an ATM signalling protocol. Formal Aspects of Computing, 10(5–6):483–508, May 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050029.

Shankland:1998:TIP

[308] Carron Shankland and Mark van der Zwaag. The tree identify protocol of IEEE 1394 in μCRL. Formal Aspects of Computing, 10(5–6):509–531, May 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050030.

Bicarregui:1998:IIT

[309] Juan Bicarregui, Brian Matthews, Brian Ritchie, and Sten Agerholm. Investigating the integration of two formal methods. Formal Aspects of Computing, 10(5–6):532–549, May 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050031.

Bowman:1998:AVL

[310] H. Bowman, G. Faconti, J.-P. Katoen, D. Latella, and M. Massink. Automatic verification of a lip-synchronisation protocol using Uppaal. Formal Aspects of Computing, 10(5–6):550–575, May 1998. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050032.

Duce:1999:CFS

[311] David Duce, David Duke, Giorgio Faconti, and Ivan Herman. The changing face of standardization: A place for formal methods? Formal Aspects of Computing, 11(1):1–20, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050033.

Hung:1999:PDC

[312] Dang Van Hung and Zhou Chaochen. Probabilistic duration calculus for continuous time. Formal Aspects of Computing, 11(1):21–44, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050034.

Hesselink:1999:VID

[313] Wim H. Hesselink. The verified incremental design of a distributed spanning tree algorithm: Extended abstract. Formal Aspects of Computing, 11(1):45–55, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050035.

Leino:1999:CPE

[314] K. Rustan M. Leino. Computing permutation encodings. Formal Aspects of Computing, 11(1):56–74, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050036.

Mahony:1999:LCR

[315] Brendan P. Mahony. The least conjunctive refinement and promotion in the refinement calculus. Formal Aspects of Computing, 11(1): 75–105, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050037.

Duke:1999:CSS

[316] David Duke, Bob Fields, and Michael D. Harrison. A case study in the specification and analysis of design alternatives for a user interface. Formal Aspects of Computing, 11(2): 107–131, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050044.

Bowman:1999:ACB

[317] Howard Bowman and Giorgio Faconti. Analysing cognitive behaviour using LOTOS and Mexitl. Formal Aspects of Computing, 11(2):132–159, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050045.

Nesi:1999:FVP

[318] Monica Nesi. Formalising a valuepassing calculus in HOL. Formal Aspects of Computing, 11(2): 160–199, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050046.

Bird:1999:GFN

[319] Richard Bird and Ross Paterson. Generalised folds for nested datatypes. Formal Aspects of Computing, 11(2): 200–222, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050047.

Anonymous:1999:E

[320] Anonymous. Editorial. Formal Aspects of Computing, 11(3):223–224, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s001650050048;   http://
link.springer.com/article/10.1007/
s001650050048.

Bertot:1999:CSD

[321] Yves Bertot. The CtCoq system: Design and architecture. Formal Aspects of Computing, 11(3): 225–243, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050049.

Bornat:1999:MGU

[322] Richard Bornat and Bernard Sufrin. A minimal graphical user interface for the Jape proof calculator. Formal Aspects of Computing, 11(3): 244–271, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050050.

Goguen:1999:SSA

[323] Joseph Goguen. Social and semiotic analyses for theorem prover user interface design 1. Formal Aspects of Computing, 11(3):272–301, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050051.

Ireland:1999:IPC

[324] Andrew Ireland, Michael Jackson, and Gordon Reid. Interactive proof critics. Formal Aspects of Computing, 11 (3):302–325, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050052.

Siekmann:1999:LUI

[325] Jörg Siekmann, Stephan Hess, Christoph Benzmüller, Lassaad Cheikhrouhou, Armin Fiedler, Helmut Horacek, Michael Kohlhase, Karsten Konrad, Andreas Meier, Erica Melis, Martin Pollet, and Volker Sorge. LΩUI: Lovely ΩMEGA user interface. Formal Aspects of Computing, 11(3): 326–342, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050053.

Takahashi:1999:PEH

[326] Koichi Takahashi and Masami Hagiya. Proving as editing HOL tactics. Formal Aspects of Computing, 11(3): 343–357, September 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050054.

Henson:1999:RZPa

[327] Martin C. Henson and Steve Reeves. Revising Z: Part I — logic and semantics. Formal Aspects of Computing, 11 (4):359–380, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050038.

Henson:1999:RZPb

[328] Martin C. Henson and Steve Reeves. Revising Z: Part II — logical development. Formal Aspects of Computing, 11 (4):381–401, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050039.

Sistla:1999:PVL

[329] A. Prasad Sistla and Viktor Gyuris. Parameterized verification of linear networks using automata as invariants. Formal Aspects of Computing, 11(4): 402–425, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050040.

Audebaud:1999:DPR

[330] Philippe Audebaud and Elena Zucca. Deriving proof rules from continuation semantics. Formal Aspects of Computing, 11(4):426–447, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650050041.

Roux:1999:HVR

[331] Olivier Roux, Vlad Rusu, and Franck Cassez. Hybrid verifications of reactive programs. Formal Aspects of Computing, 11(4):448–471, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650050042.

King:1999:SLZ

[332] Steve King. ‘The standard logic for Z’: A clarification. Formal Aspects of Computing, 11(4):472–473, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s001650050043;   http://
                                        link.springer.com/article/10.1007/
                                        s001650050043.

Felty:1999:CCS

[333] Amy Felty and Frank Stomp. Cache coherency in SCI: Specification and a sketch of correctness. Formal Aspects of Computing, 11(5):475–497, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650050055.

Banach:1999:SRM

[334] R. Banach and M. Poppleton. Sharp retrenchment, modulated refinement and simulation. Formal Aspects of Computing, 11(5):498–540, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050056.

Kleymann:1999:HLA

[335] Thomas Kleymann. Hoare logic and auxiliary variables. Formal Aspects of Computing, 11(5):541–566, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650050057.

Pitt:1999:OAP

[336] D. H. Pitt and M. W. Shields. Overtaking in asynchronous periodic systems. Formal Aspects of Computing, 11 (5):567–590, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650050058.

Jackson:1999:NAM

[337] Daniel Jackson, Yu-Chung Ng, and Jeannette Wing. A nitpick analysis of Mobile IPv6. Formal Aspects of Computing, 11(6):591–615, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001659970001.

Hesselink:1999:PTR

[338] Wim H. Hesselink. Predicate transformers for recursive procedures with local variables. Formal Aspects of Computing, 11(6):616–636, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001659970002.

Latella:1999:AVB

[339] Diego Latella, Istvan Majzik, and Mieke Massink. Automatic verification of a behavioural subset of UML statechart diagrams using the SPIN model-checker. Formal Aspects of Computing, 11(6):637–664, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001659970003.

Duke:1999:FCA

[340] David Duke and David Duce. The formalization of a cognitive architecture and its application to reasoning about human computer interaction. Formal Aspects of Computing, 11(6): 665–689, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001659970004.

Kreowski:1999:GTU

[341] Hans-Jörg Kreowski and Sabine Kuske. Graph transformation units with interleaving semantics. Formal Aspects of Computing, 11(6):690–723, December 1999. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001659970005.

Bergstra:2000:PAC

[342] J. A. Bergstra and M. E. Loots. Program algebra for component code.

Formal Aspects of Computing, 12 (1):1–17, October 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00003928.

Back:2000:CRS

[343] Ralph-Johan Back, Anna Mikhajlova, and Joakim von Wright. Class refinement as semantics of correct object substitutability. Formal Aspects of Computing, 12(1):18–40, October 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650070034.

vonMohrenschildt:2000:ACF

[344] Martin von Mohrenschildt. Algebraic composition of function tables. Formal Aspects of Computing, 12 (1):41–51, October 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070035.

Cau:2000:CRD

[345] Antonio Cau. Composing and refining dense temporal logic specifications. Formal Aspects of Computing, 12(1):52–70, October 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070036.

Diehl:2000:NSD

[346] Stephan Diehl. Natural semantics– directed generation of compilers and abstract machines. Formal Aspects of Computing, 12(2):71–99, October 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/PL00003929.

Joshi:2000:MCP

[347] Rajeev Joshi and Jayadev Misra. Maximally concurrent programs. Formal Aspects of Computing, 12(2): 100–119, October 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001650070031.

Paynter:2000:AAD

[348] Stephen Paynter, Jim Armstrong, and Jan Haveman. ADL: An activity description language for real-time networks. Formal Aspects of Computing, 12(2):120–144, October 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001650070032.

Wing:2000:FWC

[349] Jeanette Wing and Jim Woodcock. The first world Congress on formal methods in the development of computing systems. Formal Aspects of Computing, 12(3):145–146, November 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        PL00011167;   http://link.springer.
                                        com/article/10.1007/PL00011167.

Wehrheim:2000:DAT

[350] Heike Wehrheim. Data abstraction techniques in the validation of CSP– OZ specifications. Formal Aspects of Computing, 12(3):147–164, November 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650070026.

Schneider:2000:ATC

[351] Steve Schneider. Abstraction and testing in CSP. Formal Aspects of Computing, 12(3):165–181, November 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650070027.

Butler:2000:CPA

[352] Michael Butler. csp2B: A practical approach to combining CSP and B. Formal Aspects of Computing, 12(3): 182–198, November 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00003930.

Martin:2000:RZF

[353] Andrew P. Martin. Relating Z and first-order logic. Formal Aspects of Computing, 12(3):199–209, November 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650070029.

Denvir:2000:CPA

[354] Tim Denvir, José Oliveira, and Nico Plat. The cash-point (ATM) ‘problem’. Formal Aspects of Computing, 12 (4):211–215, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070012.

Larsen:2000:UVM

[355] Peter Gorm Larsen, Paul Mukherjee, and Kim Sunesen. Using VDM-Tools to model and validate the cash dispenser example. Formal Aspects of Computing, 12(4):216–217, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s001650070013;   http://
link.springer.com/article/10.1007/
s001650070013.

Browne:2000:CPS

[356] Anca Browne, Bernd Finkbeiner, Zohar Manna, and Henny Sipma. The ‘cash-point’ service: A verification case study using STeP. Formal Aspects of Computing, 12(4):218–219, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s001650070014;   http://
link.springer.com/article/10.1007/
s001650070014.

Dierks:2000:MVC

[357] Henning Dierks and Josef Tapken. Modelling and verifying of a ‘cash-point service’ using MOBY/ PLC. Formal Aspects of Computing, 12(4):220–221, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s001650070015;   http://
link.springer.com/article/10.1007/
s001650070015.

Kotkas:2000:CPS

[358] Vahur Kotkas, Peep Küngas, and Mait Harf. The cash-point service in NUT. Formal Aspects of Computing, 12(4):222–224, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003931.

Slotosch:2000:MVA

[359] Oscar Slotosch. Modelling and validation: AUTOFOCUS and quest. Formal Aspects of Computing, 12(4): 225–227, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070017.

Dupuy:2000:MFA

[360] Sophie Dupuy and Lydie du Bousquet. A multi-formalism approach for the validation of UML models. Formal Aspects of Computing, 12(4): 228–230, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070018.

Aaltonen:2000:SCP

[361] Timo Aaltonen, Pertti Kellomäki, and Risto Pitkänen. Specifying cashpoint with DisCo. Formal Aspects of Computing, 12(4):231–232, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
PL00003932;   http://link.springer.
com/article/10.1007/PL00003932.

Dissoubray:2000:DAT

[362] Sylvan Dissoubray and Bernard Dion. Design of an automatic teller machine with Esterel Studio. Formal Aspects of Computing, 12(4):233–236, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/PL00003933.

Butterworth:2000:DCP

[363] Richard Butterworth, Ann Blandford, and David Duke. Demonstrating the cognitive plausibility of interactive system specifications. Formal Aspects of Computing, 12(4): 237–259, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001650070021.

Doherty:2000:RRV

[364] Gavin J. Doherty, José C. Campos, and Michael D. Harrison. Representational reasoning and verification. Formal Aspects of Computing, 12(4):260–277, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/PL00003934.

Sere:2000:DRR

[365] Kaisa Sere and Marina Waldén. Data refinement of remote procedures. Formal Aspects of Computing, 12(4): 278–297, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003935.

Fox:2000:AMC

[366] A. C. J. Fox and N. A. Harman. Algebraic models of correctness for microprocessors. Formal Aspects of Computing, 12(4):298–312, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003936.

Back:2000:EDD

[367] Ralph-Johan Back and Joakim von Wright. Encoding, decoding and data refinement. Formal Aspects of Computing, 12(5):313–349, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650070008.

Cardell-Oliver:2000:CTR

[368] Rachel Cardell-Oliver. Conformance tests for real-time systems with timed automata specifications. Formal Aspects of Computing, 12(5): 350–371, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070009.

Staples:2000:IRR

[369] Mark Staples. Interfaces for refining recursion and procedures. Formal Aspects of Computing, 12(5): 372–391, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070010.

Tofts:2000:SAP

[370] C. Tofts. Symbolic approaches to probability distributions in process algebra. Formal Aspects of Computing, 12 (5):392–415, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00013291.

Jones:2000:E

[371] Cliff Jones. Editorial. Formal Aspects of Computing, 12(6):417, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
PL00011168;   http://link.springer.
com/article/10.1007/PL00011168.

Holcombe:2000:WXM

[372] Mike Holcombe. What are Xmachines? Formal Aspects of Computing, 12(6):418–422, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650070002.

Hierons:2000:TCQ

[373] R. M. Hierons and M. Harman. Testing conformance to a quasi–nondeterministic stream X-machine. Formal Aspects of Computing, 12(6): 423–442, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070003.

Ipate:2000:GTS

[374] Florentin Ipate and Mike Holcombe. Generating test sets from non– deterministic stream X-machines. Formal Aspects of Computing, 12(6): 443–458, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070004.

Gheorghe:2000:GSX

[375] Marian Gheorghe. Generalised stream X-machines and cooperating distributed grammar systems. Formal Aspects of Computing, 12(6): 459–472, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070005.

Balanescu:2000:GSX

[376] Tudor Bălănescu. Generalised stream X-machines with output delimited type. Formal Aspects of Computing, 12 (6):473–484, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070006.

Cowling:2000:SWU

[377] Anthony J. Cowling, Horia Georgescu, and Cristina Vertan. A structured way to use channels for communication in X-machine systems. Formal Aspects of Computing, 12(6): 485–500, December 2000. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650070007.

Grosu:2001:SBS

[378] Radu Grosu and Ketil Stølen. Streambased specification of mobile systems. Formal Aspects of Computing, 13 (1):1–31, September 2001. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00003937.

Vickers:2001:PCS

[379] Steven Vickers and Gillian Hill. Presheaves as configured specifications. Formal Aspects of Computing, 13(1): 32–49, September 2001. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003938.

Arias:2001:SMF

[380] José J. Pazos Arias and Jorge García Duque. SCTL-MUS: A formal methodology for software development of distributed systems. a case study. Formal Aspects of Computing, 13(1): 50–91, September 2001. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003939.

Cooke:2002:Ea

[381] John Cooke and Tim Denvir. Editorial. Formal Aspects of Computing, 13 (2):93, May 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s001650200000;   http://
                                        link.springer.com/article/10.1007/
                                        s001650200000.

Suhl:2002:OIF

[382] Carsten Sühl. An overview of the integrated formalism RT-Z. Formal Aspects of Computing, 13(2):94–110, May 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650200001.

Derrick:2002:CCS

[383] John Derrick and Eerke Boiten. Combining component specifications in Object-Z and CSP. Formal Aspects of Computing, 13(2):111–127, May 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200002.

Smith:2002:IRT

[384] Graeme Smith and Ian Hayes. An introduction to real-time Object-Z. Formal Aspects of Computing, 13 (2):128–141, May 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200003.

Mahony:2002:DSL

[385] Brendan Mahony and Jin Song Dong. Deep semantic links of TCSP and Object-Z: TCOZ approach. Formal Aspects of Computing, 13(2): 142–160, May 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200004.

Grosse-Rhode:2002:CCF

[386] Martin Große-Rhode. Compositional comparison of formal software specifications using transformation systems. Formal Aspects of Computing, 13(2):161–186, May 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200005.

Rydeheard:2002:CPM

[387] David E. Rydeheard and Donald T. Sannella. A collection of papers and memoirs celebrating the contribution of Rod Burstall to advances in computer science. Formal Aspects of Computing, 13(3–5):187–193, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200006.

Kerse:2002:ORB

[388] Eleanor Kerse. Ode to rod burstall. Formal Aspects of Computing, 13(3–5): 194, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s001650200007;   http://
link.springer.com/article/10.1007/
s001650200007.

Landin:2002:RBP

[389] Peter Landin. Rod burstall: A personal note. Formal Aspects of Computing, 13(3–5):195, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s001650200008;
http://link.springer.com/article/
10.1007/s001650200008.

Popplestone:2002:PBS

[390] Robin Popplestone. POP, a broadspectrum programming language, 1967–2002. Formal Aspects of Computing, 13(3–5):196–213, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200009.

MacQueen:2002:SMO

[391] David MacQueen. Should ML be object-oriented? Formal Aspects of Computing, 13(3–5):214–232, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200010.

Pettorossi:2002:LIS

[392] Alberto Pettorossi and Maurizio Proietti. The list introduction strategy for the derivation of logic programs. Formal Aspects of Computing, 13(3–5):233–251, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200011.

Bidoit:2002:ASC

[393] Michel Bidoit, Donald Sannella, and Andrzej Tarlecki. Architectural specifications in CASL. Formal Aspects of Computing, 13(3–5):252–273, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200012.

Goguen:2002:IM

[394] Joseph Goguen and Grigore Rosu. Institution morphisms. Formal Aspects of Computing, 13(3–5):274–307, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200013.

Robinson:2002:VAM

[395] Edmund Robinson. Variations on algebra: Monadicity and generalisations of equational theories. Formal Aspects of Computing, 13(3–5):308–326, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200014.

Leifer:2002:SLA

[396] James J. Leifer and Robin Milner. Shallow linear action graphs and their embeddings. Formal Aspects of Computing, 13(3–5):327–340, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200015.

Gabbay:2002:NAA

[397] Murdoch J. Gabbay and Andrew M. Pitts. A new approach to abstract syntax with variable binding. Formal Aspects of Computing, 13(3– 5):341–363, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200016.

Plotkin:2002:TIM

[398] Gordon Plotkin. Three inadequate models. Formal Aspects of Computing, 13(3–5):364–385, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200017.

Pollack:2002:DTR

[399] Robert Pollack. Dependently typed records in type theory. Formal Aspects of Computing, 13(3–5):386–402, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200018.

Whittle:2002:PPF

[400] Jon Whittle, Alan Bundy, and Richard Boulton. Proofs-as-programs as a framework for the design of an analogybased ML Editor. Formal Aspects of Computing, 13(3–5):403–421, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200019.

Barendregt:2002:ATM

[401] Henk Barendregt. The ancient theory of mind. Formal Aspects of Computing, 13(3–5):422–429, July 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200020.

Ipate:2002:TCC

[402] Florentin Ipate and Mike Holcombe. Testing conditions for communicating stream X-machine systems. Formal Aspects of Computing, 13(6): 431–446, August 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200021.

Badouel:2002:DFA

[403] Éric Badouel, Benoît Caillaud, and P. Darondeau. Distributing finite automata through Petri net synthesis. Formal Aspects of Computing, 13 (6):447–470, August 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200022.

Krogdahl:2002:VPM

[404] Stein Krogdahl and Olav Lysne. On verification of parallel message-passing processes. Formal Aspects of Computing, 13(6):471–492, August 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650200023.

Cooke:2002:Eb

[405] John Cooke. Editorial. Formal Aspects of Computing, 14(1):1, October 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s001650200024;   http://
                                        link.springer.com/article/10.1007/
                                        s001650200024.

Butler:2002:UDR

[406] Michael Butler. On the use of data refinement in the development of secure communications systems. Formal Aspects of Computing, 14 (1):2–34, October 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001650200025.

Pitt:2002:LI

[407] David H. Pitt and Michael Shields. Local invariance. Formal Aspects of Computing, 14(1):35–54, October 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650200026.

Genrich:2002:DQN

[408] Hartmann J. Genrich. Dynamical quantities in net systems. Formal Aspects of Computing, 14(1): 55–89, October 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200027.

Jones:2002:E

[409] Cliff Jones. Editorial. Formal Aspects of Computing, 14(2):91, December 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s001650200028;   http://
link.springer.com/article/10.1007/
s001650200028.

Apt:2002:EWD

[410] Krzysztof R. Apt. Edsger Wybe Dijkstra (1930–2002): A portrait of a genius. Formal Aspects of Computing, 14(2):92–98, December 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200029.

Dijkstra:2002:ENC

[411] Edsger W. Dijkstra. EWD1300: The notational conventions I adopted, and why. Formal Aspects of Computing, 14 (2):99–107, December 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200030.

Chen:2002:GC

[412] Yifeng Chen. Generic composition. Formal Aspects of Computing, 14(2): 108–122, December 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650200031.

Dingel:2002:RCS

[413] J. Dingel. A refinement calculus for shared-variable parallel and distributed programming. Formal Aspects of Computing, 14(2):123–197, December 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650200032.

Cooke:2003:E

[414] John Cooke, Savi Maharaj, Judi Romijn, and Carron Shankland. Editorial. Formal Aspects of Computing, 14 (3):199, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s001650300000;   http://
link.springer.com/article/10.1007/
s001650300000.

Maharaj:2003:ITI

[415] Savi Maharaj, Judi Romijn, and Carron Shankland. IEEE 1394 Tree Identify Protocol: Introduction to the case study. Formal Aspects of Computing, 14(3):200–214, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650300001.

Abrial:2003:MPI

[416] Jean-Raymond Abrial, Dominique Cansell, and Dominique Méry. A mechanically proved and incremental development of IEEE 1394 Tree Identify Protocol. Formal Aspects of Computing, 14(3):215–227, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s001650300002.

Verdejo:2003:SVT

[417] Alberto Verdejo, Isabel Pita, and Narciso Martí-Oliet. Specification and verification of the tree identify protocol of IEEE 1394 in rewriting logic. Formal Aspects of Computing, 14(3):228–246, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650300003.

Calder:2003:USA

[418] M. Calder and A. Miller. Using SPIN to analyse the tree identification phase of the IEEE 1394 High-Performance Serial Bus (FireWire) Protocol. Formal Aspects of Computing, 14(3): 247–266, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650300004.

Schuppan:2003:VIF

[419] Viktor Schuppan and Armin Biere. Verifying the IEEE 1394 FireWire Tree Identify Protocol with SMV. Formal Aspects of Computing, 14 (3):267–280, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001650300005.

Fidge:2003:WID

[420] Colin Fidge and Carron Shankland. But what if I don’t want to wait forever? Formal Aspects of Computing, 14(3):281–294, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001650300006.

Kwiatkowska:2003:PMC

[421] Marta Kwiatkowska, Gethin Norman, and Jeremy Sproston. Probabilistic model checking of deadline properties in the IEEE 1394 FireWire Root Contention Protocol. Formal Aspects of Computing, 14(3):295–318, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001650300007.

Romijn:2003:FLD

[422] Judi Romijn. False loop detection in the IEEE 1394 Tree Identify Phase. Formal Aspects of Computing, 14 (3):319–327, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001650300008.

Stoelinga:2003:FFC

[423] Mariëlle Stoelinga. Fun with FireWire: A comparative study of formal verification methods applied to the IEEE 1394 Root Contention Protocol. Formal Aspects of Computing, 14(3):328–337, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s001650300009.

Anonymous:2003:Ea

[424] Anonymous. Editorial. Formal Aspects of Computing, 14(4):339, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00165-003-0242-8;   http://link.
springer.com/article/10.1007/s00165-
003-0242-8.

Anonymous:2003:O

[425] Anonymous.  Obituary.  For-

mal Aspects of Computing, 14(4): 340–341, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-003-0241-
9;   http://link.springer.com/article/
10.1007/s00165-003-0241-9.

Paun:2003:CUS

[426] Dimitrie O. Paun and Marsha Chechik. On closure under stuttering. Formal Aspects of Computing, 14(4):342–368, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-002-
225-1.

Kondoh:2003:ADT

[427] Hidetaka Kondoh. Abstract data types can have inequations. Formal Aspects of Computing, 14(4):369–399, April 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00165-002-
                                        0232-2.

Derrick:2003:SRS

[428] John Derrick and Graeme Smith. Structural refinement of systems specified in Object-Z and CSP. Formal Aspects of Computing, 15(1):1–27, July 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-003-
                                        0002-9.

Oliveira:2003:ATL

[429] Marcel Oliveira, Ana Cavalcanti, and Jim Woodcock. ArcAngel: a tactic language for refinement. Formal Aspects of Computing, 15(1):28–47, July 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-003-
                                        0003-8.

Henson:2003:LSB

[430] Martin C. Henson and Steve Reeves. A logic for schema-based program development. Formal Aspects of Computing, 15(1):48–83, July 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-003-0004-7.

Jifeng:2003:AFD

[431] He Jifeng and Xu Qiwen. Advanced features of duration calculus and their applications in sequential hybrid programs. Formal Aspects of Computing, 15(1):84–99, July 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00020897.

Anonymous:2003:Eb

[432] Anonymous.  Editorial.  Formal

Aspects of Computing, 15(2–3):101– 102, November 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-003-0023-
4;   http://link.springer.com/article/
10.1007/s00165-003-0023-4.

Back:2003:CAS

[433] R. J. R. Back and J. von Wright. Compositional action system refinement. Formal Aspects of Computing, 15(2– 3):103–117, November 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-003-0005-6.

Davies:2003:CRU

[434] Jim Davies and Charles Crichton. Concurrency and refinement in the unified modeling language. Formal Aspects of Computing, 15(2–3):118–145, November 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-003-
0008-3.

Cavalcanti:2003:RSC

[435] Ana Cavalcanti, Augusto Sampaio, and Jim Woodcock. A refinement strategy for circus. Formal Aspects of Computing, 15(2–3):146–181, November 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00165-003-
                                        0006-5.

Derrick:2003:RCR

[436] John Derrick and Eerke Boiten. Relational concurrent refinement. Formal Aspects of Computing, 15(2–3): 182–214, November 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-003-0007-4.

Sekerinski:2003:ETV

[437] Emil Sekerinski. Exploring tabular verification and refinement. Formal Aspects of Computing, 15(2–3): 215–236, November 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-003-0010-9.

Borger:2003:ARM

[438] Egon Börger. The ASM refinement method. Formal Aspects of Computing, 15(2–3):237–257, November 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-003-
                                        0012-7.

Wildman:2003:VVA

[439] Luke Wildman, Colin Fidge, and David Carrington. The variety of variables in automated real-time refinement. Formal Aspects of Computing, 15(2–3): 258–279, November 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-003-0009-2.

Aichernig:2003:MTR

[440] Bernhard K. Aichernig. Mutation testing in the refinement calculus. Formal Aspects of Computing, 15(2–3): 280–295, November 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-003-0011-8.

Broy:2003:EWT

[441] Manfred Broy, Gerald Lüttgen, and Michael Mendler. Editorial: Where theory and practice meet. Formal Aspects of Computing, 15(4): 297–298, December 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-003-0020-
7;   http://link.springer.com/article/
10.1007/s00165-003-0020-7.

Gay:2003:IES

[442] Simon Gay and Rajagopal Nagarajan. Intensional and extensional semantics of dataflow programs. Formal Aspects of Computing, 15(4):299–318, December 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-003-
0018-1.

Lermer:2003:LAE

[443] Karl Lermer, Colin J. Fidge, and Ian J. Hayes. Linear approximation of execution-time constraints. Formal Aspects of Computing, 15(4):319–348, December 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00165-003-
0019-0.

Janneck:2003:ATC

[444] Jörn W. Janneck. Actors and their composition. Formal Aspects of Computing, 15(4):349–369, December 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-003-
0016-3.

Bos:2003:RSE

[445] Victor Bos and Jeroen Kleijn. Redesign of a systems engineering language: Formalisation of χ. Formal Aspects of Computing, 15(4):370–389, December 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-003-
0017-2.

Franzle:2003:SDE

[446] Martin Fränzle, Jürgen Niehaus, Alexander Metzner, and Werner Damm. A semantics for distributed execution of statemate. Formal Aspects of Computing, 15(4):390–405, December 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-003-
0015-4.

Gordon:2003:VPS

[447] Michael J. C. Gordon. Validating the PSL/Sugar semantics using automated reasoning. Formal Aspects of Computing, 15(4):406–421, December 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00165-003-
0014-5.

Anonymous:2003:CV

[448] Anonymous. Content of volume 15 — 2003. Formal Aspects of Computing, 15(4):422, December 2003. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-003-0026-
1;   http://link.springer.com/article/
10.1007/s00165-003-0026-1.

Jones:2004:Ea

[449] Cliff Jones and John Cooke. Editorial. Formal Aspects of Computing, 16(1):1, April 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-004-0029-
6;   http://link.springer.com/article/
10.1007/s00165-004-0029-6.

Fitzgerald:2004:FME

[450] John S. Fitzgerald. Formal Methods Europe update. Formal Aspects of Computing, 16(1):2–3, April 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00165-004-0030-0;   http://link.
springer.com/article/10.1007/s00165-
004-0030-0.

Jones:2004:OFP

[451] C. B. Jones, D. J. Cooke, and Christiane Notarmarco. Online first publication. Formal Aspects of Computing, 16(1):4, April 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-004-0031-
z;   http://link.springer.com/article/
10.1007/s00165-004-0031-z.

Lopes:2004:SCV

[452] Antónia Lopes and José Luiz Fiadeiro. Superposition: composition vs refinement of non-deterministic, action-based systems. Formal Aspects of Computing, 16(1):5–18, April 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-003-
0021-6.

Martin:2004:DEG

[453] Clare Martin, Jeremy Gibbons, and Ian Bayley. Disciplined, efficient, generalised folds for nested datatypes. Formal Aspects of Computing, 16 (1):19–35, April 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-003-0013-6.

Stewart:2004:BSA

[454] A. Stewart, M. Clint, and J. Gabarró. Barrier synchronisation: Axiomatisation and relaxation. Formal Aspects of Computing, 16(1):36–50, April 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-004-
0028-7.

Paige:2004:EOO

[455] Richard F. Paige and Jonathan S. Ostroff. ERC — an object-oriented refinement calculus for Eiffel. Formal Aspects of Computing, 16(1):51–79, April 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-003-
0024-3.

Carr:2004:IUV

[456] C. T. Carr, T. M. McGinnity, and L. J. McDaid. Integration of UML and VHDL–AMS for analogue system modelling. Formal Aspects of Computing, 16(1):80–94, April 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-003-0027-0.

Jones:2004:Eb

[457] Cliff Jones and Michael R. Hansen. Editorial. Formal Aspects of Computing, 16(2):95, May 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-004-0047-
4;   http://link.springer.com/article/
10.1007/s00165-004-0047-4.

Barua:2004:CCN

[458] Rana Barua. Completeness of a combination of neighbourhood logic and temporal logic. Formal Aspects of Computing, 16(2):96–103, May 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-004-0035-8.

Dierks:2004:CMC

[459] Henning Dierks. Comparing model checking and logical reasoning for real-time systems. Formal Aspects of Computing, 16(2):104–120, May 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-004-
                                        0034-9.

Franzle:2004:MCD

[460] Martin Fränzle. Model-checking densetime duration calculus. Formal Aspects of Computing, 16(2):121–139, May 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-004-
                                        0032-y.

Liu:2004:UPM

[461] Zhiming Liu, Anders P. Ravn, and Xiaoshan Li. Unifying proof methodologies of duration calculus and timed linear temporal logic. Formal Aspects of Computing, 16(2):140–154, May 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-004-
                                        0036-7.

Xuandong:2004:DCR

[462] Li Xuandong, Zhao Jianhua, Zheng Tao, Li Yong, and Zheng Guoliang. Duration-constrained regular expressions. Formal Aspects of Computing, 16(2):155–163, May 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-004-0033-x.

Broy:2004:E

[463] Manfred Broy, Gerald Lüttgen, and Michael Mendler. Editorial. Formal Aspects of Computing, 16 (3):165, August 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-004-0049-
2;   http://link.springer.com/article/
10.1007/s00165-004-0049-2.

Gallardo:2004:GSP

[464] María del Mar Gallardo, Pedro Merino, and Ernesto Pimentel. A generalized semantics of PROMELA for abstract model checking. Formal Aspects of Computing, 16(3):166–193, August 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-004-
0040-y.

Kruger:2004:BM

[465] I. Krüger, W. Prenninger, and R. Sandner. Broadcast MSCs. Formal Aspects of Computing, 16(3):194–209, August 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-004-
0046-5.

Lee:2004:BTS

[466] Edward A. Lee and Yuhong Xiong. A behavioral type system and its application in Ptolemy II. Formal Aspects of Computing, 16(3):210–237, August 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-004-
0043-8.

Lopez:2004:IFP

[467] Natalia López, Manuel Núñez, and Fernando Rubio. An integrated framework for the performance analysis of asynchronous communicating stochastic processes. Formal Aspects of Computing, 16(3):238–262, August 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-004-
                                        0044-7.

Nebut:2004:SAS

[468] Mirabelle Nebut. Specification and analysis of synchronous reactions. Formal Aspects of Computing, 16(3): 263–291, August 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-004-0039-4.

Tini:2004:TCC

[469] Simone Tini. Timed CCP compositionally embeds Argos and Lustre. Formal Aspects of Computing, 16(3):292–312, August 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-004-
                                        0041-x.

Hallerstede:2004:PAP

[470] Stefan Hallerstede and Michael Butler. Performance analysis of probabilistic action systems. Formal Aspects of Computing, 16(4):313–331, November 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-004-
                                        0037-6.

Paynter:2004:RMB

[471] S. E. Paynter, N. Henderson, and J. M. Armstrong. Ramifications of metastability in bit variables explored via Simpson’s 4-slot mechanism. Formal Aspects of Computing, 16(4): 332–351, November 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-004-0042-9.

Coupet-Grimal:2004:CCT

[472] Solange Coupet-Grimal and Line Jakubiec. Certifying circuits in Type Theory. Formal Aspects of Computing, 16 (4):352–373, November 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-004-0048-3.

Ipate:2004:CDS

[473] Florentin Ipate. Complete deterministic stream X-machine testing. Formal Aspects of Computing, 16(4): 374–386, November 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-004-0045-6.

Hesselink:2004:APC

[474] Wim H. Hesselink. An assertional proof for a construction of an atomic variable. Formal Aspects of Computing, 16 (4):387–393, November 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-004-0038-5.

Reed:2004:RIC

[475] J. N. Reed, J. E. Sinclair, and A. W. Roscoe. Responsiveness of interoperating components. Formal Aspects of Computing, 16(4):394–411, November 2004. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-004-
                                        0050-9.

Herzberg:2005:MLD

[476] D. Herzberg and M. Broy. Modeling layered distributed communication systems. Formal Aspects of Computing, 17(1):1–18, May 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-004-0051-8.

Groote:2005:CCA

[477] Jan Friso Groote, François Monin, and Jan Springintveld. A computer checked algebraic verification of a distributed summation algorithm. Formal Aspects of Computing, 17(1):19–37, May 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-004-
                                        0052-7.

Cook:2005:DAH

[478] A. Cook, A. Ireland, G. Michaelson, and N. Scaife. Discovering applications of higher order functions through proof planning. Formal Aspects of Computing, 17(1):38–57, May 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-004-0054-5.

Ehmety:2005:MCR

[479] Sidi O. Ehmety and Lawrence C. Paulson. Mechanizing compositional reasoning for concurrent systems: some lessons. Formal Aspects of Computing, 17(1):58–68, May 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-004-0053-6.

Back:2005:ATP

[480] Ralph-Johan Back and Viorel Preoteasa. An algebraic treatment of procedure refinement to support mechanical verification. Formal Aspects of Computing, 17(1):69–90, May 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-004-0060-7.

Leuschel:2005:GE

[481] Michael Leuschel. Guest editorial. Formal Aspects of Computing, 17 (2):91–92, August 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-005-0069-
6;   http://link.springer.com/article/
10.1007/s00165-005-0069-6.

Roscoe:2005:EPC

[482] A. W. Roscoe. On the expressive power of CSP refinement. Formal Aspects of Computing, 17(2):93–112, August 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0065-x.

Huth:2005:RCI

[483] Michael Huth. Refinement is complete for implementations. Formal Aspects of Computing, 17(2):113–137, August 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00165-005-
                                        0063-z.

Cattani:2005:RBP

[484] Stefano Cattani and Marta Kwiatkowska. A refinement-based process algebra for timed automata. Formal Aspects of Computing, 17(2):138–159, August 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-005-
                                        0064-y.

Norman:2005:UPM

[485] Gethin Norman, David Parker, Marta Kwiatkowska, Sandeep Shukla, and Rajesh Gupta. Using probabilistic model checking for dynamic power management. Formal Aspects of Computing, 17(2):160–176, August 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-005-
                                        0062-0.

DeWachter:2005:FDD

[486] Bram De Wachter, Alexandre Genon, Thierry Massart, and Cédric Meuter. The formal design of distributed controllers with _{d}SL and Spin. Formal Aspects of Computing, 17(2):177–200, August 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-005-
                                        0066-9.

Gunter:2005:MCT

[487] Elsa Gunter and Doron Peled. Model checking, testing and verification working together. Formal Aspects of Computing, 17(2):201–221, August 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0059-8.

Bagnara:2005:NCC

[488] Roberto Bagnara, Patricia M. Hill, and Enea Zaffanella. Not necessarily closed convex polyhedra and the double description method. Formal Aspects of Computing, 17(2):222–257, August 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0061-1.

Cooke:2005:E

[489] John Cooke.  Editorial.  For-

mal Aspects of Computing, 17(3): 259, October 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-005-0074-
9;   http://link.springer.com/article/
10.1007/s00165-005-0074-9.

Gurgens:2005:SAE

[490] S. Gürgens and C. Rudolph. Security analysis of efficient (un-) fair nonrepudiation protocols. Formal Aspects of Computing, 17(3):260–276, October 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-004-
0055-4.

Gordon:2005:VWS

[491] Andrew D. Gordon and Riccardo Pucella. Validating a web service security abstraction by typing. Formal Aspects of Computing, 17(3):277–318, October 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-004-
0058-1.

DeWulf:2005:AAS

[492] Martin De Wulf, Laurent Doyen, and Jean-François Raskin. Almost ASAP semantics: from timed models to timed implementations. Formal Aspects of Computing, 17(3):319–341, October 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0067-8.

Badban:2005:VSW

[493] Bahareh Badban, Wan Fokkink, Jan Friso Groote, Jun Pang, and Jaco van de Pol. Verification of a sliding window protocol in μCRL and PVS. Formal Aspects of Computing, 17(3):342–388, October 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-005-0070-0.

Boiten:2005:GEI

[494] E. Boiten, J. Derrick, and G. Smith. Guest editorial integrated formal methods. Formal Aspects of Computing, 17(4):389, December 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-005-0078-
5;   http://link.springer.com/article/
10.1007/s00165-005-0078-5.

Schneider:2005:CTC

[495] Steve Schneider and Helen Treharne. CSP theorems for communicating B machines. Formal Aspects of Computing, 17(4):390–422, December 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0076-7.

Beckert:2005:RRP

[496] Bernhard Beckert and Steffen Schlager. Refinement and retrenchment for programming language data types. Formal Aspects of Computing, 17(4): 423–442, December 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-005-0073-x.

He:2005:EOS

[497] Jifeng He, Dang Van Hung, Geguang Pu, Zongyan Qiu, and Wang Yi. Exploring optimal solution to hardware/software partitioning for synchronous model. Formal Aspects of Computing, 17(4):443–460, December 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0072-y.

Chaki:2005:CSV

[498] Sagar Chaki, Edmund Clarke, Joël Ouaknine, Natasha Sharygina, and Nishant Sinha. Concurrent software verification with states, events, and deadlocks. Formal Aspects of Computing, 17(4):461–483, December 2005. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-005-
                                        0071-z.

Derrick:2006:GE

[499] J. Derrick, M. Harman, and R. M. Hierons. Guest editorial. Formal Aspects of Computing, 18(1):1–2, March 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00165-005-0087-4;   http://link.
                                        springer.com/article/10.1007/s00165-
                                        005-0087-4.

Bogdanov:2006:TMX

[500] K. Bogdanov, M. Holcombe, F. Ipate, L. Seed, and S. Vanak. Testing methods for X-machines: a review. Formal Aspects of Computing, 18(1):3–30, March 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-005-
                                        0085-6.

Stannett:2006:STA

[501] Mike Stannett. Simulation testing of automata. Formal Aspects of Computing, 18(1):31–41, March 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-005-0080-y.

Vilkomir:2006:MDR

[502] Sergiy A. Vilkomir and Jonathan P. Bowen. From MC/ DC to RC/ DC: formalization and analysis of controlflow testing criteria. Formal Aspects of Computing, 18(1):42–62, March 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0084-7.

Hallal:2006:FAP

[503] H. H. Hallal, S. Boroday, A. Petrenko, and A. Ulrich. A formal approach to property testing in causally consistent distributed traces. Formal Aspects of Computing, 18(1):63–83, March 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0082-9.

Ural:2006:CCS

[504] Hasan Ural and Craig Williams. Constructing checking sequences for distributed testing. Formal Aspects of Computing, 18(1):84–101, March 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0083-8.

Binkley:2006:FRB

[505] David W. Binkley, Sebastian Danicic, Mark Harman, John Howroyd, and Lahcen Ouarbya. A formal relationship between program slicing and partial evaluation. Formal Aspects of Computing, 18(2):103–119, June 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-005-
0077-6.

McKeever:2006:PCH

[506] Steve McKeever and Wayne Luk. Provably-correct hardware compilation tools based on pass separation techniques. Formal Aspects of Computing, 18(2):120–142, June 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-005-0075-8.

Bicarregui:2006:VSR

[507] J. C. Bicarregui, C. A. R. Hoare, and J. C. P. Woodcock. The verified software repository: a step towards the verifying compiler. Formal Aspects of Computing, 18(2):143–151, June 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-005-
                                        0079-4.

Lermer:2006:PCR

[508] K. Lermer and C. J. Fidge. Procedure compilation in the refinement calculus. Formal Aspects of Computing, 18(2):152–180, June 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-006-0091-3.

Bolton:2006:SFS

[509] Christie Bolton and Jim Davies. A singleton failures semantics for Communicating Sequential Processes. Formal Aspects of Computing, 18(2):181–210, June 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-005-
                                        0081-x.

Fecher:2006:ARA

[510] H. Fecher and M. Majster-Cederbaum. Action refinement applied to late decisions. Formal Aspects of Computing, 18(2):211–230, June 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-006-0088-y.

Santiago:2006:FAC

[511] Regivan H. Nunes Santiago, Benjamín R. Callejas Bedregal, and Benedito Melo Acióly. Formal aspects of correctness and optimality of interval computations. Formal Aspects of Computing, 18(2):231–243, June 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-006-
0089-x.

Glesner:2006:FIC

[512] Sabine Glesner. Finite integer computations: An algebraic foundation for their correctness. Formal Aspects of Computing, 18(2):244–262, June 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-006-
0090-4.

Boiten:2006:GEE

[513] Eerke Boiten and Michael Butler. Guest Editorial: Editorial for the FAC special issue based on derivative papers from “Refine ’05”. Formal Aspects of Computing, 18(3): 263, September 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-006-0009-
0;   http://link.springer.com/article/
10.1007/s00165-006-0009-0.

Smith:2006:VDR

[514] Graeme Smith and John Derrick. Verifying data refinements using a model checker. Formal Aspects of Computing, 18(3):264–287, September 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-006-
                                        0002-7.

Cavalcanti:2006:ANU

[515] Ana Cavalcanti, Jim Woodcock, and Steve Dunne. Angelic nondeterminism in the unifying theories of programming. Formal Aspects of Computing, 18 (3):288–307, September 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-006-0001-8.

Schneider:2006:TMP

[516] Steve Schneider, Thai Son Hoang, Ken Robinson, and Helen Treharne. Tank monitoring: a pAMN case study. Formal Aspects of Computing, 18(3): 308–328, September 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-006-0004-5.

Deutsch:2006:ARA

[517] Moshe Deutsch and Martin C. Henson. An analysis of refinement in an abortive paradigm. Formal Aspects of Computing, 18(3):329–363, September 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-006-
                                        0006-3.

Henson:2006:SL

[518] Martin C. Henson, Moshe Deutsch, and Besnik Kajtazi. The specification logic νZ. Formal Aspects of Computing, 18 (3):364–395, September 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-006-0007-2.

Tenzer:2006:MRC

[519] Jennifer Tenzer and Perdita Stevens. On modelling recursive calls and callbacks with two variants of unified modelling language state diagrams. Formal Aspects of Computing, 18(4): 397–420, December 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-006-0003-6.

Subramani:2006:AAS

[520] K. Subramani and C. Tauras. An approximation algorithm for state minimization in 2-MDFAs. Formal Aspects of Computing, 18(4):421–431, December 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-006-
0005-4.

Vaandrager:2006:ABM

[521] F. W. Vaandrager and A. L. de Groot. Analysis of a biphase mark protocol with Uppaal and PVS. Formal Aspects of Computing, 18(4):433–458, December 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-006-
0008-1.

Bowman:2006:HST

[522] Howard Bowman and Rodolfo Gómez. How to stop time stopping. Formal Aspects of Computing, 18(4):459–493, December 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-006-
                                        0010-7.

Pym:2006:CLR

[523] David Pym and Chris Tofts. A calculus and logic of resources and processes. Formal Aspects of Computing, 18(4): 495–517, December 2006. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-006-0018-z. See errata

[553].

Cooke:2007:E

[524] John Cooke. Editorial. Formal Aspects of Computing, 19(1):1, March 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00165-007-0025-8;   http://link.
                                        springer.com/article/10.1007/s00165-
                                        007-0025-8.

Hillston:2007:FTP

[525] Jane Hillston and Leïla Kloul. Formal techniques for performance analysis: blending SAN and PEPA. Formal Aspects of Computing, 19(1):3–33, March 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-006-
                                        0011-6.

Gorogiannis:2007:MRSa

[526] Nikos Gorogiannis and Mark Ryan. Minimal refinements of specifications in model and temporal logics. Formal Aspects of Computing, 19(1):35–62, March 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-006-
0014-3.

Basin:2007:VSA

[527] David Basin, Hironobu Kuruma, Kunihiko Miyazaki, Kazuo Takaragi, and Burkhart Wolff. Verifying a signature architecture: a comparative case study. Formal Aspects of Computing, 19(1):63–91, March 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-006-0012-5.

Lanotte:2007:PPT

[528] Ruggero Lanotte, Andrea Maggiolo-Schettini, and Angelo Troina. Parametric probabilistic transition systems for system design and analysis. Formal Aspects of Computing, 19(1):93–109, March 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-006-
0015-2.

Burns:2007:EPV

[529] A. Burns and T.-M. Lin. An engineering process for the verification of real-time systems. Formal Aspects of Computing, 19(1):111–136, March 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00165-006-
                                        0021-4.

Cooke:2007:EVS

[530] J. Cooke. Editorial (VSTTE Special Issue). Formal Aspects of Computing, 19(2):137–138, June 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00165-007-0036-
                                        5;   http://link.springer.com/article/
                                        10.1007/s00165-007-0036-5.

Chalin:2007:LFV

[531] Patrice Chalin. Are the logical foundations of verifying compiler prototypes matching user expectations? Formal Aspects of Computing, 19(2):139–158, June 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-006-
                                        0016-1.

Leavens:2007:SVC

[532] Gary T. Leavens, K. Rustan M. Leino, and Peter Müller. Specification and verification challenges for sequential object-oriented programs. Formal Aspects of Computing, 19(2):159–189, June 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-007-
                                        0026-7.

Jacobs:2007:CCT

[533] Bart Jacobs, Sjaak Smetsers, and Ronny Wichers Schreur. Code-carrying theories. Formal Aspects of Computing, 19(2):191–203, June 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-006-0013-4.

Naumann:2007:ABE

[534] David A. Naumann. On assertionbased encapsulation for object invariants and simulations. Formal Aspects of Computing, 19(2):205–224, June 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-006-
0020-5.

Borger:2007:CAG

[535] Egon Börger. Construction and analysis of ground models and their refinements as a foundation for validating computer-based systems. Formal Aspects of Computing, 19(2):225–241, June 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-006-
0019-y.

Betin-Can:2007:HDC

[536] Aysu Betin-Can and Tevfik Bultan. Highly dependable concurrent programming using design for verification. Formal Aspects of Computing, 19(2):243–268, June 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-006-0017-0.

Joshi:2007:MCB

[537] Rajeev Joshi and Gerard J. Holzmann. A mini challenge: build a verifiable filesystem. Formal Aspects of Computing, 19(2):269–272, June 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-006-
0022-3.

Gorogiannis:2007:MRSb

[538] Nikos Gorogiannis and Mark Ryan. Minimal refinements of specifications in modal and temporal logics. Formal Aspects of Computing, 19(2):273, June 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-007-0038-
3;   http://link.springer.com/article/
10.1007/s00165-007-0038-3.

Lazic:2007:GE

[539] Ranko Lazic and Rajagopal Nagarajan. Guest editorial. Formal Aspects of Computing, 19(3):275, August 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00165-007-0043-6;   http://link.
springer.com/article/10.1007/s00165-
007-0043-6.

Evans:2007:ITS

[540] Neil Evans and Helen Treharne. Interactive tool support for CSP —— B consistency checking. Formal Aspects of Computing, 19(3):277–302, August 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0034-7.

Reed:2007:RSR

[541] J. N. Reed, A. W. Roscoe, and J. E. Sinclair. Responsiveness and stable revivals. Formal Aspects of Computing, 19(3):303–319, August 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-007-0032-9.

Barsotti:2007:VCS

[542] Damián Barsotti, Leonor Prensa Nieto, and Alwen Tiu. Verification of clock synchronization algorithms: experiments on a combination of deductive tools. Formal Aspects of Computing, 19(3):321–341, August 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-007-0027-6.

Slind:2007:PPS

[543] Konrad Slind, Scott Owens, Juliano Iyoda, and Mike Gordon. Proof producing synthesis of arithmetic and cryptographic hardware. Formal Aspects of Computing, 19(3):343–362, August 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0028-5.

Ceska:2007:GMP

[544] Milan Češka, Pavel Erlebach, and Tomáš Vojnar. Generalised multipattern-based verification of programs with linear linked structures. Formal Aspects of Computing, 19(3):363–374, August 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00165-007-
                                        0031-x.

Charlton:2007:PVI

[545] Nathaniel Charlton. Program verification with interacting analysis plugins. Formal Aspects of Computing, 19 (3):375–399, August 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-007-0029-4.

Kang:2007:PDV

[546] Eun-Young Kang and Stephan Merz. Predicate diagrams for the verification of real-time systems. Formal Aspects of Computing, 19(3):401–413, August 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-007-
                                        0030-y.

Jones:2007:E

[547] Cliff Jones and Jim Woodcock. Editorial. Formal Aspects of Computing, 19 (4):415–416, November 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00165-007-0046-
                                        3;   http://link.springer.com/article/
                                        10.1007/s00165-007-0046-3.

Gorogiannis:2007:MRSc

[548] Nikos Gorogiannis and Mark Ryan. Minimal refinements of specifications in modal and temporal logics. Formal Aspects of Computing, 19(4): 417–444, November 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-007-0040-9.

Bergstra:2007:TAS

[549] J. A. Bergstra and C. A. Middelburg. Thread algebra for strategic interleaving. Formal Aspects of Computing, 19 (4):445–474, November 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-007-0024-9;   http:
//link.springer.com/content/pdf/
10.1007/s00165-007-0024-9.pdf.

Vu:2007:DOB

[550] Thuy Duong Vu. Deciding orthogonal bisimulation. Formal Aspects of Computing, 19(4):475–485, November 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0023-x;   http://link.springer.
com/content/pdf/10.1007/s00165-
007-0023-x.pdf.

Brooke:2007:CME

[551] Phillip J. Brooke, Richard F. Paige, and Jeremy L. Jacob. A CSP model of Eiffel’s SCOOP. Formal Aspects of Computing, 19(4):487–512, November 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0033-8.

Curzon:2007:AFV

[552] Paul Curzon, Rimvydas Rukšėnas, and Ann Blandford. An approach to formal verification of human— computer interaction. Formal Aspects of Computing, 19(4):513–550, November 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-007-
                                        0035-6.

Collinson:2007:EBA

[553] Matthew Collinson, David Pym, and Chris Tofts. Errata for Formal Aspects of Computing (2006) 18:495–517 and their consequences. Formal Aspects of Computing, 19(4):551–554, November 2007. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-007-

0047-2. See [523].

Jones:2008:E

[554] Cliff Jones and Jim Woodcock. Editorial. Formal Aspects of Computing, 20(1):1–3, January 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-007-0064-1.

Woodcock:2008:CME

[555] Jim Woodcock, Susan Stepney, David Cooper, John Clark, and Jeremy Jacob. The certification of the Mondex electronic purse to ITSEC Level E6. Formal Aspects of Computing, 20(1):5–19, January 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-007-0060-5.

Ramananandro:2008:MEP

[556] Tahina Ramananandro. Mondex, an electronic purse: specification and refinement checks with the Alloy modelfinding method. Formal Aspects of Computing, 20(1):21–39, January 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0058-z.

Haneberg:2008:VME

[557] Dominik Haneberg, Gerhard Schellhorn, Holger Grandy, and Wolfgang Reif. Verification of Mondex electronic purses with KIV: from transactions to a security protocol. Formal Aspects of Computing, 20(1):41–59, January 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0057-0.

Butler:2008:IDM

[558] Michael Butler and Divakar Yadav. An incremental development of the Mondex system in Event-B. Formal Aspects of Computing, 20(1):61–77, January 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0061-4.

Kuhlmann:2008:MVM

[559] Mirco Kuhlmann and Martin Gogolla. Modeling and validating Mondex scenarios described in UML and OCL with USE. Formal Aspects of Computing, 20(1):79–100, January 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-007-0062-3.

George:2008:SPM

[560] Chris George and Anne E. Haxthausen. Specification, proof, and model checking of the Mondex electronic purse using RAISE. Formal Aspects of Computing, 20(1):101–116, January 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-007-
                                        0054-3.

Freitas:2008:MMZ

[561] Leo Freitas and Jim Woodcock. Mechanising Mondex with Z/Eves. Formal Aspects of Computing, 20(1): 117–139, January 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-007-0059-y.

Dongol:2008:SPB

[562] Brijesh Dongol and Arjan J. Mooij. Streamlining progress-based derivations of concurrent programs. Formal Aspects of Computing, 20(2):141–160, March 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-007-
                                        0037-4.

Moller:2008:IFM

[563] Michael Möller, Ernst-Rüdiger Olderog, Holger Rasch, and Heike Wehrheim. Integrating a formal method into a software engineering process with UML and Java. Formal Aspects of Computing, 20(2):161–204, March 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00165-007-
0042-7.

Bhaduri:2008:ISP

[564] Purandar Bhaduri and S. Ramesh. Interface synthesis and protocol conversion. Formal Aspects of Computing, 20(2):205–224, March 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-007-0045-4.

Boucheneb:2008:ITC

[565] Hanifa Boucheneb. Interval timed coloured Petri net: efficient construction of its state class space preserving linear properties. Formal Aspects of Computing, 20(2):225–238, March 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0050-7.

Bjorner:2008:JWB

[566] Dines Bjørner. John Warner Backus: 3 Dec 1924–17 March 2007. Formal Aspects of Computing, 20(3): 239–240, May 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-008-0077-
4;   http://link.springer.com/article/
10.1007/s00165-008-0077-4.

Schmaltz:2008:FFC

[567] Julien Schmaltz and Dominique Borrione. A functional formalization of on chip communications. Formal Aspects of Computing, 20(3):241–258, May 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0049-0.

Kammuller:2008:FNI

[568] Florian Kammüller. Formalizing noninterference for a simple bytecode language in Coq. Formal Aspects of Computing, 20(3):259–275, May 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0055-2.

Lowe:2008:SCP

[569] Gavin Lowe. Specification of communicating processes: temporal logic versus refusals-based refinement. Formal Aspects of Computing, 20(3):277–294, May 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0065-0.

Reeves:2008:DRS

[570] Steve Reeves and David Streader. Data refinement and singleton failures refinement are not equivalent. Formal Aspects of Computing, 20(3):295–301, May 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0076-5.

Bertolotti:2008:ERA

[571] Ivan Cibrario Bertolotti, Luca Durante, Riccardo Sisto, and Adriano Valenzano. Efficient representation of the attacker’s knowledge in cryptographic protocols analysis. Formal Aspects of Computing, 20(3):303–348, May 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0078-3.

Barkaoui:2008:GE

[572] Kamel Barkaoui, Manfred Broy, Ana Cavalcanti, and Antonio Cerone. Guest editorial. Formal Aspects of Computing, 20(4–5):349–350, July 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00165-008-0085-4;   http://link.
springer.com/article/10.1007/s00165-
008-0085-4.

Barbuti:2008:BCM

[573] Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Angelo Troina. Bisimulations in calculi modelling membranes. Formal Aspects of Computing, 20(4–5):351–377, July 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0071-x.

Bradley:2008:PDI

[574] Aaron R. Bradley and Zohar Manna. Property-directed incremental invariant generation. Formal Aspects of Computing, 20(4–5):379–405, July 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0080-9.

Delzanno:2008:RAF

[575] Giorgio Delzanno and Roberto Montagna. Reachability analysis of fragments of mobile ambients in AC term rewriting. Formal Aspects of Computing, 20(4–5):407–428, July 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-008-
                                        0081-8.

Devillers:2008:CPN

[576] Raymond Devillers, Hanna Klaudel, and Maciej Koutny. A compositional Petri net translation of general π-calculus terms. Formal Aspects of Computing, 20(4–5):429–450, July 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-008-
                                        0079-2.

Gabbay:2008:CAS

[577] Murdoch J. Gabbay and Aad Mathijssen. Capture-avoiding substitution as a nominal algebra. Formal Aspects of Computing, 20(4–5):451–479, July 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-007-
                                        0056-1.

Meyer:2008:MCD

[578] Roland Meyer, Johannes Faber, Jochen Hoenicke, and Andrey Rybalchenko. Model checking duration calculus: a practical approach. Formal Aspects of Computing, 20(4–5):481–505, July 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0082-7;   http://link.springer.
com/content/pdf/10.1007/s00165-
008-0082-7.pdf.

Slanina:2008:DVA

[579] Matteo Slanina, Henny B. Sipma, and Zohar Manna. Deductive verification of alternating systems. Formal Aspects of Computing, 20(4–5):507–560, July 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0075-6.

Jones:2008:V

[580] Cliff Jones.  Valediction.  For-

mal Aspects of Computing, 20(6): 561, December 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-008-0093-
4;   http://link.springer.com/article/
10.1007/s00165-008-0093-4.

Labbe:2008:SCA

[581] Sébastien Labbé and Jean-Pierre Gallois. Slicing communicating automata specifications: polynomial algorithms for model reduction. Formal Aspects of Computing, 20(6):563–595, December 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0086-3.

Hierons:2008:TDI

[582] Robert M. Hierons and Florentin Ipate. Testing a deterministic implementation against a non-controllable nondeterministic stream X-machine. Formal Aspects of Computing, 20(6): 597–617, December 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-008-0087-2.

Grinchtein:2008:NIR

[583] Olga Grinchtein and Martin Leucker. Network invariants for real-time systems. Formal Aspects of Computing, 20 (6):619–635, December 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-008-0089-0.

Botaschanjan:2008:CUL

[584] Jewgenij Botaschanjan, Manfred Broy, Alexander Gruler, Alexander Harhurin, Steffen Knapp, Leonid Kof, Wolfgang Paul, and Maria Spichkova. On the correctness of upper layers of automotive systems. Formal Aspects of Computing, 20(6):637–662, December 2008. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0097-0.

Boiten:2009:E

[585] Eerke Boiten. Editorial. Formal Aspects of Computing, 21(1–2):1, February 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0100-9;   http://link.springer.
com/content/pdf/10.1007/s00165-
008-0100-9.pdf.

Oliveira:2009:USC

[586] Marcel Oliveira, Ana Cavalcanti, and Jim Woodcock. A UTP semantics for circus. Formal Aspects of Computing, 21(1–2):3–32, February 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-007-0052-5.

Aichernig:2009:MTU

[587] Bernhard K. Aichernig and He Jifeng. Mutation testing in UTP. Formal Aspects of Computing, 21(1– 2):33–64, February 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-008-0083-6.

Boiten:2009:RCR

[588] Eerke Boiten, John Derrick, and Gerhard Schellhorn. Relational concurrent refinement part II: Internal operations and outputs. Formal Aspects of Computing, 21(1–2):65–102, February 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0066-z.

Zhao:2009:GTO

[589] Liang Zhao, Xiaojian Liu, Zhiming Liu, and Zongyan Qiu. Graph transformations for object-oriented refinement. Formal Aspects of Computing, 21(1– 2):103–131, February 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-007-0067-y.

Freitas:2009:FE

[590] Leo Freitas and Jim Woodcock. FDR explorer. Formal Aspects of Computing, 21(1–2):133–154, February 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0074-7.

Smith:2009:MCA

[591] Graeme Smith and Kirsten Winter. Model checking action system refinements. Formal Aspects of Computing, 21(1–2):155–186, February 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0053-4.

Groves:2009:TBD

[592] Lindsay Groves and Robert Colvin. Trace-based derivation of a scalable lock-free stack algorithm. Formal Aspects of Computing, 21(1–2): 187–223, February 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-008-0092-5.

Boca:2009:E

[593] Paul Boca, Raymond Boute, David Duce, and José Oliveira. Editorial. Formal Aspects of Computing, 21(3):225, May 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0109-8;   http://link.springer.
com/content/pdf/10.1007/s00165-
009-0109-8.pdf.

Back:2009:IBP

[594] Ralph-Johan Back. Invariant based programming: basic approach and teaching experiences. Formal Aspects of Computing, 21(3):227–244, May 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0070-y.

Larsen:2009:POC

[595] Peter Gorm Larsen, John S. Fitzgerald, and Steve Riddle. Practice-oriented courses in formal methods using VDM^{++}. Formal Aspects of Computing, 21(3):245–257, May 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-008-0068-5.

Tsay:2009:TSL

[596] Yih-Kuen Tsay, Yu-Fang Chen, Ming-Hsien Tsai, Kang-Nien Wu, Wen-Chin Chan, Chi-Jian Luo, and Jinn-Shu Chang. Tool support for learning Büchi automata and linear temporal logic. Formal Aspects of Computing, 21(3):259–275, May 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-008-0091-6.

Schreiner:2009:RPP

[597] Wolfgang Schreiner. The RISC Proof-Navigator: a proving assistant for program verification in the classroom. Formal Aspects of Computing, 21 (3):277–291, May 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-008-0069-4.

Feinerer:2009:CTT

[598] Ingo Feinerer and Gernot Salzer. A comparison of tools for teaching formal software verification. Formal Aspects of Computing, 21(3):293–301, May 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0084-5.

Paige:2009:E

[599] Richard F. Paige and Phillip J. Brooke. Editorial. Formal Aspects of Computing, 21(4):303, August 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0113-z;   http:
//link.springer.com/content/pdf/
10.1007/s00165-009-0113-z.pdf.

Nienaltowski:2009:CC

[600] Piotr Nienaltowski, Bertrand Meyer, and Jonathan S. Ostroff. Contracts for concurrency. Formal Aspects of Computing, 21(4):305–318, August 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0063-2.

Ostroff:2009:BCC

[601] Jonathan S. Ostroff, Faraz Ahmadi Torshizi, Hai Feng Huang, and Bernd Schoeller. Beyond contracts for concurrency. Formal Aspects of Computing, 21(4):319–346, August 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-008-0073-8.

Nienaltowski:2009:FAC

[602] Piotr Nienaltowski. Flexible access control policy for SCOOP. Formal Aspects of Computing, 21(4):347–362, August 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0072-9.

Brooke:2009:CAM

[603] Phillip J. Brooke and Richard F. Paige. Cameo: an alternative model of concurrency for Eiffel. Formal Aspects of Computing, 21(4):363–391, August 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0096-1.

Bornat:2009:PLC

[604] Richard Bornat. Peter Landin: a computer scientist who inspired a generation, 5th June 1930–3rd June 2009. Formal Aspects of Computing, 21(5): 393–395, October 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0122-y.

Kim:2009:FDD

[605] Soon-Kyeong Kim and David Carrington. A formalism to describe design patterns based on role concepts. Formal Aspects of Computing, 21(5):397–420, October 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-008-0090-7.

Howells:2009:CMF

[606] Paul Howells and Mark d’Inverno. A CSP model with flexible parallel termination semantics. Formal Aspects of Computing, 21(5):421–449, October 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0098-z.

Chen:2009:FFM

[607] Chunqing Chen, Jin Song Dong, and Jun Sun. A formal framework for modeling and validating Simulink diagrams. Formal Aspects of Computing, 21(5):451–483, October 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0108-9.

Kroening:2009:FSM

[608] Daniel Kroening and Ofer Strichman. A framework for Satisfiability Modulo Theories. Formal Aspects of Computing, 21(5):485–494, October 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0105-z.

Masalagiu:2009:RMS

[609] Cristian Masalagiu, Wei-Ngan Chin, Ştefan Andrei, and Vasile Alaiba. A rigorous methodology for specification and verification of business processes. Formal Aspects of Computing, 21(5):495–510, October 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0106-y.

Cerone:2009:E

[610] Antonio Cerone, Paul Curzon, and David Duce. Editorial. Formal Aspects of Computing, 21(6): 511–512, December 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-009-0121-
z;   http://link.springer.com/article/
10.1007/s00165-009-0121-z.

Su:2009:PAM

[611] Li Su, Howard Bowman, Philip Barnard, and Brad Wyble. Process algebraic modelling of attentional capture and human electrophysiology in interactive systems. Formal Aspects of Computing, 21(6):513–539, December 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0094-3;   http://link.springer.
com/content/pdf/10.1007/s00165-
008-0094-3.pdf.

Ruksenas:2009:VGM

[612] Rimvydas Ruksenas, Jonathan Back, Paul Curzon, and Ann Blandford. Verification-guided modelling of salience and cognitive load. Formal Aspects of Computing, 21(6):541–569, December 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-008-
0102-7.

Basuki:2009:MCU

[613] Thomas Anung Basuki, Antonio Cerone, Andreas Griesmayer, and Rudolf Schlatte. Model-checking user behaviour using interacting components. Formal Aspects of Computing, 21(6):571–588, December 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00165-008-0101-
                                        8.

Bowen:2009:RUI

[614] Judy Bowen and Steve Reeves. Refinement for user interface designs. Formal Aspects of Computing, 21(6): 589–612, December 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-008-0095-2.

Dix:2009:PMD

[615] Alan Dix, Masitah Ghazali, Steve Gill, Joanna Hare, and Devina Ramduny-Ellis. Physigrams: modelling devices for natural interaction. Formal Aspects of Computing, 21(6):613–641, December 2009. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-008-
                                        0099-y.

Boiten:2010:E

[616] Eerke Boiten, Michael Butler, John Derrick, and Graeme Smith. Editorial. Formal Aspects of Computing, 22(1):1, January 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.

1007/s00165-009-0147-2;   http:
//link.springer.com/content/pdf/
10.1007/s00165-009-0147-2.pdf.

Meinicke:2010:RAP

[617] Larissa Meinicke and Kim Solin. Refinement algebra for probabilistic programs. Formal Aspects of Computing, 22(1):3–31, January 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0111-1.

Banach:2010:AAT

[618] Richard Banach and Gerhard Schellhorn. Atomic actions, and their refinements to isolated protocols. Formal Aspects of Computing, 22(1):33–61, January 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0103-1.

Mooij:2010:IBR

[619] Arjan J. Mooij. Invariant-based reasoning about parameterized security protocols. Formal Aspects of Computing, 22(1):63–81, January 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0104-0;   http:
//link.springer.com/content/pdf/
10.1007/s00165-009-0104-0.pdf.

Collinson:2010:ALA

[620] Matthew Collinson and David Pym. Algebra and logic for access control. Formal Aspects of Computing, 22(2):83–104, March 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0107-x. See erra-

tum [636].

Kroening:2010:VFP

[621] Daniel Kroening and Georg Weissenbacher. Verification and falsification of programs with loops using predicate abstraction. Formal Aspects of Computing, 22(2):105–128, March 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0110-2.

Mathis:2010:FGC

[622] Pascal Mathis and Simon E. B. Thierry. A formalization of geometric constraint systems and their decomposition. Formal Aspects of Computing, 22(2):129–151, March 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0117-8.

Sherif:2010:PAF

[623] Adnan Sherif, Ana Cavalcanti, He Jifeng, and Augusto Sampaio. A process algebraic framework for specification and validation of real-time systems. Formal Aspects of Computing, 22(2):153–191, March 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0119-6.

Aransay:2010:GCC

[624] Jesús Aransay, Clemens Ballarin, and Julio Rubio. Generating certified code from formal proofs: a case study in homological algebra. Formal Aspects of Computing, 22(2):193–213, March 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0120-0.

Fiadeiro:2010:E

[625] J. L. Fiadeiro. Editorial. Formal Aspects of Computing, 22(3– 4):215–216, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-010-0151-
6;   http://link.springer.com/article/
10.1007/s00165-010-0151-6.

Alrajeh:2010:DNZ

[626] D. Alrajeh, J. Kramer, A. Russo, and S. Uchitel. Deriving non-Zeno behaviour models from goal models using ILP. Formal Aspects of Computing, 22(3–4):217–241, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0128-5.

Bocchi:2010:SPS

[627] Laura Bocchi, Stephen Gorton, and Stephan Reiff-Marganiec. From StPowla processes to SRML models. Formal Aspects of Computing, 22 (3–4):243–268, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0118-7.

Boronat:2010:ASM

[628] Artur Boronat and José Meseguer. An algebraic semantics for MOF.

Formal Aspects of Computing, 22 (3–4):269–296, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0140-9.

deLara:2010:ATB

[629] Juan de Lara and Hans Vangheluwe. Automating the transformation-based analysis of visual languages. Formal Aspects of Computing, 22(3– 4):297–326, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0114-y.

Ehrig:2010:CIM

[630] Hartmut Ehrig, Karsten Ehrig, Claudia Ermel, and Ulrike Prange. Consistent integration of models based on views of meta models. Formal Aspects of Computing, 22(3–4):327–344, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0127-6.

Moha:2010:DAS

[631] Naouel Moha, Yann-Gaël Guéhéneuc, Anne-Françoise Le Meur, Laurence Duchien, and Alban Tiberghien. From a domain analysis to the specification and detection of code and design smells. Formal Aspects of Computing, 22(3–4):345–361, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0115-x.

Mossakowski:2010:GCD

[632] Till Mossakowski, Lutz Schröder, and Sergey Goncharov. A generic complete dynamic logic for reasoning about purity and effects. Formal Aspects of Computing, 22(3–4):363–384, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-010-
0153-4.

Orejas:2010:RGC

[633] Fernando Orejas, Hartmut Ehrig, and Ulrike Prange. Reasoning with graph constraints. Formal Aspects of Computing, 22(3–4):385–422, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0116-9.

Smans:2010:AVJ

[634] Jan Smans, Bart Jacobs, Frank Piessens, and Wolfram Schulte. Automatic verification of Java programs with dynamic frames. Formal Aspects of Computing, 22(3–4):423–457, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-010-
0148-1.

vanderAalst:2010:PCD

[635] Wil M. P. van der Aalst, Marlon Dumas, Florian Gottschalk, Arthur H. M. ter Hofstede, Marcello La Rosa, and Jan Mendling. Preserving correctness during business process model configuration. Formal Aspects of Computing, 22(3–4):459–482, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-009-0112-0;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-009-0112-0.pdf.

Collinson:2010:EAL

[636] Matthew Collinson and David Pym. Erratum to: Algebra and logic for access control. Formal Aspects of Computing, 22(3–4):483–484, May 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-010-
                                        0155-2;   http://link.springer.
                                        com/content/pdf/10.1007/s00165-
                                        010-0155-2.pdf. See [620].

Anonymous:2010:RMJ

[637] Anonymous. Robin Milner: 13 January 1934—20 March 2010. Formal Aspects of Computing, 22(5):485–487, September 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-010-
                                        0162-3.

Duran:2010:AAD

[638] Adolfo Duran, Ana Cavalcanti, and Augusto Sampaio. An algebraic approach to the design of compilers for object-oriented languages. Formal Aspects of Computing, 22(5): 489–535, September 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-009-0124-9.

Hesselink:2010:SEL

[639] Wim H. Hesselink. Solutions of equations in languages. Formal Aspects of Computing, 22(5):537–545, September 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0123-x;   http://link.springer.
com/content/pdf/10.1007/s00165-
009-0123-x.pdf.

Filipovic:2010:BCD

[640] Ivana Filipović, Peter O’Hearn, Noah Torp-Smith, and Hongseok Yang. Blaming the client: on data refinement in the presence of pointers. Formal Aspects of Computing, 22(5): 547–583, September 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0125-8.

Saleh:2010:GTF

[641] Mohamed Saleh and Mourad Debbabi. A game-theoretic framework for specification and verification of cryptographic protocols. Formal Aspects of Computing, 22(5):585–609, September 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0129-4.

BohorquezV:2010:EUA

[642] Jaime A. Bohórquez V. An elementary and unified approach to program correctness. Formal Aspects of Computing, 22(5):611–627, September 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-009-
                                        0137-4.

Back:2010:SDU

[643] Ralph-Johan Back. Structured derivations: a unified proof style for teaching mathematics. Formal Aspects of Computing, 22(5):629–661, September 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-009-
                                        0136-5.

Harel:2010:APG

[644] David Harel. Amir Pnueli: A gentle giant: Lord of the ϕ’s and the ψ’s. Formal Aspects of Computing, 22(6): 663–665, November 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-010-0165-0.

Jourdan:2010:LBL

[645] Guy-Vincent Jourdan, Hasan Ural, Hüsnü Yenigün, and Ji Chao Zhang. Lower bounds on lengths of checking sequences. Formal Aspects of Computing, 22(6):667–679, November 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-009-
                                        0135-6.

Lang:2010:TFL

[646] Frédéric Lang, Gwen Salaün, Rémi Hérilier, Jeff Kramer, and Jeff Magee. Translating FSP into LOTOS and networks of automata. Formal Aspects of Computing, 22(6):681–711, November 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0133-8.

Wahl:2010:LAS

[647] Thomas Wahl and Vijay D’Silva. A lazy approach to symmetry reduction. Formal Aspects of Computing, 22(6): 713–733, November 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0131-x.

Bornat:2010:IPB

[648] Richard Bornat and Hasan Amjad. Inter-process buffers in separation logic with rely-guarantee. Formal Aspects of Computing, 22(6):735–772, November 2010. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0141-8.

Borger:2011:E

[649] Egon Börger.  Editorial.  For-

mal Aspects of Computing, 23(1): 1–2, January 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-010-0168-
x;   http://link.springer.com/article/
10.1007/s00165-010-0168-x.

Julliand:2011:GTB

[650] Jacques Julliand, Pierre-Alain Masson, Régis Tissot, and Pierre-Christophe Bué. Generating tests from B specifications and dynamic selection criteria. Formal Aspects of Computing, 23(1):3–19, January 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-009-0132-9.

Cavarra:2011:DFA

[651] Alessandra Cavarra. A data-flow approach to test multi-agent ASMs. Formal Aspects of Computing, 23 (1):21–41, January 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-009-0134-7.

Derrick:2011:ZTB

[652] John Derrick, Siobhán North, and Anthony J. H. Simons. Z2SAL: a translation-based model checker for Z. Formal Aspects of Computing, 23 (1):43–71, January 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-009-0126-7.

Wright:2011:UEB

[653] Stephen Wright and Kerstin Eder. Using Event-B to construct instruction set architectures. Formal Aspects of Computing, 23(1):73–89, January 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-009-
                                        0142-7.

Baumler:2011:PLT

[654] Simon Bäumler, Gerhard Schellhorn, Bogdan Tofan, and Wolfgang Reif. Proving linearizability with temporal logic. Formal Aspects of Computing, 23(1):91–112, January 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0130-y.

Banach:2011:REB

[655] Richard Banach. Retrenchment for Event-B: UseCase-wise development and Rodin integration. Formal Aspects of Computing, 23(1):113–131, January 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0139-2.

Hallerstede:2011:PEB

[656] Stefan Hallerstede. On the purpose of Event-B proof obligations. Formal Aspects of Computing, 23(1): 133–150, January 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0138-3.

Liu:2011:E

[657] Zhiming Liu and Jim Woodcock. Editorial. Formal Aspects of Computing, 23(2):151, March 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0173-8;   http:
//link.springer.com/content/pdf/
10.1007/s00165-011-0173-8.pdf.

Butterfield:2011:DSH

[658] Andrew Butterfield. A denotational semantics for Handel-C. Formal Aspects of Computing, 23(2):153–170, March 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0146-3.

Xia:2011:SDP

[659] Bican Xia, Lu Yang, Naijun Zhan, and Zhihai Zhang. Symbolic decision procedure for termination of linear programs. Formal Aspects of Computing, 23(2):171–190, March 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-009-0144-5.

Haxthausen:2011:FAC

[660] Anne E. Haxthausen, Jan Peleska, and Sebastian Kinder. A formal approach for the construction and verification of railway control systems. Formal Aspects of Computing, 23(2):191–219, March 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0143-6.

Mosses:2011:VSP

[661] Peter D. Mosses. VDM semantics of programming languages: combinators and monads. Formal Aspects of Computing, 23(2):221–238, March 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-009-
0145-4.

vanLamsweerde:2011:HHR

[662] Axel van Lamsweerde. The humble humorous researcher: A tribute to Michel Sintzoff. Formal Aspects of Computing, 23(3):239–242, May 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0176-5.

Brown:2011:AVR

[663] Geoffrey M. Brown and Lee Pike. Automated verification and refinement for physical-layer protocols. Formal Aspects of Computing, 23(3):243–266, May 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-010-
0149-0.

Kassios:2011:DFT

[664] I. T. Kassios. The dynamic frames theory. Formal Aspects of Computing, 23(3):267–288, May 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-010-0152-5.

Jones:2011:ECA

[665] Cliff B. Jones and Ken G. Pierce. Elucidating concurrent algorithms via layers of abstraction and reification. Formal Aspects of Computing, 23(3):289–306, May 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-010-
0156-1.

Sinnig:2011:POS

[666] Daniel Sinnig, Ferhat Khendek, and Patrice Chalin. Partial order semantics for use case and task models. Formal Aspects of Computing, 23(3):307–332, May 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00165-010-
                                        0158-z.

vanderAalst:2011:SWN

[667] W. M. P. van der Aalst, K. M. van Hee, A. H. M. ter Hofstede, N. Sidorova, H. M. W. Verbeek, M. Voorhoeve, and M. T. Wynn. Soundness of workflow nets: classification, decidability, and analysis. Formal Aspects of Computing, 23(3):333–363, May 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-010-0161-4;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-010-0161-4.pdf.

Holzmann:2011:MCB

[668] Gerard J. Holzmann and Mihai Florian. Model checking with bounded context switching. Formal Aspects of Computing, 23(3):365–389, May 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-010-
                                        0160-5.

Hehner:2011:PP

[669] Eric C. R. Hehner. A probability perspective. Formal Aspects of Computing, 23(4):391–419, July 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-010-0157-0.

Stewart:2011:PMB

[670] Alan Stewart. A programming model for BSP with partitioned synchronisation. Formal Aspects of Computing, 23(4):421–432, July 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-010-0163-2;   http:
//link.springer.com/content/pdf/
10.1007/s00165-010-0163-2.pdf.

Fiadeiro:2011:AMS

[671] José Luiz Fiadeiro, Antónia Lopes, and Laura Bocchi. An abstract model of service discovery and binding. Formal Aspects of Computing, 23(4):433–463, July 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-010-
0166-z.

Cavalcanti:2011:CLD

[672] Ana Cavalcanti, Phil Clayton, and Colin O’Halloran. From control law diagrams to Ada via Circus. Formal Aspects of Computing, 23(4):465–512, July 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-010-
0170-3.

Yang:2011:IGB

[673] Linmin Yang, Zhe Dang, and Thomas R. Fischer. Information gain of black-box testing. Formal Aspects of Computing, 23(4):513–539, July 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0175-6.

Gao:2011:CPS

[674] Han Gao, Flemming Nielson, and Hanne Riis Nielson. CaPiTo: protocol stacks for services. Formal Aspects of Computing, 23(4):541–565, July 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-011-
                                        0174-7.

Bergstra:2011:TAP

[675] J. A. Bergstra and C. A. Middelburg. Thread algebra for polythreading. Formal Aspects of Computing, 23(4):567–583, July 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0178-3;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-011-0178-3.pdf.

Kroening:2011:E

[676] Daniel Kroening, Tiziana Margaria, and Jim Woodcock. Editorial. Formal Aspects of Computing, 23(5): 585–588, September 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0201-8;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-011-0201-8.pdf.

Jorges:2011:APC

[677] Sven Jörges, Tiziana Margaria, and Bernhard Steffen. Assuring property conformance of code generators via model checking. Formal Aspects of Computing, 23(5):589–606, September 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-010-
                                        0169-9.

Sitaraman:2011:BPB

[678] Murali Sitaraman, Bruce Adcock, Jeremy Avigad, Derek Bronish, Paolo Bucci, David Frazier, Harvey M. Friedman, Heather Harton, Wayne Heym, Jason Kirschenbaum, Joan Krone, Hampton Smith, and Bruce W. Weide. Building a push-button RESOLVE verifier: Progress and challenges. Formal Aspects of Computing, 23(5): 607–626, September 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-010-0154-3.

Braghin:2011:MCB

[679] Chiara Braghin, Natasha Sharygina, and Katerina Barone-Adesi. A model checking-based approach for security policy verification of mobile systems. Formal Aspects of Computing, 23(5): 627–648, September 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-010-0159-y.

Chandy:2011:VDS

[680] K. Mani Chandy, Brian Go, Sayan Mitra, Concetta Pilotto, and Jerome White. Verification of distributed systems with local—global predicates. Formal Aspects of Computing, 23(5): 649–679, September 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-010-0150-7.

Cavalcanti:2011:E

[681] Ana Cavalcanti, Dennis Dams, and Marie-Claude Gaudel. Editorial. Formal Aspects of Computing, 23 (6):681, November 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0211-6;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-011-0211-6.pdf.

Leuschel:2011:APV

[682] Michael Leuschel, Jérôme Falampin, Fabian Fritz, and Daniel Plagge. Automated property verification for large scale B models with ProB. Formal Aspects of Computing, 23(6): 683–709, November 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-010-0172-1;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-010-0172-1.pdf.

McIver:2011:CRA

[683] A. K. McIver and C. C. Morgan. Compositional refinement in agent-based security protocols. Formal Aspects of Computing, 23(6):711–737, November 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-010-
                                        0164-1.

Reynolds:2011:TBD

[684] Mark Reynolds. A tableau-based decision procedure for CTL*. Formal Aspects of Computing, 23(6): 739–779, November 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0193-4.

Wang:2011:SPA

[685] Chao Wang, Sudipta Kundu, Rhishikesh Limaye, Malay Ganai, and Aarti Gupta. Symbolic predictive analysis for concurrent programs. Formal Aspects of Computing, 23(6):781–805, November 2011. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0179-2.

Boiten:2012:E

[686] Eerke Boiten, John Derrick, Jin Song Dong, and Steve Reeves. Editorial. Formal Aspects of Computing, 24(1):1, January 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0220-5;   http:
//link.springer.com/content/pdf/
10.1007/s00165-011-0220-5.pdf.

Morgan:2012:CNF

[687] Carroll Morgan. Compositional noninterference from first principles. Formal Aspects of Computing, 24(1):3–26, January 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-010-
0167-y.

Hesselink:2012:FHF

[688] Wim H. Hesselink and Muhammad Ikram Lali. Formalizing a hierarchical file system. Formal Aspects of Computing, 24(1):27–44, January 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-010-
0171-2;   http://link.springer.
com/content/pdf/10.1007/s00165-
010-0171-2.pdf.

Sanders:2012:ER

[689] J. W. Sanders and Graeme Smith. Emergence and refinement. Formal Aspects of Computing, 24(1):45–65, January 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-011-
                                        0190-7.

Preoteasa:2012:IDD

[690] Viorel Preoteasa and Ralph-Johan Back. Invariant diagrams with data refinement. Formal Aspects of Computing, 24(1):67–95, January 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00165-011-0195-
                                        2.

Hallerstede:2012:EPV

[691] Stefan Hallerstede and Michael Leuschel. Experiments in program verification using Event-B. Formal Aspects of Computing, 24(1):97–125, January 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-011-
                                        0205-4.

Zeyda:2012:MSS

[692] Frank Zeyda, Marcel Oliveira, and Ana Cavalcanti. Mechanised support for sound refinement tactics. Formal Aspects of Computing, 24(1): 127–160, January 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0218-z.

Fiadeiro:2012:E

[693] J. L. Fiadeiro, S. Gnesi, and T. Maibaum. Editorial. Formal Aspects of Computing, 24(2):161–162, March 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0224-9;   http://link.springer.
com/content/pdf/10.1007/s00165-
012-0224-9.pdf.

Filieri:2012:FAA

[694] Antonio Filieri, Carlo Ghezzi, and Giordano Tamburrelli. A formal approach to adaptive software: continuous assurance of non-functional requirements. Formal Aspects of Computing, 24(2):163–186, March 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s00165-011-0207-
2.

Kokash:2012:RMF

[695] Natallia Kokash, Christian Krause, and Erik de Vink. Reo + mCRL2: A framework for model-checking dataflow in service compositions. Formal Aspects of Computing, 24(2):187–216, March 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0191-6.

Barros:2012:ABS

[696] José Bernardo Barros, Daniela da Cruz, Pedro Rangel Henriques, and Jorge Sousa Pinto. Assertion-based slicing and slice graphs. Formal Aspects of Computing, 24(2):217–248, March 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0196-1.

Lindsay:2012:CSA

[697] Peter A. Lindsay, Nisansala Yatapanage, and Kirsten Winter. Cut set analysis using behavior trees and model checking. Formal Aspects of Computing, 24(2):249–266, March 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-011-
                                        0181-8.

Massink:2012:SCD

[698] M. Massink, D. Latella, A. Bracciali, M. D. Harrison, and J. Hillston. Scalable context-dependent analysis of emergency egress models. Formal Aspects of Computing, 24(2):267–302, March 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-011-
                                        0188-1.

Woodcock:2012:E

[699] Jim Woodcock. Editorial. Formal Aspects of Computing, 24(3):303, May 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-012-
                                        0229-4;   http://link.springer.
                                        com/content/pdf/10.1007/s00165-
                                        012-0229-4.pdf.

Jones:2012:JM

[700] Cliff B. Jones. John McCarthy (1927— 2011). Formal Aspects of Computing, 24(3):305–306, May 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-012-0226-
7;   http://link.springer.com/article/
10.1007/s00165-012-0226-7.

Risi:2012:UFF

[701] Michele Risi, Giuseppe Scanniello, and Genoveffa Tortora. Using foldin and fold-out in the architecture recovery of software systems. Formal Aspects of Computing, 24(3):307–330, May 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0199-y.

Andrade:2012:TIR

[702] Wilkerson L. Andrade and Patrícia D. L. Machado. Testing interruptions in reactive systems. Formal Aspects of Computing, 24(3):331–353, May 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0197-0.

Kim:2012:CTM

[703] Moonzoo Kim, Yunho Kim, and Yunja Choi. Concolic testing of the multisector read operation for flash storage platform software. Formal Aspects of Computing, 24(3):355–374, May 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0200-9.

Muhlberg:2012:VCF

[704] Jan Tobias Mühlberg and Gerald Lüttgen. Verifying compiled file system code. Formal Aspects of Computing, 24(3):375–391, May 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0198-z.

Derrick:2012:TLP

[705] John Derrick and Graeme Smith. Temporal-logic property preservation under Z refinement. Formal Aspects of Computing, 24(3):393–416, May 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-011-
                                        0177-4.

Hofner:2012:P

[706] P. Höfner. Preface. Formal Aspects of Computing, 24(4–6):417–422, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-012-
                                        0257-0;   http://link.springer.
                                        com/content/pdf/10.1007/s00165-
                                        012-0257-0.pdf.

Hoare:2012:PA

[707] Tony Hoare and Stephan van Staden. In praise of algebra. Formal Aspects of Computing, 24(4–6):423–431, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-012-
                                        0249-0.

Oliveira:2012:TLA

[708] José N. Oliveira. Towards a linear algebra of programming. Formal Aspects of Computing, 24(4–6):433–458, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0240-9.

Hofner:2012:DFW

[709] Peter Höfner and Bernhard Möller. Dijkstra, Floyd and Warshall meet Kleene. Formal Aspects of Computing, 24(4–6):459–476, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0245-4.

Swaminathan:2012:LRR

[710] Mani Swaminathan, Joost-Pieter Katoen, and Ernst-Rüdiger Olderog. Layered reasoning for randomized distributed algorithms. Formal Aspects of Computing, 24(4–6):477–496, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0231-x.

Markovski:2012:RRS

[711] J. Markovski, P. R. D’Argenio, J. C. M. Baeten, and E. P. de Vink. Reconciling real and stochastic time: the need for probabilistic refinement. Formal Aspects of Computing, 24(4–6):497–518, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0230-y;   http://link.springer.

                                        com/content/pdf/10.1007/s00165-
                                        012-0230-y.pdf.

Leino:2012:SRH

[712] K. Rustan M. Leino and Kuat Yessenov. Stepwise refinement of heap-manipulating code in Chalice. Formal Aspects of Computing, 24 (4–6):519–535, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0254-3.

Jonsson:2012:URC

[713] Bengt Jonsson. Using refinement calculus techniques to prove linearizability. Formal Aspects of Computing, 24 (4–6):537–554, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0250-7.

Butler:2012:EIC

[714] Michael Butler. External and internal choice with event groups in Event-B. Formal Aspects of Computing, 24 (4–6):555–567, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0239-2.

Pettorossi:2012:CBC

[715] Alberto Pettorossi, Maurizio Proietti, and Valerio Senni. Constraintbased correctness proofs for logic program transformations. Formal Aspects of Computing, 24(4–6):569–594, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00165-012-
0233-8.

Bouyer:2012:TIF

[716] Patricia Bouyer, Nicolas Markey, Joël Ouaknine, Philippe Schnoebelen, and James Worrell. On termination and invariance for faulty channel machines. Formal Aspects of Computing, 24 (4–6):595–607, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0234-7.

Bird:2012:BCS

[717] Richard S. Bird. On building cyclic and shared structures in Haskell. Formal Aspects of Computing, 24 (4–6):609–621, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0243-6.

Emoto:2012:FES

[718] Kento Emoto, Sebastian Fischer, and Zhenjiang Hu. Filter-embedding semiring fusion for programming with MapReduce. Formal Aspects of Computing, 24(4–6):623–645, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s00165-012-0241-
8.

Winskel:2012:DCS

[719] Glynn Winskel. Deterministic concurrent strategies. Formal Aspects of Computing, 24(4–6):647–660, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00165-012-
                                        0235-6.

Kwiatkowska:2012:PVH

[720] Marta Kwiatkowska, Gethin Norman, and David Parker. Probabilistic verification of Herman’s self-stabilisation algorithm. Formal Aspects of Computing, 24(4–6):661–670, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0227-6.

Kiefer:2012:TTH

[721] Stefan Kiefer, Andrzej S. Murawski, Joël Ouaknine, Björn Wachter, and James Worrell. Three tokens in Herman’s algorithm. Formal Aspects of Computing, 24(4–6):671–678, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-012-
                                        0228-5.

Hierons:2012:UST

[722] Robert M. Hierons and Manuel Núñez. Using schedulers to test probabilistic distributed systems. Formal Aspects of Computing, 24(4–6):679–699, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-012-
                                        0244-5.

Deng:2012:CTP

[723] Yuxin Deng and Alwen Tiu. Characterisations of testing preorders for a finite probabilistic π-calculus. Formal Aspects of Computing, 24(4–6):701–726, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0238-3.

Georgievska:2012:PMM

[724] Sonja Georgievska and Suzana Andova. Probabilistic may/must testing: retaining probabilities by restricted schedulers. Formal Aspects of Computing, 24(4–6):727–748, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0236-5;   http://link.springer.
com/content/pdf/10.1007/s00165-
012-0236-5.pdf.

Hennessy:2012:EPB

[725] Matthew Hennessy. Exploring probabilistic bisimulations, part I. Formal Aspects of Computing, 24(4– 6):749–768, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0242-7.

vanderMeyden:2012:ARN

[726] Ron van der Meyden. Architectural refinement and notions of intransitive noninterference. Formal Aspects of Computing, 24(4–6):769–792, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0247-2.

Misra:2012:SVS

[727] Jayadev Misra. A secure voting scheme based on rational self-interest. Formal Aspects of Computing, 24 (4–6):793–805, July 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0248-1.

Anonymous:2013:P

[728] Anonymous. Preface. Formal Aspects of Computing, 25(1):1–2, January 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00165-012-0271-2;   http://link.
                                        springer.com/article/10.1007/s00165-
                                        012-0271-2.

Roscoe:2013:CNT

[729] A. W. Roscoe and Jian Huang. Checking noninterference in Timed CSP. Formal Aspects of Computing, 25 (1):3–35, January 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0251-6.

Cavalcanti:2013:SCJ

[730] Ana Cavalcanti, Andy Wellings, and Jim Woodcock. The Safety-Critical Java memory model formalised. Formal Aspects of Computing, 25(1): 37–57, January 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0253-4.

Hoang:2013:SID

[731] Thai Son Hoang. Security invariants in discrete transition systems. Formal Aspects of Computing, 25(1):59–87, January 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0256-1.

Chen:2013:SIU

[732] Yifeng Chen. Semantic inheritance in unifying theories of programming. Formal Aspects of Computing, 25 (1):89–106, January 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-007-0051-6.

Stoddart:2013:UPC

[733] Bill Stoddart and Frank Zeyda. A unification of probabilistic choice within a design-based model of reversible computation. Formal Aspects of Computing, 25(1):107–131, January 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0048-1.

Oliveira:2013:UTP

[734] Marcel Oliveira, Ana Cavalcanti, and Jim Woodcock. Unifying theories in ProofPower-Z. Formal Aspects of Computing, 25(1):133–158, January 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-007-
0044-5.

Runde:2013:RCS

[735] Ragnhild Kobro Runde, Atle Refsdal, and Ketil Stølen. Relating computer systems to sequence diagrams: the impact of underspecification and inherent nondeterminism. Formal Aspects of Computing, 25(2):159–187, March 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-011-
                                        0192-5.

Johnson:2013:DTS

[736] Kenneth Johnson and John V. Tucker. The data type of spatial objects. Formal Aspects of Computing, 25 (2):189–218, March 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0182-7.

Murray:2013:LRT

[737] Toby Murray. On the limits of refinement-testing for model-checking CSP. Formal Aspects of Computing, 25(2):219–256, March 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0183-6.

Damgaard:2013:ICM

[738] Troels C. Damgaard, Arne J. Glenstrup, Lars Birkedal, and Robin Milner. An inductive characterization of matching in binding bigraphs. Formal Aspects of Computing, 25(2):257–288, March 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-011-
                                        0184-5.

Gomez:2013:MCT

[739] Rodolfo Gómez. Model-checking timed automata with deadlines with Uppaal. Formal Aspects of Computing, 25(2):289–318, March 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0185-4.

Subramani:2013:IAO

[740] K. Subramani, Matthew Williamson, and Xiaofeng Gu. Improved algorithms for optimal length resolution refutation in difference constraint systems. Formal Aspects of Computing, 25(2):319–341, March 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0186-3.

Bowen:2013:E

[741] Jonathan P. Bowen, Michael Butler, Steve Reeves, and Mike Hinchey. Editorial. Formal Aspects of Computing, 25(3):343, May 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0281-8;   http:
//link.springer.com/content/pdf/
10.1007/s00165-013-0281-8.pdf.

Arthan:2013:HLL

[742] Rob Arthan, Ursula Martin, and Paulo Oliva. A Hoare logic for linear systems. Formal Aspects of Computing, 25(3):345–363, May 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0180-9.

Perna:2013:MCR

[743] Juan I. Perna and Chris George. Model checking RAISE applicative specifications. Formal Aspects of Computing, 25(3):365–388, May 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0217-0.

Babic:2013:PTN

[744] Domagoj Babić, Byron Cook, Alan J. Hu, and Zvonimir Rakamarić. Proving termination of nonlinear command sequences. Formal Aspects of Computing, 25(3):389–403, May 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0252-5.

Beckert:2013:DLD

[745] Bernhard Beckert and Vladimir Klebanov. A Dynamic Logic for deductive verification of multi-threaded programs. Formal Aspects of Computing, 25(3):405–437, May 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0261-4.

Banach:2013:AFR

[746] Richard Banach, Czeslaw Jeske, Anthony Hall, and Susan Stepney. Atomicity failure and the retrenchment atomicity pattern. Formal Aspects of Computing, 25(3):439–464, May 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0216-1.

Li:2013:IAS

[747] Yongjian Li and Jun Pang. An inductive approach to strand spaces. Formal Aspects of Computing, 25(4):465–501, July 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0187-2.

Galpin:2013:HHM

[748] Vashti Galpin, Luca Bortolussi, and Jane Hillston. HYPE: Hybrid modelling by composition of flows. Formal Aspects of Computing, 25(4):503–541, July 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0189-0.

Poch:2013:TBP

[749] Tomás Poch, Ondrej Serý, Frantisek Plásil, and Jan Kofron. Threaded behavior protocols. Formal Aspects of Computing, 25(4):543–572, July 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0194-3.

Banach:2013:MGFa

[750] Richard Banach and Marco Bozzano. The mechanical generation of fault trees for reactive systems via retrenchment I: combinational circuits. Formal Aspects of Computing, 25(4):573–607, July 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0202-7.

Banach:2013:MGFb

[751] Richard Banach and Marco Bozzano. The mechanical generation of fault trees for reactive systems via retrenchment II: clocked and feedback circuits. Formal Aspects of Computing, 25(4):609–657, July 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0203-6.

Eshuis:2013:SPN

[752] Rik Eshuis. Statechartable Petri nets. Formal Aspects of Computing, 25(5): 659–681, September 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0204-5;   http:
//link.springer.com/content/pdf/
10.1007/s00165-011-0204-5.pdf.

Brucker:2013:TPB

[753] Achim D. Brucker and Burkhart Wolff. On theorem prover-based testing. Formal Aspects of Computing, 25(5): 683–721, September 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0222-y.

John:2013:CVC

[754] Mathias John, Hans-Jörg Schulz, Heidrun Schumann, Adelinde M. Uhrmacher, and Andrea Unger. Constructing and visualizing chemical reaction networks from pi-calculus models. Formal Aspects of Computing, 25(5): 723–742, September 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-011-0209-0.

Rabanal:2013:TRS

[755] Pablo Rabanal, Ismael Rodríguez, and Fernando Rubio. Testing restorable systems: formal definition and heuristic solution based on river formation dynamics. Formal Aspects of Computing, 25(5):743–768, September 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0206-3.

Doherty:2013:TFS

[756] Simon Doherty, Lindsay Groves, Victor Luchangco, and Mark Moir. Towards formally specifying and verifying transactional memory. Formal Aspects of Computing, 25(5):769–799, September 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0225-8.

Merro:2013:CTA

[757] Massimo Merro and Eleonora Sibilio. A calculus of trustworthy ad hoc networks. Formal Aspects of Computing, 25(5):801–832, September 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s00165-011-0210-
7.

Stewart:2013:RAO

[758] Alan Stewart, Joaquim Gabarro, and Anthony Keenan. Reasoning about orchestrations of web services using partial correctness. Formal Aspects of Computing, 25(6):833–846, November 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-011-
                                        0212-5;   http://link.springer.
                                        com/content/pdf/10.1007/s00165-
                                        011-0212-5.pdf.

Fu:2013:SLS

[759] Xiang Fu, Michael C. Powell, Michael Bantegui, and Chung-Chih Li. Simple linear string constraints. Formal Aspects of Computing, 25(6): 847–891, November 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0214-3.

Bornat:2013:ETN

[760] Richard Bornat and Hasan Amjad. Explanation of two non-blocking sharedvariable communication algorithms. Formal Aspects of Computing, 25(6): 893–931, November 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0213-4.

Bientinesi:2013:DDL

[761] Paolo Bientinesi, John A. Gunnels, Margaret E. Myers, Enrique S. Quintana-Ortí, Tyler Rhodes, Robert A. van de Geijn, and Field G. Van Zee. Deriving dense linear algebra libraries. Formal Aspects of Computing, 25(6): 933–945, November 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-011-0221-4.

Hesselink:2013:SFM

[762] Wim H. Hesselink and Mark IJbema. Starvation-free mutual exclusion with semaphores. Formal Aspects of Computing, 25(6):947–969, November 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-011-
0219-y;   http://link.springer.
com/content/pdf/10.1007/s00165-
011-0219-y.pdf.

Abed:2013:AVR

[763] Sa’ed Abed, Otmane Ait Mohamed, and Ghiath Al Sammane. Automatic verification of reduction techniques in higher order logic. Formal Aspects of Computing, 25(6):971–991, November 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0223-x.

Gottliebsen:2013:ARN

[764] Hanne Gottliebsen, Ruth Hardy, Olga Lightfoot, and Ursula Martin. Applications of real number theorem proving in PVS. Formal Aspects of Computing, 25 (6):993–1016, November 2013. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0232-9;   http:
//link.springer.com/content/pdf/
10.1007/s00165-012-0232-9.pdf.

Boiten:2014:Ea

[765] Eerke Boiten and Steve Schneider. Editorial. Formal Aspects of Computing, 26(1):1–2, January 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-013-0290-7;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-013-0290-7.pdf.

Bella:2014:ISC

[766] Giampaolo Bella. Inductive study of confidentiality: for everyone. Formal Aspects of Computing, 26(1):3–36, January 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-012-
                                        0246-3.

Heather:2014:CPE

[767] James Heather, Steve Schneider, and Vanessa Teague. Cryptographic protocols with everyday objects. Formal Aspects of Computing, 26(1):37–62, January 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-013-
                                        0274-7.

Moran:2014:VAV

[768] Murat Moran, James Heather, and Steve Schneider. Verifying anonymity in voting systems using CSP. Formal Aspects of Computing, 26(1): 63–98, January 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0268-x.

Avalle:2014:FVS

[769] Matteo Avalle, Alfredo Pironti, and Riccardo Sisto. Formal verification of security protocol implementations: a survey. Formal Aspects of Computing, 26(1):99–123, January 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0269-9.

Pironti:2014:SAD

[770] Alfredo Pironti and Riccardo Sisto. Safe abstractions of data encodings in formal security protocol models. Formal Aspects of Computing, 26(1): 125–167, January 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0267-y.

Hoang:2014:ANI

[771] T. S. Hoang, A. K. McIver, L. Meinicke, C. C. Morgan, A. Sloane, and E. Susatyo. Abstractions of noninterference security: probabilistic versus possibilistic. Formal Aspects of Computing, 26(1):169–194, January 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0237-4.

Boiten:2014:Eb

[772] Eerke A. Boiten, John Derrick, and Steve Reeves. Editorial. Formal Aspects of Computing, 26(2):195, March 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0293-z;   http://link.springer.
com/content/pdf/10.1007/s00165-
014-0293-z.pdf.

Petre:2014:KSM

[773] Luigia Petre, Elena Troubitsyna, and Marina Waldén. Kaisa Sere: In memoriam. Formal Aspects of Computing, 26(2):197–201, March 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0292-5;   http:
//link.springer.com/content/pdf/
10.1007/s00165-013-0292-5.pdf.

Llano:2014:DIT

[774] Maria Teresa Llano, Andrew Ireland, and Alison Pease. Discovery of invariants through automated theory formation. Formal Aspects of Computing, 26(2):203–249, March 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0264-1.

Schneider:2014:BSE

[775] Steve Schneider, Helen Treharne, and Heike Wehrheim. The behavioural semantics of Event-B refinement. Formal Aspects of Computing, 26(2):251–280, March 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0265-0.

Bostrom:2014:DCP

[776] Pontus Boström, Fredrik Degerlund, Kaisa Sere, and Marina Waldén. Derivation of concurrent programs by stepwise scheduling of Event-B models. Formal Aspects of Computing, 26(2):281–303, March 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0260-5.

Boiten:2014:IEO

[777] Eerke A. Boiten. Introducing extra operations in refinement. Formal Aspects of Computing, 26(2):305–317, March 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0266-z.

Banach:2014:CKA

[778] Richard Banach, Huibiao Zhu, Wen Su, and Runlei Huang. Continuous KAOS, ASM, and formal control system design across the continuous/ discrete modeling interface: a simple train stopping application. Formal Aspects of Computing, 26(2):319–366, March 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0263-2.

Miyazawa:2014:RBV

[779] Alvaro Miyazawa and Ana Cavalcanti. Refinement-based verification of implementations of Stateflow charts. Formal Aspects of Computing, 26(2):367–405, March 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-013-
0291-6.

Derrick:2014:RCR

[780] John Derrick and Eerke Boiten. Relational concurrent refinement part III: traces, partial relations and automata. Formal Aspects of Computing, 26(2):407–432, March 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0262-3.

Jones:2014:Ea

[781] C. B. Jones. Editorial. Formal Aspects of Computing, 26(3):433, May 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-014-
                                        0298-7;   http://link.springer.
                                        com/content/pdf/10.1007/s00165-
                                        014-0298-7.pdf.

Brookes:2014:ER

[782] Stephen Brookes, Peter W. O’Hearn, and Uday Reddy. The essence of Reynolds. Formal Aspects of Computing, 26(3):435–439, May 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-014-0295-x.

Nogueira:2014:TGS

[783] Sidney Nogueira, Augusto Sampaio, and Alexandre Mota. Test generation from state based use case models. Formal Aspects of Computing, 26(3):441–490, May 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-012-0258-z.

Colvin:2014:OSO

[784] Robert J. Colvin. An operational semantics for object-oriented concepts based on the class hierarchy. Formal Aspects of Computing, 26(3):491–535, May 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00165-012-
0259-y.

Calder:2014:MIC

[785] Muffy Calder and Michele Sevegnani. Modelling IEEE 802.11 CSMA/ CA RTS/CTS with stochastic bigraphs with sharing. Formal Aspects of Computing, 26(3):537–561, May 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-012-
0270-3;   http://link.springer.
com/content/pdf/10.1007/s00165-
012-0270-3.pdf.

Dongol:2014:RAG

[786] Brijesh Dongol, Ian J. Hayes, and Peter J. Robinson. Reasoning about goaldirected real-time teleo-reactive programs. Formal Aspects of Computing, 26(3):563–589, May 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0272-1.

Dixit:2014:TBC

[787] Manoj G. Dixit, S. Ramesh, and Pallab Dasgupta. Time-budgeting: a component based development methodology for real-time embedded systems. Formal Aspects of Computing, 26(3):591–621, May 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-012-0273-0.

Liu:2014:CD

[788] Xi Liu, Shaofa Yang, and J. W. Sanders. Compensation by design.

Formal Aspects of Computing, 26 (4):623–676, July 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0275-6.

Konur:2014:FVP

[789] Savas Konur, Michael Fisher, Simon Dobson, and Stephen Knox. Formal verification of a pervasive messaging system. Formal Aspects of Computing, 26(4):677–694, July 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0277-4.

Sampath:2014:ETP

[790] Prahladavaradan Sampath. An elementary theory of product-line variations. Formal Aspects of Computing, 26(4):695–727, July 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0276-5.

Dong:2014:TVC

[791] Jin Song Dong, Yang Liu, Jun Sun, and Xian Zhang. Towards verification of computation orchestration. Formal Aspects of Computing, 26(4):729–759, July 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-013-
0280-9.

Damiani:2014:VTI

[792] Ferruccio Damiani, Johan Dovland, Einar Broch Johnsen, and Ina Schaefer. Verifying traits: an incremental proof system for fine-grained reuse. Formal Aspects of Computing, 26(4):761–793, July 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-013-
0278-3.

Cavalcanti:2014:TDG

[793] Ana Cavalcanti, Steve King, Colin O’Halloran, and Jim Woodcock. Testdata generation for control coverage by proof. Formal Aspects of Computing, 26(4):795–823, July 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0279-2.

Chatterjee:2014:AGS

[794] Krishnendu Chatterjee and Vishwanath Raman. Assume-guarantee synthesis for digital contract signing. Formal Aspects of Computing, 26(4):825–859, July 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0283-6.

Jones:2014:Eb

[795] Cliff B. Jones. Editorial. Formal Aspects of Computing, 26(5):861, September 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0314-y;   http://link.springer.
com/content/pdf/10.1007/s00165-
014-0314-y.pdf.

Tarlecki:2014:WMT

[796] Andrzej Tarlecki. Władysław Marek Turski (1938—2013). Formal Aspects of Computing, 26(5):863–864, September 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00165-014-0312-0;   http://link.
springer.com/article/10.1007/s00165-
014-0312-0.

Cruz-Filipe:2014:SBS

[797] Luís Cruz-Filipe, Ivan Lanese, Francisco Martins, António Ravara, and Vasco Thudichum Vasconcelos. The stream-based service-centred calculus: a foundation for service-oriented programming. Formal Aspects of Computing, 26(5):865–918, September 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-013-
0284-5.

Frappier:2014:RPA

[798] Marc Frappier, Frédéric Gervais, Régine Laleau, and Jérémy Milhau. Refinement patterns for ASTDs. Formal Aspects of Computing, 26(5): 919–941, September 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0286-3.

Yamagata:2014:FSE

[799] Yoriyuki Yamagata, Weiqiang Kong, Akira Fukuda, Nguyen Van Tang, Hitoshi Ohsaki, and Kenji Taguchi. A formal semantics of extended hierarchical state transition matrices using CSP#. Formal Aspects of Computing, 26(5): 943–962, September 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0282-7.

Banks:2014:ICF

[800] Michael J. Banks and Jeremy L. Jacob. On integrating confidentiality and functionality in a formal method. Formal Aspects of Computing, 26(5): 963–992, September 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-013-0285-4.

Ward:2014:PCD

[801] Martin Ward and Hussein Zedan. Provably correct derivation of algorithms using FermaT. Formal Aspects of Computing, 26(5):993–1031, September 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-013-
0287-2.

Ruksenas:2014:CHE

[802] Rimvydas Rukšėnas, Paul Curzon, Ann Blandford, and Jonathan Back. Combining human error verification and timing analysis: a case study on an infusion pump. Formal Aspects of Computing, 26(5):1033–1076, September 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-013-
0288-1.

Riccobene:2014:FFS

[803] Elvinia Riccobene and Patrizia Scandurra. A formal framework for service modeling and prototyping. Formal Aspects of Computing, 26(6):

1077–1113, November 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-013-0289-0.

Rossini:2014:FDM

[804] Alessandro Rossini, Juan de Lara, Esther Guerra, Adrian Rutle, and Uwe Wolter. A formalisation of deep metamodelling. Formal Aspects of Computing, 26(6):1115–1152, November 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-014-
                                        0307-x.

Gunicen:2014:RBP

[805] Canan Güniçen, Kemal İnan, Uraz Cengiz Türker, and Hüsnü Yenigün. The relation between preset distinguishing sequences and synchronizing sequences. Formal Aspects of Computing, 26(6): 1153–1167, November 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-014-0297-8.

Su:2014:ANM

[806] Li Su, Rodolfo Gomez, and Howard Bowman. Analysing neurobiological models using communicating automata. Formal Aspects of Computing, 26(6):1169–1204, November 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-014-
                                        0294-y.

Daum:2014:CUU

[807] Matthias Daum, Nelson Billing, and Gerwin Klein. Concerned with the unprivileged: user programs in kernel refinement. Formal Aspects of Computing, 26(6):1205–1229, November 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0296-9.

Wu:2014:FRD

[808] Nicolas Wu and Andrew Simpson. Formal relational database design: an exercise in extending the formal template language. Formal Aspects of Computing, 26(6):1231–1269, November 2014. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0299-6.

Bonakdarpour:2015:SBT

[809] Borzoo Bonakdarpour and Sandeep S. Kulkarni. Synthesizing bounded-time 2-phase fault recovery. Formal Aspects of Computing, 27(1):1–31, January 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0325-8.

Yazid:2015:MEI

[810] Mohand Yazid, Djamil Aïssani, Louiza Bouallouche-Medjkoune, Nassim Amrouche, and Kamel Bakli. Modeling and enhancement of the IEEE 802.11 RTS/CTS scheme in an error-prone channel. Formal Aspects of Computing, 27(1):33–52, January 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-014-0300-4.

Tarasyuk:2015:ISR

[811] Anton Tarasyuk, Elena Troubitsyna, and Linas Laibinis. Integrating stochastic reasoning into Event-B development. Formal Aspects of Computing, 27(1):53–77, January 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00165-014-0305-
                                        z.

Elleuch:2015:FPA

[812] Maissa Elleuch, Osman Hasan, Sofiène Tahar, and Mohamed Abid. Formal probabilistic analysis of detection properties in wireless sensor networks. Formal Aspects of Computing, 27(1):79–102, January 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-014-0304-0.

Ma:2015:VDS

[813] Qian Ma, Zhenhua Duan, Nan Zhang, and Xiaobing Wang. Verification of distributed systems with the axiomatic system of MSVL. Formal Aspects of Computing, 27(1):103–131, January 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-014-
                                        0303-1.

Zhu:2015:DSA

[814] H. Zhu, Jifeng He, Shengchao Qin, and Phillip J. Brooke. Denotational semantics and its algebraic derivation for an event-driven system-level language. Formal Aspects of Computing, 27(1):133–166, January 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-014-0309-8.

Cheng:2015:UFR

[815] Shu Cheng, Jim Woodcock, and Deepak D’Souza. Using formal reasoning on a model of tasks for FreeR-TOS. Formal Aspects of Computing, 27(1):167–192, January 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-014-0308-9.

Lano:2015:FMT

[816] Kevin Lano, T. Clark, and S. Kolahdouz-Rahimi. A framework for model transformation verification. Formal Aspects of Computing, 27(1):193–235, January 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0313-z.

Jones:2015:MPH

[817] Cliff B. Jones. In memoriam: Professor Heinz Zemanek (1920–2014). Formal Aspects of Computing, 27(2):237, March 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-015-
0332-4;   http://link.springer.
com/content/pdf/10.1007/s00165-
015-0332-4.pdf.

Zave:2015:PCA

[818] P. Zave. A practical comparison of Alloy and Spin. Formal Aspects of Computing, 27(2):239–253, March 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-014-
                                        0302-2.

Li:2015:FPA

[819] Yongjian Li and Jun Pang. Formalizing provable anonymity in Isabelle/ HOL. Formal Aspects of Computing, 27(2):255–282, March 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-014-0315-x.

Macedo:2015:LAA

[820] Hugo Daniel Macedo and José Nuno Oliveira. A linear algebra approach to OLAP. Formal Aspects of Computing, 27(2):283–307, March 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-014-0316-9.

Sotudeh:2015:AAF

[821] Gholamreza Sotudeh and Ali Movaghar. Abstraction and approximation in fuzzy temporal logics and models. Formal Aspects of Computing, 27 (2):309–334, March 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-014-0318-7.

Mammar:2015:PBV

[822] Amel Mammar and Marc Frappier. Proof-based verification approaches for dynamic properties: application to the information system domain. Formal Aspects of Computing, 27(2):335–374, March 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0323-x.

Madeira:2015:RHI

[823] Alexandre Madeira, Manuel A. Martins, Luís S. Barbosa, and Rolf Hennicker. Refinement in hybridised institutions. Formal Aspects of Computing, 27(2):375–395, March 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-014-0327-6.

Song:2015:MCD

[824] Fu Song and Tayssir Touili. Model checking dynamic pushdown networks. Formal Aspects of Computing, 27 (2):397–421, March 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-014-0330-y.

Zeyda:2015:LMB

[825] Frank Zeyda and Ana Cavalcanti. Laws of mission-based programming. Formal Aspects of Computing, 27(2):423–472, March 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0317-8.

Eleftherakis:2015:E

[826] George Eleftherakis, Michael Butler, and Mike Hinchey. Editorial. Formal Aspects of Computing, 27(3):473, May 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-015-
0335-1;   http://link.springer.
com/content/pdf/10.1007/s00165-
015-0335-1.pdf.

Jones:2015:BEF

[827] Cliff B. Jones, Ian J. Hayes, and Robert J. Colvin. Balancing expressiveness in formal approaches to concurrency. Formal Aspects of Computing, 27(3):475–497, May 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-014-0310-2;   http:
//link.springer.com/content/pdf/
10.1007/s00165-014-0310-2.pdf.

Fathabadi:2015:LTS

[828] Asieh Salehi Fathabadi, Michael Butler, and Abdolbaghi Rezazadeh. Language and tool support for event refinement structures in Event-B. Formal Aspects of Computing, 27(3):499–523, May 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0311-1.

Ciobanu:2015:SBS

[829] Gabriel Ciobanu, Maciej Koutny, and Jason Steggles. Strategy based semantics for mobility with time and access permissions. Formal Aspects of Computing, 27(3):525–549, May 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0324-9.

Din:2015:CRA

[830] Crystal Chang Din and Olaf Owe. Compositional reasoning about active objects with shared futures. Formal Aspects of Computing, 27(3):551–572, May 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0322-y.

Kirchner:2015:FCS

[831] Florent Kirchner, Nikolai Kosmatov, Virgile Prevosto, Julien Signoles, and Boris Yakobowski. Frama-C: A software analysis perspective. Formal Aspects of Computing, 27(3):573–609, May 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0326-7.

Butler:2015:E

[832] Michael Butler, Einar Broch Johnsen, and Luigia Petre. Editorial. Formal Aspects of Computing, 27(4):611–612, July 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-015-
0338-y;   http://link.springer.
com/content/pdf/10.1007/s00165-
015-0338-y.pdf.

Laneve:2015:ATW

[833] Cosimo Laneve and Luca Padovani. An algebraic theory for web service contracts. Formal Aspects of Computing, 27(4):613–640, July 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-015-0334-2.

Taha:2015:CAB

[834] Safouan Taha, Jacques Julliand, Frédéric Dadeau, Kalou Cabrera Castillos, and Bilal Kanso. A compositional automata-based semantics and preserving transformation rules for testing property patterns. Formal Aspects of Computing, 27(4):641–664, July 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0328-5.

Albert:2015:QAC

[835] Elvira Albert, Jesús Correas, Germán Puebla, and Guillermo Román-Díez. Quantified abstract configurations of distributed systems. Formal Aspects of Computing, 27(4):665–699, July 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0321-z.

Lucanu:2015:PEC

[836] Dorel Lucanu and Vlad Rusu. Program equivalence by circular reasoning. Formal Aspects of Computing, 27(4):701–726, July 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-014-0319-6.

Olderog:2015:STD

[837] Ernst-Rüdiger Olderog and Mani Swaminathan. Structural transformations for data-enriched real-time systems. Formal Aspects of Computing, 27(4):727–750, July 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-014-0306-y.

Woodcock:2015:E

[838] Jim Woodcock and Cliff Jones. Editorial. Formal Aspects of Computing, 27(5–6):751–752, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00165-
015-0344-0.pdf.

Polyvyanyy:2015:UNA

[839] Artem Polyvyanyy, Marcello La Rosa, Chun Ouyang, and Arthur H. M. ter Hofstede. Untanglings: a novel approach to analyzing concurrent systems. Formal Aspects of Computing, 27(5–6):753–788, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-014-
0329-4.

Dabaghchian:2015:MCO

[840] Maryam Dabaghchian and Mohammad Abdollahi Azgomi. Model checking the observational determinism security property using PROMELA and SPIN. Formal Aspects of Computing, 27(5–6):789–804, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-014-
                                        0331-x.

Rebiha:2015:GIN

[841] Rachid Rebiha, Arnaldo Vieira Moura, and Nadir Matringe. Generating invariants for non-linear loops by linear algebraic methods. Formal Aspects of Computing, 27(5–6):805–829, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-015-
                                        0333-3.

Castro:2015:CFS

[842] Pablo F. Castro, Nazareno Aguirre, Carlos L. Pombo, and T. S. E. Maibaum. Categorical foundations for structured specifications in Z. Formal Aspects of Computing, 27(5–6): 831–865, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-015-0336-0.

Rosaci:2015:FSA

[843] Domenico Rosaci. Finding semantic associations in hierarchically structured groups of Web data. Formal Aspects of Computing, 27(5–6):867–884, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-015-
                                        0337-z.

Pasqualin:2015:TCA

[844] Douglas Pereira Pasqualin, Juliana Kaizer Vizzotto, and Eduardo Kessler Piveta.

Typed context awareness Ambient Calculus for pervasive applications. Formal Aspects of Computing, 27(5–6): 885–916, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-015-0339-x.

Al-Bataineh:2015:AWC

[845] Omar Al-Bataineh, Mark Reynolds, and Tim French. Accelerating worst case execution time analysis of timed automata models with cyclic behaviour. Formal Aspects of Computing, 27(5–6):917–949, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-015-
0340-4.

Bouneb:2015:RML

[846] Messaouda Bouneb, Djamel Eddine Saidouni, and Jean Michel Ilie. A reduced maximality labeled transition system generation for recursive Petri nets. Formal Aspects of Computing, 27(5–6):951–973, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-015-
0341-3.

Mota:2015:MCC

[847] A. Mota, A. Farias, J. Woodcock, and P. G. Larsen. Model checking CML: tool development and industrial applications. Formal Aspects of Computing, 27(5–6):975–1001, November 2015. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00165-015-
                                        0342-2.

Ipate:2016:UIC

[848] Florentin Ipate and Dimitris Dranidis. A unified integration and component testing approach from deterministic stream X-machine specifications. Formal Aspects of Computing, 28(1):1–20, March 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00165-015-0345-
                                        z;   http://link.springer.com/article/
                                        10.1007/s00165-015-0345-z.

Khakpour:2016:SSB

[849] Narges Khakpour, Farhad Arbab, and Eric Rutten. Synthesizing structural and behavioral control for reconfigurations in component-based systems. Formal Aspects of Computing, 28(1):21–43, March 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00165-015-0346-
                                        y;   http://link.springer.com/article/
                                        10.1007/s00165-015-0346-y.

Barnat:2016:ASR

[850] Jirí Barnat, Petr Bauch, Nikola Benes, Lubos Brim, Jan Beran, and Tomás Kratochvíla. Analysing sanity of requirements for avionics systems. Formal Aspects of Computing, 28(1):45–63, March 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00165-015-0348-
                                        9;   http://link.springer.com/article/
                                        10.1007/s00165-015-0348-9.

Paiva:2016:GCT

[851] Sofia Costa Paiva and Adenilso Simao. Generation of complete test suites from mealy input/output transition systems. Formal Aspects of Computing, 28(1):65–78, March 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-015-0350-
2;   http://link.springer.com/article/
10.1007/s00165-015-0350-2.

Wiik:2016:CBV

[852] Jonatan Wiik and Pontus Boström. Contract-based verification of MATLABstyle matrix programs. Formal Aspects of Computing, 28(1):79–107, March 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00165-015-0353-z;   http://link.
springer.com/article/10.1007/s00165-
015-0353-z.

Fioriti:2016:DPA

[853] Luis María Ferrer Fioriti, Vahid Hashemi, Holger Hermanns, and Andrea Turrini. Deciding probabilistic automata weak bisimulation: theory and practice. Formal Aspects of Computing, 28(1):109–143, March 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0356-
4;   http://link.springer.com/article/
10.1007/s00165-016-0356-4.

Vekris:2016:VEE

[854] Dimitris Vekris, Frédéric Lang, Catalin Dima, and Radu Mateescu. Verification of EB^{3} specifications using CADP. Formal Aspects of Computing, 28(1):145–178, March 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-016-0362-6.

Vekris:2016:VES

[855] Dimitris Vekris, Frédéric Lang, Catalin Dima, and Radu Mateescu. Verification of EB^{3} specifications using CADP. Formal Aspects of Computing, 28(1):145–178, March 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0362-
6;   http://link.springer.com/article/
10.1007/s00165-016-0362-6.

Butler:2016:E

[856] Michael Butler. Editorial. Formal Aspects of Computing, 28(2):179–180, April 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0368-0;   http://link.springer.
com/content/pdf/10.1007/s00165-
016-0368-0.pdf.

Giannakopoulou:2016:E

[857] Dimitra Giannakopoulou, Gwen Salaün, and Michael Butler. Editorial. Formal Aspects of Computing, 28(2):179–180, April 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0368-0;   http://link.springer.
com/content/pdf/10.1007/s00165-
016-0368-0.pdf.

Keshishzadeh:2016:FTC

[858] Sarmen Keshishzadeh and Arjan J. Mooij. Formalizing and testing the consistency of DSL transformations. Formal Aspects of Computing, 28 (2):181–206, April 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-016-0359-1;   http:
//link.springer.com/content/pdf/
10.1007/s00165-016-0359-1.pdf.

Attie:2016:GFA

[859] Paul Attie, Eduard Baranov, Simon Bliudze, Mohamad Jaber, and Joseph Sifakis. A general framework for architecture composability. Formal Aspects of Computing, 28(2):207–231, April 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-015-
0349-8;   http://link.springer.
com/content/pdf/10.1007/s00165-
015-0349-8.pdf.

Cassel:2016:ALE

[860] Sofia Cassel, Falk Howar, Bengt Jonsson, and Bernhard Steffen. Active learning for extended finite state machines. Formal Aspects of Computing, 28(2):233–263, April 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0355-
5;   http://link.springer.com/article/
10.1007/s00165-016-0355-5.

Armstrong:2016:BPC

[861] Alasdair Armstrong, Victor B. F. Gomes, and Georg Struth. Building program construction and verification tools from algebraic principles. Formal Aspects of Computing, 28(2):265–293, April 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-015-0343-
1;   http://link.springer.com/article/
10.1007/s00165-015-0343-1.

Dobrikov:2016:OPM

[862] Ivaylo Dobrikov and Michael Leuschel. Optimising the ProB model checker for B using partial order reduction. Formal Aspects of Computing, 28 (2):295–323, April 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-015-0351-
1;   http://link.springer.com/article/
10.1007/s00165-015-0351-1.

Hesselink:2016:CCC

[863] Wim H. Hesselink. Correctness and concurrent complexity of the Black– White Bakery Algorithm. Formal Aspects of Computing, 28(2):325–341, April 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0364-4;   http://link.springer.
com/content/pdf/10.1007/s00165-
016-0364-4.pdf.

Merz:2016:Ea

[864] Stephan Merz, Jun Pang, and Jin Song Dong. Editorial. Formal Aspects of Computing, 28(3):343–344, May 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0378-y;   http://link.springer.
com/content/pdf/10.1007/s00165-
016-0378-y.pdf.

Merz:2016:Eb

[865] Stephan Merz, Jun Pang, and Jin Song Dong. Editorial. Formal Aspects of Computing, 28(3):343–344, May 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0378-y;   http://link.springer.
com/content/pdf/10.1007/s00165-
016-0378-y.pdf.

Molnar:2016:CWI

[866] Vince Molnár, András Vörös, Dániel Darvas, Tamás Bartha, and István Majzik. Component-wise incremental LTL model checking. Formal Aspects of Computing, 28(3): 345–379, May 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-015-0347-
x;   http://link.springer.com/article/
10.1007/s00165-015-0347-x.

Boulgakov:2016:CMW

[867] Alexandre Boulgakov, Thomas Gibson-Robinson, and A. W. Roscoe. Computing maximal weak and other bisimulations. Formal Aspects of Computing, 28(3):381–407, May 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0366-

2;   http://link.springer.com/article/
10.1007/s00165-016-0366-2.

Hamiaz:2016:CCM

[868] Mounira Kezadri Hamiaz, Marc Pantel, Xavier Thirioux, and Benoit Combemale. Correct-by-construction model driven engineering composition operators. Formal Aspects of Computing, 28(3):409–440, May 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0354-
6;   http://link.springer.com/article/
10.1007/s00165-016-0354-6.

Bagheri:2016:MDS

[869] Hamid Bagheri and Kevin Sullivan. Model-driven synthesis of formally precise, stylized software architectures. Formal Aspects of Computing, 28 (3):441–467, May 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0360-
8;   http://link.springer.com/article/
10.1007/s00165-016-0360-8.

Ciobaca:2016:LIP

[870] Stefan Ciobâca, Dorel Lucanu, Vlad Rusu, and Grigore Rosu. A language-independent proof system for full program equivalence. Formal Aspects of Computing, 28(3): 469–497, May 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0361-
7;   http://link.springer.com/article/
10.1007/s00165-016-0361-7.

Feo-Arenis:2016:RTE

[871] Sergio Feo-Arenis, Bernd Westphal, Daniel Dietsch, Marco Muñiz, Siyar Andisha, and Andreas Podelski. Ready for testing: ensuring conformance to industrial standards through formal verification. Formal Aspects of Computing, 28(3):499–527, May 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0365-
3;   http://link.springer.com/article/
10.1007/s00165-016-0365-3.

Carbone:2016:E

[872] Marco Carbone, Thomas Hildebrandt, Joachim Parrow, and Matthias Weidlich. Editorial. Formal Aspects of Computing, 28(4):529–530, July 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0386-y;   http://link.springer.
com/content/pdf/10.1007/s00165-
016-0386-y.pdf.

DiGiusto:2016:EBR

[873] Cinzia Di Giusto and Jorge A. Pérez. Event-based run-time adaptation in communication-centric systems. Formal Aspects of Computing, 28(4):531–566, July 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-016-0377-z;   http:
//link.springer.com/content/pdf/
10.1007/s00165-016-0377-z.pdf.

Arcaini:2016:ABF

[874] Paolo Arcaini, Roxana-Maria Holom, and Elvinia Riccobene. ASMbased formal design of an adaptivity component for a cloud system. Formal Aspects of Computing, 28 (4):567–595, July 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0371-
5;   http://link.springer.com/article/
10.1007/s00165-016-0371-5.

Polyvyanyy:2016:EPB

[875] Artem Polyvyanyy, Abel Armas-Cervantes, Marlon Dumas, and Luciano García-Bañuelos. On the expressive power of behavioral profiles. Formal Aspects of Computing, 28(4):597–613, July 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0372-
4;   http://link.springer.com/article/
10.1007/s00165-016-0372-4.

Montali:2016:MCP

[876] Marco Montali and Andrey Rivkin. Model checking Petri nets with names using data-centric dynamic systems. Formal Aspects of Computing, 28 (4):615–641, July 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0370-
6;   http://link.springer.com/article/
10.1007/s00165-016-0370-6.

Ghilezan:2016:DRA

[877] Silvia Ghilezan, Svetlana Jaksić, Jovanka Pantović, Jorge A. Pérez, and Hugo Torres Vieira. Dynamic role authorization in multiparty conversations. Formal Aspects of Computing, 28(4):643–667, July 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-016-0363-5;   http:
//link.springer.com/content/pdf/
10.1007/s00165-016-0363-5.pdf.

Castellani:2016:SAS

[878] Ilaria Castellani, Mariangiola Dezani-Ciancaglini, and Jorge A. Pérez. Selfadaptation and secure information flow in multiparty communications. Formal Aspects of Computing, 28(4):669–696, July 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0381-3;   http://link.springer.
com/content/pdf/10.1007/s00165-
016-0381-3.pdf.

Barbanera:2016:RCS

[879] Franco Barbanera, Mariangiola Dezani-Ciancaglini, and Ugo de’Liguoro. Reversible client/server interactions. Formal Aspects of Computing, 28 (4):697–722, July 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0358-
2;   http://link.springer.com/article/
10.1007/s00165-016-0358-2.

Merz:2016:Ec

[880] Stephan Merz, Jun Pang, and Jin Song Dong. Editorial. Formal Aspects of Computing, 28(5):723–724, September 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-016-
                                        0390-2;   http://link.springer.
                                        com/content/pdf/10.1007/s00165-
                                        016-0390-2.pdf.

Carvalho:2016:MTR

[881] Gustavo Carvalho, Ana Cavalcanti, and Augusto Sampaio. Modelling timed reactive systems from naturallanguage requirements. Formal Aspects of Computing, 28(5):725– 765, September 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00165-016-0387-
                                        x;   http://link.springer.com/article/
                                        10.1007/s00165-016-0387-x.

Jebali:2016:FMV

[882] Fatma Jebali, Frédéric Lang, and Radu Mateescu. Formal modelling and verification of GALS systems using GRL and CADP. Formal Aspects of Computing, 28(5):767– 804, September 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00165-016-0373-
                                        3;   http://link.springer.com/article/
                                        10.1007/s00165-016-0373-3.

Andre:2016:FCU

[883] Étienne André, Mohamed Mahdi Benmoussa, and Christine Choppy. Formalising concurrent UML state machines using coloured Petri nets. Formal Aspects of Computing, 28(5): 805–845, September 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0388-
9;   http://link.springer.com/article/
10.1007/s00165-016-0388-9.

Chevrou:2016:DAC

[884] Florent Chevrou, Aurélie Hurault, and Philippe Quéinnec. On the diversity of asynchronous communication. Formal Aspects of Computing, 28(5): 847–879, September 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0379-
x;   http://link.springer.com/article/
10.1007/s00165-016-0379-x.

Siddique:2016:FAG

[885] Umair Siddique and Sofiène Tahar. On the formal analysis of Gaussian optical systems in HOL. Formal Aspects of Computing, 28(5): 881–907, September 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0367-
1;   http://link.springer.com/article/
10.1007/s00165-016-0367-1.

Hoang:2016:FUL

[886] Thai Son Hoang, Steve Schneider, Helen Treharne, and David M. Williams. Foundations for using linear temporal logic in Event-B refinement. Formal Aspects of Computing, 28(6): 909–935, November 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-016-0376-0;   http:
//link.springer.com/content/pdf/
10.1007/s00165-016-0376-0.pdf.

Oliveira:2016:RDC

[887] M. V. M. Oliveira, P. Antonino, R. Ramos, A. Sampaio, A. Mota, and A. W. Roscoe. Rigorous development of component-based systems using component metadata and patterns. Formal Aspects of Computing, 28(6): 937–1004, November 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00165-016-0375-
                                        1;   http://link.springer.com/article/
                                        10.1007/s00165-016-0375-1.

Aransay:2016:FCE

[888] Jesús Aransay and Jose Divasón. Formalisation of the computation of the echelon form of a matrix in Isabelle/HOL. Formal Aspects of Computing, 28(6):1005– 1026, November 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00165-016-0383-
                                        1;   http://link.springer.com/article/
                                        10.1007/s00165-016-0383-1.

Kirwan:2016:MCL

[889] Ryan Kirwan, Alice Miller, and Bernd Porr. Model checking learning agent systems using Promela with embedded C code and abstraction. Formal Aspects of Computing, 28(6): 1027–1056, November 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-016-0382-2;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-016-0382-2.pdf.

Hayes:2016:GRG

[890] Ian J. Hayes. Generalised relyguarantee concurrency: an algebraic foundation. Formal Aspects of Computing, 28(6):1057– 1078, November 2016. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0384-
0;   http://link.springer.com/article/
10.1007/s00165-016-0384-0.

Proietti:2017:E

[891] Maurizio Proietti, Hirohisa Seki, and Jim Woodcock. Editorial. Formal Aspects of Computing, 29(1):1–2, January 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0408-9;   http://link.springer.
com/content/pdf/10.1007/s00165-
016-0408-9.pdf.

Giacobazzi:2017:MIO

[892] Roberto Giacobazzi, Isabella Mastroeni, and Mila Dalla Preda. Maximal incompleteness as obfuscation potency. Formal Aspects of Computing, 29(1):3–31, January 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0374-
2;   http://link.springer.com/article/
10.1007/s00165-016-0374-2.

Chawdhary:2017:PES

[893] Aziem Chawdhary, Ranjeet Singh, and Andy King. Partial evaluation of string obfuscations for Java malware detection. Formal Aspects of Computing, 29(1):33–55, January 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0357-
3;   http://link.springer.com/article/
10.1007/s00165-016-0357-3.

Christiansen:2017:PCM

[894] Henning Christiansen and Maja H. Kirkeby. On proving confluence modulo equivalence for constraint handling rules. Formal Aspects of Computing, 29(1):57–95, January 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0396-
9;   http://link.springer.com/article/
10.1007/s00165-016-0396-9.

Arias:2017:CLP

[895] Emilio Jesús Gallego Arias, James Lipton, and Julio Mariño. Constraint logic programming with a relational machine. Formal Aspects of Computing, 29(1):97–124, January 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0369-
z;   http://link.springer.com/article/
10.1007/s00165-016-0369-z.

Nys:2017:ACP

[896] Vincent Nys and Danny De Schreye. Abstract conjunctive partial deduction for the analysis and compilation of coroutines. Formal Aspects of Computing, 29(1): 125–153, January 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/

article/10.1007/s00165-016-0389-
8;   http://link.springer.com/article/
10.1007/s00165-016-0389-8.

Drabent:2017:PCL

[897] Włodzimierz Drabent. Proving completeness of logic programs with the cut. Formal Aspects of Computing, 29 (1):155–172, January 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0392-
0;   http://link.springer.com/article/
10.1007/s00165-016-0392-0.

Anonymous:2017:OAL

[898] Anonymous. [Obituary:] Amílcar Sernadas. Formal Aspects of Computing, 29(2):173, March 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0423-5;   http:
//link.springer.com/content/pdf/
10.1007/s00165-017-0423-5.pdf.

Bjorner:2017:MDA

[899] Dines Bjørner. Manifest domains: analysis and description. Formal Aspects of Computing, 29(2): 175–225, March 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0385-
z;   http://link.springer.com/article/
10.1007/s00165-016-0385-z.

Li:2017:RAA

[900] Qin Li and Graeme Smith. Refining autonomous agents with declarative beliefs and desires. Formal Aspects of Computing, 29(2):

227–249, March 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0391-
1;   http://link.springer.com/article/
10.1007/s00165-016-0391-1.

Bhattacharyya:2017:ECF

[901] Anirban Bhattacharyya, Andrey Mokhov, and Ken Pierce. An empirical comparison of formalisms for modelling and analysis of dynamic reconfiguration of dependable systems. Formal Aspects of Computing, 29(2):251–307, March 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0405-z;   http://link.springer.
com/content/pdf/10.1007/s00165-
016-0405-z.pdf.

Riesco:2017:MEC

[902] Adrián Riesco, Kazuhiro Ogata, and Kokichi Futatsugi. A Maude environment for CafeOBJ. Formal Aspects of Computing, 29(2): 309–334, March 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0398-
7;   http://link.springer.com/article/
10.1007/s00165-016-0398-7.

Huang:2017:CMB

[903] Wen ling Huang and Jan Peleska. Complete model-based equivalence class testing for nondeterministic systems. Formal Aspects of Computing, 29 (2):335–364, March 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0402-
2;   http://link.springer.com/article/
10.1007/s00165-016-0402-2.

Banerjee:2017:DBR

[904] Kunal Banerjee, Dipankar Sarkar, and Chittaranjan Mandal. Deriving bisimulation relations from path based equivalence checkers. Formal Aspects of Computing, 29(2): 365–379, March 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00165-016-0406-
y;   http://link.springer.com/article/
10.1007/s00165-016-0406-y.

Falaschi:2017:E

[905] Moreno Falaschi and Augusto Sampaio. Editorial. Formal Aspects of Computing, 29(3):381–382, May 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL https://link.springer.com/
content/pdf/10.1007/s00165-017-
0427-1.pdf.

Miller:2017:PCL

[906] Dale Miller. Proof checking and logic programming. Formal Aspects of Computing, 29(3):383–399, May 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Seki:2017:DPC

[907] Hirohisa Seki. On dual programs in co-logic programming and the Horn μcalculus. Formal Aspects of Computing, 29(3):401–421, May 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Meseguer:2017:EFP

[908] José Meseguer and Stephen Skeirik. Equational formulas and pattern operations in initial order-sorted algebras. Formal Aspects of Computing, 29(3):423–452, May 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Fu:2017:OSR

[909] Peng Fu and Ekaterina Komendantskaya. Operational semantics of resolution and productivity in Horn clause logic. Formal Aspects of Computing, 29(3):453–474, May 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL https://link.springer.com/
                                        content/pdf/10.1007/s00165-016-
                                        0403-1.pdf.

Antoy:2017:TBE

[910] Sergio Antoy and Michael Hanus. Transforming Boolean equalities into constraints. Formal Aspects of Computing, 29(3):475–494, May 2017. CO-DEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Chaudhari:2017:APT

[911] Dipak L. Chaudhari and Om Damani. Assumption propagation through annotated programs. Formal Aspects of Computing, 29(3):495–530, May 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Comini:2017:PAF

[912] Marco Comini, María del Mar Gallardo, Laura Titolo, and Alicia Villanueva. A program analysis framework for tccp based on abstract interpretation. Formal Aspects of Computing, 29(3):531–557, May 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Codish:2017:OSA

[913] Michael Codish, Luís Cruz-Filipe, Markus Nebel, and Peter Schneider-Kamp. Optimizing sorting algorithms by using sorting networks. Formal Aspects of Computing, 29(3):559–579, May 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Li:2017:E

[914] Xuandong Li and Zhiming Liu. Editorial. Formal Aspects of Computing, 29(4):581–582, July 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/content/pdf/
10.1007/s00165-017-0431-5.pdf.

Bochmann:2017:SVC

[915] Gregor V. Bochmann, Martin Hilscher, Sven Linker, and Ernst-Rüdiger Olderog. Synthesizing and verifying controllers for multi-lane traffic maneuvers. Formal Aspects of Computing, 29(4):583–600, July 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Faitelson:2017:IDD

[916] David Faitelson and Shmuel Tyszberowicz. Improving design decomposition (extended version). Formal Aspects of Computing, 29(4):601–627, July 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Hatefi:2017:CVT

[917] Hassan Hatefi, Ralf Wimmer, Bettina Braitling, Luis María Ferrer Fioriti, Bernd Becker, and Holger Hermanns. Cost vs. time in stochastic games and Markov automata. Formal Aspects of Computing, 29(4):629–649, July 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Junges:2017:FTD

[918] Sebastian Junges, Dennis Guck, Joost-Pieter Katoen, Arend Rensink, and Mariëlle Stoelinga. Fault trees on a diet: automated reduction by graph rewriting. Formal Aspects of Computing, 29(4):651–703, July 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Moszkowski:2017:ATP

[919] Ben Moszkowski and Dimitar P. Guelev. An application of temporal projection to interleaving concurrency. Formal Aspects of Computing, 29(4):705–750, July 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Wang:2017:CMV

[920] Shuling Wang, Naijun Zhan, and Lijun Zhang. A compositional modelling and verification framework for stochastic hybrid systems. Formal Aspects of Computing, 29(4):751–775, July 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Jancik:2017:PSM

[921] Pavel Jancík and Jan Kofron. On partial state matching. Formal Aspects of Computing, 29(5):777–803, September 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0413-z.

Gaina:2017:BSC

[922] Daniel Gaina. Birkhoff style calculi for hybrid logics. Formal Aspects of Computing, 29(5):805–832, September 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0414-y.

Hesselink:2017:TME

[923] Wim H. Hesselink. Tournaments for mutual exclusion: verification and concurrent complexity. Formal Aspects of Computing, 29(5):833–852, September 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-016-
0407-x;   https://link.springer.
com/content/pdf/10.1007/s00165-
016-0407-x.pdf.

Colvin:2017:DSM

[924] Robert J. Colvin, Ian J. Hayes, and Larissa A. Meinicke. Designing a semantic model for a wide-spectrum language with concurrency. Formal Aspects of Computing, 29(5): 853–875, September 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0416-4.

Neykova:2017:TRM

[925] Rumyana Neykova, Laura Bocchi, and Nobuko Yoshida. Timed runtime monitoring for multiparty conversations. Formal Aspects of Computing, 29(5): 877–910, September 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0420-8;   https:
//link.springer.com/content/pdf/
10.1007/s00165-017-0420-8.pdf.

Schrammel:2017:IBM

[926] Peter Schrammel, Daniel Kroening, Martin Brain, Ruben Martins, Tino Teige, and Tom Bienmüller. Incremental bounded model checking for embedded software. Formal Aspects of Computing, 29(5):911–931, September 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0419-1;   https://link.springer.
com/content/pdf/10.1007/s00165-
017-0419-1.pdf.

Anonymous:2017:MJC

[927] Anonymous. Michael J. C. Gordon, FRS, Professor of Computer Assisted Reasoning (28 February 1948– 22 August 2017). Formal Aspects of Computing, 29(6):933, November 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0438-y;   https://link.springer.
com/content/pdf/10.1007/s00165-
017-0438-y.pdf.

Smith:2017:RTR

[928] Graeme Smith and Kirsten Winter. Relating trace refinement and linearizability. Formal Aspects of Computing, 29 (6):935–950, November 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0418-2.

Nazarpour:2017:CPS

[929] Hosein Nazarpour, Yliès Falcone, Saddek Bensalem, and Marius Bozga. Concurrency-preserving and sound monitoring of multi-threaded componentbased systems: theory, algorithms, implementation, and evaluation. Formal Aspects of Computing, 29(6): 951–986, November 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0422-6.

Graiet:2017:VDA

[930] Mohamed Graiet, Lazhar Hamel, Amel Mammar, and Samir Tata. A verification and deployment approach for elastic component-based applications. Formal Aspects of Computing, 29(6): 987–1011, November 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0425-3.

Demasi:2017:SRF

[931] Ramiro Demasi, Pablo F. Castro, Thomas S. E. Maibaum, and Nazareno Aguirre. Simulation relations for fault-tolerance. Formal Aspects of Computing, 29(6):1013–1050, November 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0426-2.

Yousefi:2017:MEV

[932] Behnaz Yousefi, Fatemeh Ghassemi, and Ramtin Khosravi. Modeling and efficient verification of wireless ad hoc networks. Formal Aspects of Computing, 29(6):1051–1086, November 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0429-z.

Eggert:2017:DIN

[933] Sebastian Eggert and Ron van der Meyden. Dynamic intransitive noninterference revisited. Formal Aspects of Computing, 29(6):1087–1120, November 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0430-6.

Jones:2017:BRT

[934] Cliff B. Jones. Book review: The Turing Guide, By Jack Copeland, Jonathan Bowen, Mark Sprevak, Robin Wilson and others. Oxford University Press, Oxford, UK, 26 January 2017, xv + 576 pp, 246 × 189 mm, ISBN: 978-0-19-874782-6 (Hardback, $75.00), ISBN: 978-0-19-874783-3 (Paperback, $19.99). Formal Aspects of Computing, 29(6):1121–1122, November 2017. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0446-y.

Denney:2018:E

[935] Ewen Denney, Perdita Stevens, and Andrzej Wasowski. Editorial. Formal Aspects of Computing, 30(1):1, January 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0448-9;   https://link.springer.
com/content/pdf/10.1007/s00165-
017-0448-9.pdf.

dePutter:2018:FVT

[936] Sander de Putter and Anton Wijs. A formal verification technique for behavioural model-to-model transformations. Formal Aspects of Computing, 30(1):3–43, January 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0437-z;   https:
//link.springer.com/content/pdf/
10.1007/s00165-017-0437-z.pdf.

Chrszon:2018:PFO

[937] Philipp Chrszon, Clemens Dubslaff, Sascha Klüppelholz, and Christel Baier. ProFeat: feature-oriented engineering for family-based probabilistic model checking. Formal Aspects of Computing, 30(1):45–75, January 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0432-4.

Gerhold:2018:MBT

[938] Marcus Gerhold and Mariëlle Stoelinga. Model-based testing of probabilistic systems. Formal Aspects of Computing, 30(1):77–106, January 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0440-4;   https:
//link.springer.com/content/pdf/
10.1007/s00165-017-0440-4.pdf.

Lechenet:2018:CBB

[939] Jean-Christophe Léchenet, Nikolai Kosmatov, and Pascale Le Gall. Cut branches before looking for bugs: certifiably sound verification on relaxed slices. Formal Aspects of Computing, 30(1):107–131, January 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0439-x.

Struber:2018:VBM

[940] D. Strüber, J. Rubin, T. Arendt, M. Chechik, G. Taentzer, and J. Plöger. Variability-based model transformation: formal foundation and application. Formal Aspects of Computing, 30(1):133–162, January 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0441-3.

Corrodi:2018:SCW

[941] Claudio Corrodi, Alexander Heußner, and Christopher M. Poskitt. A semantics comparison workbench for a concurrent, asynchronous, distributed programming language. Formal Aspects of Computing, 30(1):163–192, January 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0443-1.

Li:2018:ETS

[942] Jianwen Li, Lijun Zhang, Shufang Zhu, Geguang Pu, Moshe Y. Vardi, and Jifeng He. An explicit transition system construction approach to LTL satisfiability checking. Formal Aspects of Computing, 30(2):193–217, March 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0442-2.

Zhang:2018:FVQ

[943] Hui Zhang and Jinzhao Wu. Formal verification and quantitative metrics of MPSoC data dynamics. Formal Aspects of Computing, 30(2):219–237, March 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0444-0.

Pedersen:2018:SCV

[944] Jan B. Pedersen and Peter H. Welch. The symbiosis of concurrency and verification: teaching and case studies. Formal Aspects of Computing, 30(2):239–277, March 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0447-x;   https:
//link.springer.com/content/pdf/
10.1007/s00165-017-0447-x.pdf.

Sisto:2018:FSI

[945] Riccardo Sisto, Piergiuseppe Bettassa Copet, Matteo Avalle, and Alfredo Pironti. Formally sound implementations of security protocols with JavaSPI. Formal Aspects of Computing, 30(2):279–317, March 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0449-8.

El-Fakih:2018:ADT

[946] Khaled El-Fakih, Nina Yevtushenko, and Natalia Kushik. Adaptive distinguishing test cases of nondeterministic finite state machines: test case derivation and length estimation. Formal Aspects of Computing, 30(2):319–332, March 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0450-2.

Santos:2018:ASM

[947] Filipe Santos, Krystian Kwiecinski, Ana de Almeida, Sara Eloy, and Bruno Taborda. Alternative shaper: a model for automatic design generation. Formal Aspects of Computing, 30 (3–4):333–349, August 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0452-8.

Shi:2018:USC

[948] Ling Shi, Yongxin Zhao, Yang Liu, Jun Sun, Jin Song Dong, and Shengchao Qin. A UTP semantics for communicating processes with shared variables and its formal encoding in PVS. Formal Aspects of Computing, 30(3– 4):351–380, August 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0453-7.

Bao:2018:USL

[949] Yuyan Bao, Gary T. Leavens, and Gidon Ernst. Unifying separation logic and region logic to allow interoperability. Formal Aspects of Computing, 30 (3–4):381–441, August 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0455-5.

Jing:2018:CTM

[950] Yaping Jing and Andrew S. Miner. Computation tree measurement language (CTML). Formal Aspects of Computing, 30(3–4):443–462, August 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
0457-3.

Chane-Yack-Fa:2018:PVM

[951] Raphaël Chane-Yack-Fa, Marc Frappier, Amel Mammar, and Alain Finkel. Parameterized verification of monotone information systems. Formal Aspects of Computing, 30(3–4):463–489, August 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
0460-8.

Hallerstede:2018:BRT

[952] Stefan Hallerstede. Book review: Tobias Nipkow and Gerwin Klein: Concrete Semantics with Isabelle/HOL, Springer Verlag, 2014, x + 289 pp, EUR 63,29 (Hardback), ISBN 978-3-319-10542-0, http://www.concretesemantics.org/. Formal Aspects of Computing, 30(3–4):491–492, August 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
0454-6.

Bjorner:2018:E

[953] Nikolaj Bjørner, Frank de Boer, and Andrew Butterfield. Editorial. Formal Aspects of Computing, 30(5): 493–494, September 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0463-5;   https:
//link.springer.com/content/pdf/
10.1007/s00165-018-0463-5.pdf.

Polikarpova:2018:FVC

[954] Nadia Polikarpova, Julian Tschannen, and Carlo A. Furia. A fully verified container library. Formal Aspects of Computing, 30(5):495–523, September 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0435-1;   https://link.springer.
com/content/pdf/10.1007/s00165-
017-0435-1.pdf.

Bagheri:2018:FAD

[955] Hamid Bagheri, Eunsuk Kang, Sam Malek, and Daniel Jackson. A formal approach for detection of security flaws in the Android permission system. Formal Aspects of Computing, 30 (5):525–544, September 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-017-0445-z.

Schneider:2018:MBP

[956] David Schneider, Michael Leuschel, and Tobias Witt. Model-based problem solving for university timetable validation and improvement. Formal Aspects of Computing, 30(5):545–569, September 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
0461-7.

Elkader:2018:ACA

[957] Karam Abd Elkader, Orna Grumberg, Corina S. Păsăreanu, and Sharon Shoham. Automated circular assumeguarantee reasoning. Formal Aspects of Computing, 30(5):571–595, September 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0436-0.

Derrick:2018:MPO

[958] John Derrick, Simon Doherty, Brijesh Dongol, Gerhard Schellhorn, Oleg Travkin, and Heike Wehrheim. Mechanized proofs of opacity: a comparison of two techniques. Formal Aspects of Computing, 30(5):597–625, September 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-017-
0433-3;   https://link.springer.
com/content/pdf/10.1007/s00165-
017-0433-3.pdf.

Aichernig:2018:SST

[959] Bernhard K. Aichernig, Carlo A. Furia, Marie-Claude Gaudel, and Rob Hierons. Special section of Tests and Proofs 2016. Formal Aspects of Computing, 30(6):627–628, November 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
0474-2;   https://link.springer.
com/content/pdf/10.1007/s00165-
018-0474-2.pdf.

Petiot:2018:HTH

[960] Guillaume Petiot, Nikolai Kosmatov, Bernard Botella, Alain Giorgetti, and Jacques Julliand. How testing helps to diagnose proof failures. Formal Aspects of Computing, 30(6):629–657, November 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
0456-4.

Dubois:2018:TPC

[961] Catherine Dubois and Alain Giorgetti. Tests and proofs for custom data generators. Formal Aspects of Computing, 30 (6):659–684, November 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0459-1.

Bruni:2018:COA

[962] Roberto Bruni, Roberto Giacobazzi, and Roberta Gori. Code obfuscation against abstraction refinement attacks. Formal Aspects of Computing, 30(6): 685–711, November 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0462-6.

Xie:2018:UAR

[963] Wanling Xie, Shuangqing Xiang, and Huibiao Zhu. A UTP approach for rTiMo. Formal Aspects of Computing, 30(6):713–738, November 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s00165-018-0467-
1.

Shi:2018:FAK

[964] Zhiping Shi, Aixuan Wu, Xiumei Yang, Yong Guan, Yongdong Li, and Xiaoyu Song. Formal analysis of the kinematic Jacobian in screw theory. Formal Aspects of Computing, 30(6): 739–757, November 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0468-0.

Monahan:2018:BRD

[965] Rosemary Monahan. Book review: Daniel Kroening and Ofer Strichman: Decision procedures, Springer Verlag, 2016, xxi + 356 pp. ISBN 978-3-662-50496-3 (hardback, EUR 69,67), http://www.decision-procedures.org/. Formal Aspects of Computing, 30 (6):759, November 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0466-2;   https:
//link.springer.com/content/pdf/
10.1007/s00165-018-0466-2.pdf.

Bowen:2018:BRE

[966] Jonathan P. Bowen. Book review: Egon Börger and Alexander Raschke: Modeling companion for software practitioners, Springer, 2018, xxi + 349 pp, ISBN: 978-3-662-56639-8 (Paperback, £46.99), eISBN: 978-3-662-56641-1 (eBook, £36.99), https://doi.org/10.1007/978-3-662-56641-1. Formal Aspects of Computing, 30(6):761–762, November 2018. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
0472-4.

Franzle:2019:E

[967] Martin Fränzle, Deepak Kapur, Heike Wehrheim, and Naijun Zhan. Editorial. Formal Aspects of Computing, 31(1):1, February 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-00477-6;   https:
//link.springer.com/content/pdf/
10.1007/s00165-018-00477-6.pdf.

Ying:2019:TAV

[968] Mingsheng Ying. Toward automatic verification of quantum programs. Formal Aspects of Computing, 31 (1):3–25, February 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0465-3.

Mizera:2019:GAS

[969] Andrzej Mizera, Jun Pang, and Qixia Yuan. GPU-accelerated steady-state computation of large probabilistic Boolean networks. Formal Aspects of Computing, 31(1):27–46, February 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00165-018-
0470-6.

Dong:2019:EPN

[970] Xiaoju Dong, Yuxi Fu, and Daniele Varacca. Extensional Petri net. Formal Aspects of Computing, 31(1): 47–58, February 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0473-3.

Bozzano:2019:FRA

[971] Marco Bozzano, Alessandro Cimatti, and Cristian Mattarei. Formal reliability analysis of redundancy architectures. Formal Aspects of Computing, 31(1):59–94, February 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-018-0475-1.

Lin:2019:AEB

[972] Yuhui Lin, Alan Bundy, Gudmund Grov, and Ewen Maclean. Automating Event-B invariant proofs by rippling and proof patching. Formal Aspects of Computing, 31(1):95–129, February 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
00476-7;   https://link.springer.
com/content/pdf/10.1007/s00165-
018-00476-7.pdf.

Gnesi:2019:E

[973] Stefania Gnesi, Ana Cavalcanti, John Fitzgerald, and Constance Heitmeyer. Editorial. Formal Aspects of Computing, 31(2):131–132, April 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-019-
00481-4;   https://link.springer.
com/content/pdf/10.1007/s00165-
019-00481-4.pdf.

Hayes:2019:SPA

[974] Ian J. Hayes, Larissa A. Meinicke, Kirsten Winter, and Robert J. Colvin. A synchronous program algebra: a basis for reasoning about shared-memory and event-based concurrency. Formal Aspects of Computing, 31(2):133–163, April 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
0464-4.

Biondi:2019:HSE

[975] Fabrizio Biondi, Yusuke Kawamoto, Axel Legay, and Louis-Marie Traonouez. Hybrid statistical estimation of mutual information and its application to information flow. Formal Aspects of Computing, 31(2):165–206, April 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-018-
0469-z.

Ta:2019:AMI

[976] Quang-Trung Ta, Ton Chanh Le, Siau-Cheng Khoo, and Wei-Ngan Chin. Automated mutual induction proof in separation logic. Formal Aspects of Computing, 31(2):207–230, April 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00165-018-
0471-5.

Dimovski:2019:FSV

[977] Aleksandar S. Dimovski, Claus Brabrand, and Andrzej Wasowski. Finding suitable variability abstractions for lifted analysis. Formal Aspects of Computing, 31(2):231–259, April 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-019-00479-y.

Bisgaard:2019:BAS

[978] Morten Bisgaard, David Gerhardt, Holger Hermanns, Jan Krčál, Gilles Nies, and Marvin Stenger. Batteryaware scheduling in low orbit: the GomX-3 case. Formal Aspects of Computing, 31(2):261–285, April 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s00165-018-0458-
2.

Kobayashi:2019:CPR

[979] Tsutomu Kobayashi, Fuyuki Ishikawa, and Shinichi Honiden. Consistencypreserving refactoring of refinement structures in Event-B models. Formal Aspects of Computing, 31(3):287–320, June 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-019-
00478-z.

Beillahi:2019:MVF

[980] Sidi Mohamed Beillahi, Mohamed Yousri Mahmoud, and Sofiène Tahar. A modeling and verification framework for optical quantum circuits. Formal Aspects of Computing, 31(3):321–351, June 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-019-
00480-5.

Jones:2019:ILR

[981] Cliff B. Jones and Nisansala Yatapanage. Investigating the limits of rely/ guarantee relations based on a concurrent garbage collector example. Formal Aspects of Computing, 31(3):353–374, June 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-019-
00482-3;   https://link.springer.
com/content/pdf/10.1007/s00165-
019-00482-3.pdf.

Antonino:2019:EVC

[982] Pedro Antonino, Thomas Gibson-Robinson, and A. W. Roscoe. Efficient verification of concurrent systems using local-analysis-based approximations and SAT solving. Formal Aspects of Computing, 31(3):375–409, June 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-019-
00483-2;   https://link.springer.
com/content/pdf/10.1007/s00165-
019-00483-2.pdf.

Lowe:2019:DCD

[983] Gavin Lowe. Discovering and correcting a deadlock in a channel implementation. Formal Aspects of Computing, 31(4):411–419, August 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-019-
00487-y;   https://link.springer.
com/content/pdf/10.1007/s00165-
019-00487-y.pdf.

Brogi:2019:ECM

[984] Antonio Brogi, Andrea Corradini, and Jacopo Soldani. Estimating costs of multi-component enterprise applications. Formal Aspects of Computing, 31(4):421–451, August 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-019-00491-2.

Banach:2019:BRJ

[985] Richard Banach. Book review: John Fitzgerald, Peter Gorm Larsen, Marcel Verhoef (eds): Collaborative design for embedded systems, Springer, Berlin Heidelberg, 2014, xxii + 385 pp, 6×2 mm, ISBN: 978-3-642-54117-9 (hardback, $119.99), ISBN: 978-3-662-52444-2 (softcover, $129.00). Formal Aspects of Computing, 31(4):453–454, August 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-019-
00485-0.

Konnov:2019:BRE

[986] Igor Konnov. Book review: Edmund M. Clarke, Thomas A. Henzinger, Helmut Veith, and Roderick Bloem (eds): Handbook of model checking. Formal Aspects of Computing, 31(4):455–456, August 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-019-
00486-z.

Russo:2019:E

[987] Alessandra Russo, Andy Schürr, and Heike Wehrheim. Editorial. Formal Aspects of Computing, 31(5):457–458, November 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-019-
                                        00495-y;   https://link.springer.
                                        com/content/pdf/10.1007/s00165-
                                        019-00495-y.pdf.

Menghi:2019:VDF

[988] Claudio Menghi, Paola Spoletini, Marsha Chechik, and Carlo Ghezzi. A verification-driven framework for iterative design of controllers. Formal Aspects of Computing, 31(5): 459–502, November 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-019-00484-1;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-019-00484-1.pdf.

Liu:2019:RAT

[989] Si Liu, Peter Csaba Ölveczky, Qi Wang, Indranil Gupta, and José Meseguer. Read atomic transactions with prevention of lost updates: ROLA and its formal analysis. Formal Aspects of Computing, 31(5):503–540, November 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-019-
                                        00489-w.

Marmsoler:2019:IVA

[990] Diego Marmsoler and Habtom Kashay Gidey. Interactive verification of architectural design patterns in FACTum. Formal Aspects of Computing, 31(5): 541–610, November 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-019-00488-x.

Diskin:2019:MMS

[991] Zinovy Diskin, Harald König, and Mark Lawford. Multiple model synchronization with multiary delta lenses with amendment and K-Putput. Formal Aspects of Computing, 31(5): 611–640, November 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-019-00493-0. See cor-

rection [1006].

Michaelson:2019:BRB

[992] Greg Michaelson. Book review: Bernhard Steffen, Oliver Rüthing, and Michael Huth: Mathematical Foundations of Advanced Informatics — Volume 1: Inductive Approaches. Formal Aspects of Computing, 31(5): 641–642, November 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-019-00496-x.

Dershowitz:2019:ZM

[993] Nachum Dershowitz and Richard Waldinger. Zohar Manna (1939–2018). Formal Aspects of Computing, 31(6): 643–660, December 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-019-00500-4.

Jones:2019:E

[994] Cliff Jones and José Oliveira. Editorial. Formal Aspects of Computing, 31(6):661, December 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-019-00498-9;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00165-019-00498-9.pdf.

Haigh:2019:APF

[995] Thomas Haigh. Assembling a prehistory for formal methods: a personal view. Formal Aspects of Computing, 31 (6):663–674, December 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-019-00494-z.

Paulson:2019:LIH

[996] Lawrence C. Paulson, Tobias Nipkow, and Makarius Wenzel. From LCF to Isabelle/HOL. Formal Aspects of Computing, 31(6):675–698, December 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-019-
                                        00492-1;   https://link.springer.
                                        com/content/pdf/10.1007/s00165-
                                        019-00492-1.pdf.

Moore:2019:MPL

[997] J. Strother Moore. Milestones from the Pure Lisp theorem prover to ACL2. Formal Aspects of Computing, 31(6): 699–732, December 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-019-00490-3.

Pous:2019:BCE

[998] Damien Pous and Davide Sangiorgi. Bisimulation and coinduction enhancements: A historical perspective. Formal Aspects of Computing, 31(6): 733–749, December 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-019-00497-w.

Apt:2019:FYH

[999] Krzysztof R. Apt and Ernst-Rüdiger Olderog. Fifty years of Hoare’s logic. Formal Aspects of Computing, 31(6): 751–807, December 2019. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-019-00501-3.

Smith:2020:LHW

[1000] Graeme Smith, Kirsten Winter, and Robert J. Colvin. Linearizability on hardware weak memory models. Formal Aspects of Computing, 32 (1):1–32, February 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-019-00499-8.

Baresi:2020:UFV

[1001] L. Baresi, M. M. Bersani, F. Marconi, G. Quattrocchi, and M. Rossi. Using formal verification to evaluate the execution time of Spark applications. Formal Aspects of Computing, 32(1):33–70, February 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

//link.springer.com/article/10.
1007/s00165-020-00505-4.

Ahmad:2020:FRF

[1002] Waqar Ahmad, Osman Hasan, and Sofiène Tahar. Formal reliability and failure analysis of Ethernet based communication networks in a smart grid substation. Formal Aspects of Computing, 32(1):71–111, February 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-019-
                                        00503-1.

Huang:2020:MVT

[1003] Yanhong Huang, Haiping Pang, and Jianqi Shi. Modeling and verification of a timing protection mechanism in the OSEK/VDX OS using CSP. Formal Aspects of Computing, 32 (1):113–145, February 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00165-020-00511-6.

Bowen:2020:BRG

[1004] Jonathan P. Bowen. Book review: Gerard O’Regan: Concise Guide to Formal Methods: Theory, Fundamentals and Industry Applications. Formal Aspects of Computing, 32(1):147–148, February 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00165-020-
                                        00506-3.

Knapp:2020:BRA

[1005] Alexander Knapp and Markus Roggenbach. Book review: André Platzer: Logical foundations of cyber-physical systems. Formal Aspects of Computing, 32(1):149–151, February 2020.

CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL http://link.springer.com/
article/10.1007/s00165-020-00510-
7;   https://link.springer.com/
content/pdf/10.1007/s00165-020-
00510-7.pdf.

Diskin:2020:CMM

[1006] Zinovy Diskin, Harald König, and Mark Lawford. Correction to: Multiple model synchronization with multiary delta lenses with amendment and K-Putput. Formal Aspects of Computing, 32(1):153, February 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00165-019-
00502-2;   https://link.springer.
com/content/pdf/10.1007/s00165-
019-00502-2.pdf. See [991].

Cavalcanti:2020:E

[1007] Ana Cavalcanti and Pedro Ribeiro. Editorial. Formal Aspects of Computing, 32(2–3):155, July 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00516-1.

Fraser:2020:CMA

[1008] Douglas Fraser, Ruben Giaquinta, and Gethin Norman. Collaborative models for autonomous systems controller synthesis. Formal Aspects of Computing, 32(2–3):157–186, July 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00508-1.

Bersani:2020:PSR

[1009] Marcello M. Bersani, Matteo Soldo, and Matteo Rossi. PuRSUE — from specification of robotic environments to synthesis of controllers. Formal Aspects of Computing, 32 (2–3):187–227, July 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-020-00509-0.

Rashid:2020:FVR

[1010] Adnan Rashid and Osman Hasan. Formal verification of robotic cell injection systems up to 4-DOF using HOL light. Formal Aspects of Computing, 32(2–3):229–250, July 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-020-00514-3.

Williams:2020:LDD

[1011] David M. Williams, Salaheddin Darwish, and David R. Michael. Legislation-driven development of a Gift Aid system using Event-B. Formal Aspects of Computing, 32(2– 3):251–273, July 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-020-00512-5.

Sheng:2020:TPA

[1012] Feng Sheng, Huibiao Zhu, and Jonathan P. Bowen. Theoretical and practical approaches to the denotational semantics for MDESL based on UTP. Formal Aspects of Computing, 32(2–3):275–314, July 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00513-4.

Renard:2020:RET

[1013] Matthieu Renard, Antoine Rollet, and Yliès Falcone. Runtime enforcement of timed properties using games. Formal Aspects of Computing, 32(2–3):315–360, July 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00515-2.

Lahouij:2020:EBB

[1014] Aida Lahouij, Lazhar Hamel, and Béchir el Ayeb. An Event-B based approach for cloud composite services verification. Formal Aspects of Computing, 32(4–6):361– 393, November 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00517-0.

Zhuo:2020:TEP

[1015] Xilong Zhuo and Chenyi Zhang.

TFA: an efficient and precise virtual method call resolution for Java. Formal Aspects of Computing, 32(4–6): 395–416, November 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00518-z.

Chen:2020:FCP

[1016] Shanyan Chen, Guohui Wang, and Yong Guan. Formalization of camera pose estimation algorithm based on Rodrigues formula. Formal Aspects of Computing, 32(4–6):417– 437, November 2020. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00520-5.

deVink:2021:E

[1017] Erik de Vink and Ana Cavalcanti. Editorial. Formal Aspects of Computing, 33(1):1–2, January 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00522-3.

Bacci:2021:ORC

[1018] Giovanni Bacci, Patricia Bouyer, and Pierre-Alain Reynier. Optimal and robust controller synthesis using energy timed automata with uncertainty. Formal Aspects of Computing, 33(1):3–25, January 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00521-4.

Cavezza:2021:WMG

[1019] Davide G. Cavezza, Dalal Alrajeh, and András György. A weakness measure for GR(1) formulae. Formal Aspects of Computing, 33(1): 27–63, January 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00519-y.

Dutle:2021:FAC

[1020] Aaron Dutle, Mariano Moscato, and François Bobot. Formal analysis of the compact position reporting algorithm. Formal Aspects of Computing, 33(1):65–86, January 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-019-00504-0.

Geisler:2021:SDM

[1021] S. Geisler and A. E. Haxthausen.

Stepwise development and model checking of a distributed interlocking system using RAISE. Formal Aspects of Computing, 33(1): 87–125, January 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00507-2.

Letan:2021:MVP

[1022] Thomas Letan, Yann Régis-Gianas, and Guillaume Hiet. Modular verification of programs with effects and effects handlers. Formal Aspects of Computing, 33(1): 127–150, January 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00523-2.

Zhang:2021:IPH

[1023] Wenbo Zhang, Xian Xu, and Huan Long. On the interactive power of higher-order processes extended with parameterization. Formal Aspects of Computing, 33(2):151–183, March 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL https://link.springer.com/
article/10.1007/s00165-020-00524-

Hesselink:2021:UBA

[1024] Wim H. Hesselink. UNITY and Büchi automata. Formal Aspects of Computing, 33(2):185–205, March 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00165-020-00528-
                                        x.

Xie:2021:PCB

[1025] Wanling Xie, Huibiao Zhu, and Qiwen Xu. A process calculus BigrTiMo of mobile systems and its formal semantics. Formal Aspects of Computing, 33(2):207–249, March 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00530-x.

Bozzano:2021:MBS

[1026] Marco Bozzano, Alessandro Cimatti, and Cristian Mattarei. Model-based safety assessment of a triple modular generator with xSAP. Formal Aspects of Computing, 33(2): 251–295, March 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00532-9.

Rossi:2021:MAC

[1027] Matteo Rossi. Modeling and analysis of communicating systems. Formal Aspects of Computing, 33(2): 297–298, March 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00533-8.

Chen:2021:E

[1028] Xiaoping Chen, Zhiming Liu, and Jim Woodcock. Editorial. Formal Aspects of Computing, 33(3): 299–300, June 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00552-5.

Bu:2021:MLS

[1029] Lei Bu, Yongjuan Liang, and Xuandong Li. Machine learning steered symbolic execution framework for complex software code. Formal Aspects of Computing, 33(3):301–323, June 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL https://link.springer.com/
article/10.1007/s00165-021-00538-
3.

Wu:2021:SVS

[1030] Huihui Wu, Deyun Lv, and Weiqiang Kong. SDLV: Verification of steering angle safety for self-driving cars. Formal Aspects of Computing, 33 (3):325–341, June 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00539-2.

Yang:2021:EAI

[1031] Zhibin Yang, Yang Bao, and Zonghua Gu. Exploiting augmented intelligence in the modeling of safetycritical autonomous systems. Formal Aspects of Computing, 33(3): 343–384, June 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00543-6.

Jin:2021:ISN

[1032] Xiangyu Jin, Jie An, and Miaomiao Zhang. Inferring switched nonlinear dynamical systems. Formal Aspects of Computing, 33(3): 385–406, June 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00542-7.

Yang:2021:ERV

[1033] Pengfei Yang, Jianlin Li, and Lijun Zhang. Enhancing robustness verification for deep neural networks via symbolic propagation. Formal Aspects of Computing, 33 (3):407–435, June 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00548-1.

Zhao:2021:LSN

[1034] Hengjun Zhao, Xia Zeng, and Jim Woodcock. Learning safe neural network controllers with barrier certificates. Formal Aspects of Computing, 33(3):437–455, June 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00544-5.

Loreti:2021:SPL

[1035] Michele Loreti. Semantics of the probabilistic Lambda calculus by Dirk Draheim. Formal Aspects of Computing, 33(3):457–458, June 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00534-7.

McIver:2021:E

[1036] Annabelle McIver and Maurice H ter Beek. Editorial. Formal Aspects of Computing, 33(4-5):459–460, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL https://link.springer.com/
article/10.1007/s00165-021-00559-
y.

Tan:2021:AAE

[1037] Yong Kiam Tan and André Platzer.

An axiomatic approach to existence and liveness for differential equations. Formal Aspects of Computing, 33(4-5):461–518, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00525-0.

Tran:2021:VPD

[1038] Hoang-Dung Tran, Neelanjana Pal, and Taylor T. Johnson. Verification of piecewise deep neural networks: a star set approach with zonotope prefilter. Formal Aspects of Computing, 33 (4-5):519–545, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00553-4.

Derrick:2021:VCP

[1039] John Derrick, Simon Doherty, and Heike Wehrheim. Verifying correctness of persistent concurrent data structures: a sound and complete method. Formal Aspects of Computing, 33(4-5):547–573, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00541-8.

Tappler:2021:BLM

[1040] Martin Tappler, Bernhard K. Aichernig, and Kim G. Larsen. L^{∗}- based learning of Markov decision processes (extended version). Formal Aspects of Computing, 33(4-5): 575–615, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00536-5.

deBoer:2021:SEF

[1041] Frank S. de Boer and Marcello Bonsangue. Symbolic execution formally explained. Formal Aspects of Computing, 33(4-5):617–636, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00165-020-00527-
                                        y.

Ceska:2021:CGI

[1042] Milan Ceska, Christian Hensel, and Joost-Pieter Katoen. Counterexampleguided inductive synthesis for probabilistic systems. Formal Aspects of Computing, 33(4-5):637–667, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00165-021-00547-
                                        2.

Evangelidis:2021:QVK

[1043] Alexandros Evangelidis and David Parker. Quantitative verification of Kalman filters. Formal Aspects of Computing, 33(4-5):669–693, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL https://link.springer.com/
article/10.1007/s00165-020-00529-
w.

Wissmann:2021:GPR

[1044] Thorsten Wißmann, Hans-Peter Deifel, and Lutz Schröder. From generic partition refinement to weighted tree automata minimization. Formal Aspects of Computing, 33(4-5): 695–727, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-020-00526-z.

Amram:2021:GGS

[1045] Gal Amram, Shahar Maoz, and Or Pistiner. GR(1)*: GR(1) specifications extended with existential guarantees. Formal Aspects of Computing, 33 (4-5):729–761, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00535-6.

Gleirscher:2021:RDA

[1046] Mario Gleirscher, Radu Calinescu, and Jim Woodcock. RiskStructures : A design algebra for risk-aware machines. Formal Aspects of Computing, 33(4-5):763–802, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00545-4.

Ngondi:2021:DSC

[1047] Gerard Ekembe Ngondi. Denotational semantics of channel mobility in UTP-CSP. Formal Aspects of Computing, 33 (4-5):803–826, August 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00546-3.

Ahrendt:2021:E

[1048] Wolfgang Ahrendt, Silvia Lizeth Tapia Tarifa, and Heike Wehrheim. Editorial. Formal Aspects of Computing, 33(6):827, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00569-w.

Luteberget:2021:DSF

[1049] Bjørnar Luteberget and Christian Johansen. Drawing with SAT: four methods and a tool for producing railway infrastructure schematics. Formal Aspects of Computing, 33(6): 829–854, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00566-z.

Foster:2021:IFP

[1050] Simon Foster, Yakoub Nemouchi, and Tim Kelly. Integration of formal proof into unified assurance cases with Isabelle/SACM. Formal Aspects of Computing, 33(6): 855–884, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00537-4.

Dubslaff:2021:EPM

[1051] Clemens Dubslaff, Patrick Koopmann, and Anni-Yasmin Turhan. Enhancing probabilistic model checking with ontologies. Formal Aspects of Computing, 33(6):885–921, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

URL https://link.springer.com/
article/10.1007/s00165-021-00549-
0.

Fantechi:2021:E

[1052] Alessandro Fantechi, Anne E. Haxthausen, and Jim Woodcock. Editorial. Formal Aspects of Computing, 33 (6):923–924, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00568-x.

Peleska:2021:EDV

[1053] Jan Peleska, Niklas Krafczyk, and Ralf Pinger. Efficient data validation for geographical interlocking systems. Formal Aspects of Computing, 33(6): 925–955, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00551-6.

Basile:2021:AAT

[1054] Davide Basile, Alessandro Fantechi, and Gianluca Mandò. Analysing an autonomous tramway positioning system with the Uppaal statistical model checker. Formal Aspects of Computing, 33(6):957– 987, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00556-1.

Flammini:2021:CMR

[1055] Francesco Flammini, Stefano Marrone, and Valeria Vittorini. Compositional modeling of railway virtual coupling with stochastic activity networks. Formal Aspects of Computing, 33(6): 989–1007, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00560-5.

Stankaitis:2021:RBD

[1056] Paulius Stankaitis, Alexei Iliasov, and Alexander Romanovsky. A refinement-based development of a distributed signalling system. Formal Aspects of Computing, 33(6): 1009–1036, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00567-y.

Cabot:2021:E

[1057] Jordi Cabot, Heike Wehrheim, and Eerke Boiten. Editorial. Formal Aspects of Computing, 33(6): 1037, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00565-0.

Menghi:2021:TWM

[1058] Claudio Menghi, Alessandro Maria Rizzi, and Paola Spoletini. TOr-PEDO: witnessing model correctness with topological proofs. Formal Aspects of Computing, 33(6):

1039–1066, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00564-1.

Stunkel:2021:CSF

[1059] Patrick Stünkel, Harald König, and Adrian Rutle. Comprehensive systems: A formal foundation for multimodel consistency management. Formal Aspects of Computing, 33(6): 1067–1114, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00555-2.

Weidmann:2021:SCC

[1060] Nils Weidmann and Anthony Anjorin. Schema compliant consistency management via triple graph grammars and integer linear programming. Formal Aspects of Computing, 33(6): 1115–1145, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00557-0.

Cordy:2021:SMC

[1061] Maxime Cordy, Sami Lazreg, and Axel Legay. Statistical model checking for variability-intensive systems: applications to bug detection and minimization. Formal Aspects of Computing, 33(6):1147– 1172, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

//link.springer.com/article/10.
1007/s00165-021-00563-2.

deLara:2021:LFE

[1062] Juan de Lara and Esther Guerra. Language family engineering with product lines of multi-level models. Formal Aspects of Computing, 33(6): 1173–1208, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00554-3.

Hennicker:2021:HDL

[1063] Rolf Hennicker, Alexander Knapp, and Alexandre Madeira. Hybrid dynamic logic institutions for event/ data-based systems. Formal Aspects of Computing, 33(6):1209– 1248, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00550-7.

Archibald:2021:TTG

[1064] Blair Archibald, Géza Kulcsár, and Michele Sevegnani. A tale of two graph models: a case study in wireless sensor networks. Formal Aspects of Computing, 33(6):1249– 1277, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00165-021-00558-z.

Wilhelm:2021:FPL

[1065] Reinhard Wilhelm. Foundations of programming languages. Formal Aspects of Computing, 33(6): 1279–1280, December 2021. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic). URL https:

REFERENCES 154

//link.springer.com/article/10.
1007/s00165-021-00561-4.